### v2

# Minimax and adaptive estimation of general linear functionals under sparsity

Abstract
We study estimation of the linear functional $\eta^{\top}\theta$ of a high-dimensional $s$-sparse mean vector $\theta$ when the loading vector $\eta$ is arbitrary and the noise is symmetric with exponentially decaying tails. Previous analyses for equal loadings treat coordinates as exchangeable and do not yield sharp rates when loadings vary. We give a sharp nonasymptotic characterization of the oracle minimax rate that makes explicit its dependence on $s$, $\eta$, and the noise tail parameter. To attain this rate, we construct an estimator that treats large and small loadings differently with a cutoff calibrated to $\eta$, and we prove a matching lower bound using a sparse prior whose inclusion probabilities and signal magnitudes depend on $\eta$. For unknown sparsity, we identify an $\eta$-dependent threshold for a Lepski type selection and show that the resulting estimator achieves the oracle minimax rate up to a logarithmic factor, and that it cannot be improved for a broad, verifiable class of loading vectors. In analytic examples, we demonstrate how heterogeneity in $\eta$ changes the minimax and adaptive rates. We also extend the theory to non-symmetric noise, hypothesis testing, and estimation with unknown noise variance.
1 Introduction
We consider the model
$y_{j}=\theta_{j}+\sigma\xi_{j},\qquad j=1,\ldots,d,$ (1)
where $\theta=(\theta_{1},\ldots,\theta_{d})\in\mathbb{R}^{d}$ is an unknown parameter vector, $\xi_{j}$ are independent centered noise random variables with unit variance, and $\sigma>0$ is the noise level. We study the problem of estimating the following linear functional:
$L(\theta)=\eta^{\top}\theta=\sum_{j=1}^{d}\eta_{j}\theta_{j},$ (2)
based on the observations $y_{1},\ldots,y_{d}$, where $\eta=(\eta_{1},\ldots,\eta_{d})\in\mathbb{R}^{d}$ is a fixed loading vector. Because coordinates with zero loadings can be dropped, we assume that all $\eta_{j}$ ( $j=1,\ldots,d$) are nonzero and sorted such that $|\eta_{1}|\geq|\eta_{2}|\geq\cdots\geq|\eta_{d}|>0$.
Throughout this paper, we assume that $\theta$ is $s$-sparse, i.e.,
$\|\theta\|_{0}=\sum_{j=1}^{d}\mathbbm{1}\{\theta_{j}\neq 0\}\leq s,$
for some integer $s\in\{1,\ldots,d\}$. This sparsity assumption arises in many applications, including spectroscopy, astronomy, and interferometry [12]. In these applications, the observed signal is typically close to zero with a few rare spikes, often described as a nearly black object. Most theoretical work has focused on Gaussian or sub-Gaussian noise, although in practice other noise distributions may also be relevant.
In this paper, we denote by $\mathcal{P}_{j}$ the distribution of noise $\xi_{j}$ ($j=1,\ldots,d$) and by $\mathcal{P}_{\xi}=(\mathcal{P}_{1},\ldots,\mathcal{P}_{d})$ the joint noise distributions. We denote by $\mathbb{P}_{\theta,\mathcal{P}_{\xi}}$ the distribution of $(y_{1},\ldots,y_{d})$ when the signal is $\theta$ and the joint noise distribution is $\mathcal{P}_{\xi}$. We also write $\mathbb{E}_{\theta,\mathcal{P}_{\xi}}$ for the corresponding expectation. The classical Gaussian sequence model corresponds to the case where the noise variables $\xi_{i}$ are standard Gaussian, i.e., $\mathcal{P}_{1}=\cdots=\mathcal{P}_{d}=\mathcal{N}(0,1)$, in which case we write $\mathcal{P}_{\xi}=\mathcal{N}^{\otimes}$.
Beyond the Gaussian setting, we study a broader family of $\mathcal{P}_{\xi}$ that retains symmetry and fast-decaying tails. Specifically, we introduce the following class.
Definition 1 (Symmetric distributions with exponentially decaying tails).
For some $\alpha,\tau>0$, let $\mathcal{G}_{\alpha,\tau}$ denote the class of distributions on $\mathbb{R}$ such that for any $P\in\mathcal{G}_{\alpha,\tau}$ and any random variable $\sim P, WW$ is symmetric around 0, $\mathbb{E}(W^{2})=1$, and
$\forall t\geq 0,\quad\mathbb{P}(|W|\geq t)\leq 2\exp\!\left\{-2\left(\frac{t}{\tau}\right)^{\alpha}\right\}.$
Distributions in $\mathcal{G}_{\alpha,\tau}$ exhibit sub-Weibull tails and have recently attracted considerable attention [11, 23, 20]. The additional symmetry condition is important for our analysis. The class parameter $\alpha$ specifies the tail behavior, and the parameter $\tau$ specifies the spread of the tail. In particular, when $\alpha=2$, the class reduces to a sub-Gaussian class; when $\alpha=2$ and $\tau=2$, it includes the standard Gaussian distribution.
We denote by $\mathcal{G}_{\alpha,\tau}^{\otimes}$ the class of (independent) product distributions on $\mathbb{R}^{d}$ whose marginals all belong to $\mathcal{G}_{\alpha,\tau}$. In other words, $\mathcal{G}_{\alpha,\tau}^{\otimes}=\{\otimes_{j=1}^{d}P_{j}:P_{j}\in\mathcal{G}_{\alpha,\tau},~~j=1,\ldots,d\}$. We assume the noise distribution $\mathcal{P}_{\xi}$ lies in $\mathcal{G}_{\alpha,\tau}^{\otimes}$ with fixed $\alpha$ and $\tau$, but the marginals of $\mathcal{P}_{\xi}$ may differ and remain unknown.
As a measure of the quality of an estimator $\hat{T}$ of the functional $L(\theta)$, we consider the maximum mean squared error over $s$-sparse vectors:
$\sup_{\theta\in\Theta_{s}}\sup_{{\mathcal{P}}_{\xi}\in\mathcal{G}_{\alpha,\tau}^{\otimes}}\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left(\hat{T}-L(\theta)\right)^{2},$
where $\Theta_{s}=\{\theta:\|\theta\|_{0}\leq s\}$. In this paper, we propose rate-optimal estimators in a nonasymptotic minimax sense, that is, estimators $\widetilde{T}$ such that
$\sup_{\theta\in\Theta_{s}}\sup_{{\mathcal{P}}_{\xi}\in\mathcal{G}_{\alpha,\tau}^{\otimes}}\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left(\widetilde{T}-L(\theta)\right)^{2}\asymp\inf_{\hat{T}}\sup_{\theta\in\Theta_{s}}\sup_{{\mathcal{P}}_{\xi}\in\mathcal{G}_{\alpha,\tau}^{\otimes}}\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left(\hat{T}-L(\theta)\right)^{2},$
where $\inf_{\hat{T}}$ denotes the infimum over all estimators.
The minimax rate of estimation under the Gaussian sequence model has been extensively studied in the literature; see, for example, [15, 4, 5, 13, 14, 21]. The most closely related work to ours is [9], which established the nonasymptotic minimax rate for estimating the linear functional with a homogeneous loading vector (i.e., $\eta_{j}=1$ for all $j$) in the Gaussian sequence model. Specifically, for estimating the sum $\sum_{j=1}^{d}\theta_{j}$, they show that
$\inf_{\hat{T}}\sup_{\theta\in\Theta_{s}}\mathbb{E}_{\theta,{\cal N}^{\otimes}}\!\biggl(\hat{T}-\sum_{j=1}^{d}\theta_{j}\biggr)^{2}\;\asymp\;\sigma^{2}s^{2}\log\!\left(1+\frac{d}{s^{2}}\right),$ (3)
and the optimal rate is attained by an estimator of the form
$\hat{T}_{s}=\begin{cases}\sum_{j=1}^{d}y_{j}\,\mathbbm{1}\!\left\{|y_{j}|>\sigma\sqrt{2\log\!\left(1+\frac{d}{s^{2}}\right)}\right\},&s<\sqrt{d},\\[6.0pt] \sum_{j=1}^{d}y_{j},&s\geq\sqrt{d}.\end{cases}$ (4)
This estimator includes all $y_{j}$ when $s$ is large and applies thresholding when $s$ is small. However, their analysis assumes knowledge of the sparsity level $s$, which is typically unknown in practice. To address this issue, [10] revisited the problem under the same setting but without assuming knowledge of $s$, and established that the optimal rate is
$\sigma^{2}s^{2}\log\!\left(1+\frac{d\log d}{s^{2}}\right).$
While the aforementioned theoretical works on the homogeneous loading case yield valuable insights into the roles of sparsity and dimensionality, they do not yield sharp minimax rates for estimating linear functionals with heterogeneous loadings (i.e., unequal $\eta_{j}$ values). Heterogeneous loadings commonly arise in practice, as illustrated in the following two examples.
Example 1 (Estimation of Integrals).
Let $\mathcal{H}$ be the class of square-integrable functions on $[0,1]$. Suppose $f\in\mathcal{H}$ is unknown and $Y$ is an observation from the white noise model
${\rm d\,}Y(t)=f(t)\,{\rm d\,}t+\sigma\,{\rm d\,}W(t)$
where $W(t)$ is a standard Brownian motion. It is often of interest to estimate the integral $\int fg$ for some test function $g\in\mathcal{H}$. To see the connection to our problem, let $\{\varphi_{i}(t)\}$ be any orthonormal basis for $\mathcal{H}$ and define
$y_{j}=\int_{0}^{1}\varphi_{j}\,{\rm d\,}Y,~~\theta_{j}=\int_{0}^{1}f\varphi_{j},~~\xi_{j}=\int_{0}^{1}\varphi_{j}\,{\rm d\,}W,~~\text{ and }\eta_{j}=\int_{0}^{1}\varphi_{j}g.$
It then follows that $\{\xi_{i}\}$ are independent standard normal r.v.s and $\int fg=\sum_{j}\eta_{j}\theta_{j}$. If it is believed that $\theta_{i}=0$ for $i>d$ and $\{\theta_{i}\}_{i=1}^{d}$ has at most $s$ nonzero components, this model coincides with (1) and the estimation reduces to estimating $L(\theta)$ in (2); see [17] for further discussions. In many problems, the test function $g$ is arbitrary, so the induced sequence $\{\eta_{j}\}$ may lack any assumed structural properties. ∎
Example 2 (Prediction in linear regression).
Given $n$ covariate vectors $x_{i}\in\mathbb{R}^{d}$, suppose we observe $z_{i}=x_{i}^{\top}\beta+\sigma\varepsilon_{i}$ where $\beta\in\mathbb{R}^{d}$ is the unknown parameter and $\varepsilon_{i}$ are i.i.d. samples from $\mathcal{N}(0,1)$. A common task is to predict the response for a new fixed covariate vector $x_{0}$ by estimating $x_{0}^{\top}\beta$. To see the connection to our problem, consider an orthogonal design, where the design matrix $\mathbf{X}$ can be written as $\mathbf{X}=\sum_{j\leq d}b_{j}u_{j}e_{j}^{\top}$, with singular values $b_{j}>0$, singular vectors $u_{j}\in\mathbb{R}^{n}$, and canonical basis vectors $e_{j}\in\mathbb{R}^{d}$. For any $j\leq d$, define the transformed variables and parameters as
$y_{j}=\sum_{i\leq n}u_{j,i}z_{i},~~\theta_{j}=b_{j}\beta_{j},~~\xi_{j}=\sum_{i\leq n}u_{j,i}\varepsilon_{i},\text{ and }~~\eta_{j}=b_{j}^{-1}x_{0,j}.$
It follows that $y_{j}=\theta_{j}+\sigma\xi_{j}$, $\xi_{j}$ are i.i.d. from $\mathcal{N}(0,1)$, and $x_{0}^{\top}\beta=\sum_{j=1}^{d}\eta_{j}\theta_{j}=L(\theta)$. The standard assumption that $\beta$ is $s$-sparse is now equivalent to $\theta$ being $s$-sparse. In this case, the model coincides with the sequence model in (1), and the prediction reduces to estimating $L(\theta)$, where the loading vector $\eta=(b_{j}^{-1}x_{0,j})$ is determined by an arbitrary $x_{0}$ and non-uniform singular values $b_{j}$. ∎
In both examples and many other applications, the loading vector is arbitrary and heterogeneous. This presents a critical gap in the existing minimax theories: the sequence model literature [4, 9, 10] assumes equal loadings, while the linear model literature [2, 6] considers special loading structures (specifically, loadings with bounded coordinate ratios or polynomial decay rates). Consequently, the optimal rates for general $\eta$ remain unknown. To our knowledge, the current paper is the first to address this problem.
Contributions. We study estimation of $L(\theta)=\eta^{\top}\theta$ for $s$-sparse $\theta$ with an arbitrary loading vector $\eta$ under any noise distribution $\mathcal{P}_{\xi}\in\mathcal{G}_{\alpha,\tau}^{\otimes}$. All results are nonasymptotic and explicit in $(s,d,\eta,\alpha,\sigma)$.
In the oracle setting where $s$ is known, we establish the minimax rate $\Phi_{o}(s;\eta)$, which is governed by an important quantity $\lambda_{o}$ that depends on $(s,\eta,\alpha)$ in Equation 6. Our theory not only recovers the homogeneous case, but also reveals how heterogeneity in $\eta$ changes the minimax rate. To obtain sharp upper and lower bounds, the derivation should take into account the structure of the loading vector. However, this aspect is completely missing in existing works on homogeneous loadings [4, 9, 10]. To fill this gap, we make the following technical contributions:
  • Upper bounds. The estimator in Equation 4 for homogeneous loadings is no longer optimal for heterogeneous loadings. To achieve minimax optimality with general $\eta$, it is necessary to treat large and small loadings differently. We use $\lambda_{o}$ as the cutoff value for $|\eta_{j}|$ and propose a hybrid estimator that handles $\theta_{j}$ with large $|\eta_{j}|$ by plug-in and those with small $|\eta_{j}|$ by thresholding. The cutoff and the threshold are precisely calibrated to the loading vector $\eta$ and the noise tail parameter $\alpha$; these choices are critical for achieving minimax optimality.
  • Lower bounds. The lower bound arguments in [9] are not sharp for heterogeneous loadings, as they rely on a uniform prior distribution that ignores the differences in $|\eta_{j}|$. Furthermore, their analysis is confined to Gaussian noise. To address these limitations, we develop new techniques that apply to general noise distributions in $\mathcal{G}_{\alpha,\tau}$ and construct a non-uniform sparse prior where the nonzero probabilities and the magnitudes of coordinates are specifically designed based on $\eta$ and $\alpha$. This refined construction yields sharp minimax lower bounds for general loading vectors and a broader family of noise distributions.
For unknown $s$, we investigate the price of adaptation. Unlike the Lepski type procedure for homogeneous loadings in [10], where the threshold for each candidate $s$ is independent of $\eta$ and is not optimal, we identify a new $\eta$-dependent threshold such that the resulting adaptive estimator attains a rate that matches the minimax rate up to a logarithmic factor in $s$. We further show that this rate is the unimprovable adaptive rate for a broad class of loading vectors satisfying a mild and verifiable condition (1).
In addition, we study non-symmetric noise, showing that skewness can increase the minimax rate in certain examples. We also consider extensions to the setting with unknown noise variance $\sigma^{2}$ and to hypothesis testing for the linear functional.
Organization. The rest of the paper is organized as follows. In Section 2, we study minimax estimation of the linear functional $L(\theta)$ under symmetric sub-Weibull noise when the sparsity level $s$ is known. Section 3 considers the more realistic case where $s$ is unknown and establishes the optimal adaptive rate for estimating $L(\theta)$. In Section 4, we extend the analysis to general (not necessarily symmetric) sub-Weibull noise with known sparsity. Section 5 provides analytically tractable examples of the derived rates. Extensions and future directions are discussed in Section 6 and Section 7, respectively. The last section contains the proofs of the lower bounds. Other proofs are provided in the supplementary material.
Notation: We denote by $\mathbb{N}^{*}$ the set of positive integers. For any $k\in\mathbb{N}^{*}$, we write $k]=\{1,\ldots,k\}$. For $\alpha\in\mathbb{R}$, we write $\lfloor\alpha\rfloor$ and $\lceil\alpha\rceil$ for the floor and ceiling functions of $\alpha$, that is, the greatest integer not exceeding $\alpha$ and the smallest integer not less than $\alpha$, respectively. For two real-valued functions $f$ and $g$, we write fgf*g for their convolution. We use $f\lesssim g$ (resp. $f\gtrsim g$) to mean that there exists a constant $C>0$ (resp. $c>0$) such that $f\leq Cg$ (resp. $f\geq cg$). We write $f\asymp g$ if both $f\lesssim g$ and $g\lesssim f$ hold. We respectively denote by $x\vee y$ and $x\wedge y$ the maximum and minimum of the two real values $x$ and $y$, and we set $x_{+}=x\vee 0$. We use $\mathbbm{1}\{\cdot\}$ to denote an indicator function. For any $d\in\mathbb{N}^{*}$, and for any property $P(j)$ over index $j\in[d]$, we set $\max\left\{j\in[d]\mid P(j)\right\}=0$ if for any $j\in[d]$, $P(j)$ is false.
2 Minimax estimation with known sparsity
In this section, we assume that the sparsity level $s$ and the variance $\sigma^{2}$ are known. To present the minimax rate for estimating the linear functional, we first introduce a few definitions.
Given any $s\in[1,d]$, let $\beta=\beta(s;\eta,\alpha)\in\mathbb{R}$ be the unique solution to the equation
$\frac{\sum_{j=1}^{d}|\eta_{j}|\exp(-\beta/|\eta_{j}|^{\alpha})}{\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})}}=\frac{s}{2},$ (5)
and define
$\lambda_{o}=\lambda_{o}(s;\eta,\alpha)=\beta_{+}^{1/\alpha}.$ (6)
We first verify that $\beta$ is well defined. Indeed, the left-hand side of Equation 5 is continuous and strictly decreasing (see Lemma 11 in the supplementary material), diverges to $\infty$ as $\beta\to$-$\infty$, and converges to 0 as $\beta\to+\infty$. It follows that Equation 5 admits a unique solution $\beta$ and we have $\beta_{+}=\beta\vee 0\geq 0$. When $\alpha=2$, Equation 5 coincides with Equation (5) in [8], so our formulation extends their definition from Gaussian noise to symmetric sub-Weibull noise. The parameter $\lambda_{o}$ captures the dependence of the linear functional estimation problem on the three key components: the sparsity level $s$, the loading vector $\eta$, and the noise tail parameter $\alpha$. $\lambda_{o}$ plays a crucial role in characterizing the minimax rate, as it appears both in the expression of our lower bound and in the construction of our optimal estimator.
Our estimator is constructed as follows. Define $j_{1}(s)=\max\left\{j\in[d]:|\eta_{j}|\geq\lambda_{o}\right\}$ and
$\hat{L}_{s}=\sum_{j\leq j_{1}(s)}\eta_{j}y_{j}+\sum_{j>j_{1}(s)}\eta_{j}y_{j}\,\mathbbm{1}\!\left\{\,|\eta_{j}y_{j}|>{~}\sigma\,\tau\,\lambda_{o}\right\}.$ (7)
The following theorem establishes that the minimax rate is given by $\sigma^{2}\Phi_{\rm o}(s;\eta)$, where
$\Phi_{\rm o}(s;\eta)=\left(\lambda_{o}s+\nu\right)^{2}\quad\text{and}\quad\nu=\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp\left(-(\lambda_{o}/|\eta_{j}|)^{\alpha}\right)}$ (8)
and the estimator $\hat{L}_{s}$ is a minimax rate-optimal estimator.
Theorem 1.
Let $\lambda_{o}$ and $\nu$ be defined as in (6) and (8), respectively.
(1). [Lower bound] For $\alpha>0$, there exist constants $c,\tau_{\alpha}>0$ such that for any $\tau\geq\tau_{\alpha}$ and any integer $s\in[d]$, it holds that
$\inf_{\hat{T}}\sup_{\theta\in\Theta_{s}}\sup_{{\mathcal{P}}_{\xi}\in\mathcal{G}_{\alpha,\tau}^{\otimes}}\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\!\left(\hat{T}-L(\theta)\right)^{2}\;\geq\;c\sigma^{2}\Phi_{\rm o}(s;\eta).$ (9)
(2). [Upper bound] For the estimator $\hat{L}_{s}$ defined in (7) and $\alpha,\tau>0$, there exists a constant $C>0$ such that
$\sup_{\theta\in\Theta_{s}}\sup_{{\mathcal{P}}_{\xi}\in\mathcal{G}_{\alpha,\tau}^{\otimes}}\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\!\left(\hat{L}_{s}-L(\theta)\right)^{2}\;\leq\;C\sigma^{2}\Phi_{\rm o}(s;\eta).$ (10)
Theorem 1 may seem related to Theorem 1 of [8], who investigated the minimax separation distance in the heteroscedastic Gaussian sequence model for testing
$H_{0}:\;\theta=0\quad\text{against}\quad H_{1}:\;\|\theta\|_{t}\geq\varepsilon,$
where $\|\theta\|_{t}=(\sum_{j=1}^{d}|\theta_{j}|^{t})^{1/t}$ for $t\geq 1$. As discussed in their work, the theory of testing in this framework is closely connected to minimax estimation of the norm $\|\theta\|_{t}$. However, the case $t=1$ considered there fundamentally differs from our setting. The functional $\|\theta\|_{1}=\sum_{j=1}^{d}|\theta_{j}|$ is nonlinear and invariant to sign changes, whereas our target $L(\theta)$ is a linear and sign-sensitive functional: flipping the sign of any nonzero coordinate leaves $\|\theta\|_{1}$ unchanged but alters $L(\theta)$. This structural difference implies that the analytical tools and lower bound constructions used in norm estimation problems cannot be directly extended to our setting. In addition, the analysis in [8] does not accommodate non-Gaussian noise.
Theorem 1 extends the seminal result of [9], which focused on the homogeneous setting where all $\eta_{j}=1$, to the more general and practically relevant case of arbitrary loading vectors. This generalization reveals how heterogeneity in the loadings fundamentally alters the minimax separation rate and highlights the role of $\eta$ in governing the intrinsic difficulty of linear functional estimation. To illustrate these insights concretely, we provide in Section 5 explicit expressions for the minimax rate under several representative examples of loading structures, including the classical homogeneous case.
To interpret the minimax rate, we state an alternative expression for $\Phi_{\rm o}(s;\eta)$, which is analogous to [8, Corollary 1].
Proposition 1.
Recall the definitions of $\lambda_{o}$, $\nu$, and $j_{1}(s)$. For any $s\in[d]$, we have
$\Phi_{\rm o}(s;\eta)=\left(\lambda_{o}s+\nu\right)^{2}\;\asymp\;\lambda_{o}^{2}s^{2}+\nu^{2}\;\asymp\;\lambda_{o}^{2}s^{2}+\sum_{j\leq j_{1}(s)}\eta_{j}^{2}.$
In view of Proposition 1, we can decompose the estimation error of $\hat{L}_{s}$ into two parts:
  • Large loading components at $1\leq j\leq j_{1}(s)$. For components with large absolute loadings, the sparsity structure is not utilized and we estimate $\sum_{j\leq j_{1}(s)}\eta_{j}\theta_{j}$ by the plug-in estimator $\sum_{j\leq j_{1}(s)}\eta_{j}y_{j}$. The squared error of this term is of order $\sigma^{2}\sum_{j\leq j_{1}(s)}\eta_{j}^{2}$.
  • Small loading components at $j>j_{1}(s)$. For components with small absolute loadings, we exploit the sparsity assumption by thresholding. Specifically, we retain indices where the observations are sufficiently strong, i.e., $|\eta_{j}y_{j}|>{~}\tau\sigma\lambda_{o}$. This threshold is crucial for achieving minimax optimality. The variance after thresholding is dominated by the variance from the large loading part, and the main error is the squared bias induced by thresholding, which is of order $\sigma^{2}\lambda_{o}^{2}s^{2}$.
The above decomposition is crucial in the analysis of $\hat{L}_{s}$ for general loading vectors, where the magnitudes of the loadings can vary widely. In this case, components with large loadings are treated using a plug-in strategy while those with small loadings require thresholding. In contrast, in the homogeneous case (where $\eta_{j}=1,\forall j$) studied in [9], we have either $j_{1}(s)=0$ or $j_{1}(s)=d$. In this case, depending on the sparsity level $s$, the optimal estimator $\hat{L}_{s}$ reduces to applying a single estimation strategy (either plug-in or thresholding) uniformly across all components; see Equation 4 for $\alpha=2$ and $\tau=2$. Hence, the decomposition is only necessary in the heterogeneous setting.
For the lower bound, we apply Le Cam’s method (also known as the “method of two fuzzy hypotheses” [27]), which is standard in the functional estimation literature. The idea is to construct two priors for $\theta$, say $\mu_{1}$ and $\mu_{2}$, that stochastically separate $L(\theta^{(1)})$ from $L(\theta^{(2)})$ as much as possible for $\theta^{(1)}\sim\mu_{1}$ and $\theta^{(2)}\sim\mu_{2}$, while ensuring that the total variation distance between their induced sampling distributions of $(y_{j})_{j=1}^{d}$ remains small. The difficulty of deriving a sharp lower bound lies in maximizing the separation.
In previous works [9, 10], the priors were chosen such that $\mu_{1}$ is the point mass at 0, while $\mu_{2}$ is the uniform distribution over the set of exact $s$-sparse vectors with nonzero entries equal to a fixed constant $\rho$. This construction fulfills the sparsity assumption and ensures that $L(\theta^{(2)})$ equals $s\rho$ for $\theta^{(2)}\sim\mu_{2}$ while $L(\theta^{(1)})=0$ for $\theta^{(1)}\sim\mu_{1}$. This approach yields the sharp lower bound in the homogeneous loading case, but it is not satisfactory for heterogeneous loading vectors. In order to maximize the separation of $L(\theta^{(2)})$ from 0 under $\mu_{2}$, coordinates with larger absolute loadings should have higher chances to be nonzero.
To address this, we adopt the random sparsity strategy of [8], where each coordinate is independently nonzero with probability $\pi_{j}$ and, if nonzero, takes value $\gamma_{j}$. This approach offers two advantages: (i) each coordinate is treated independently, which enables computation of the $\chi^{2}$-divergence using our new tool in Lemma 2, specifically developed for sub-Weibull noise distributions, thereby allowing us to control the total variation distance; and (ii) the probabilities $\pi_{j}$ and magnitudes $\gamma_{j}$ can be tailored according to the loading vector $\eta$. A drawback of this construction is that the values of $\|\theta^{(2)}\|_{0}$ and $L(\theta^{(2)})$ become random under $\theta^{(2)}\sim\mu_{2}$, but we can establish probability inequalities to control both quantities. Full details are provided in Section 8.
3 Adaptation to unknown sparsity
In the previous section, we proposed a rate-optimal estimator that requires knowledge of the sparsity level $s$. When $s$ is unknown, it is not implementable; if the supplied $s$ is misspecified, it may underperform. In this section, we turn to the more realistic setting where the sparsity level $s$ is unknown, and our purpose is to construct an adaptive estimator that achieves the optimal rate of convergence for estimating the linear functional $L(\theta)$.
To construct the adaptive estimator, we first define a class of nonadaptive estimators $\{\hat{L}_{s}^{*}:s\in[d]\}$ as follows. For any $s\in[1,d]$, let $\beta_{*}(s)$ be the solution to the following equation:
$\frac{\sum_{j=1}^{d}|\eta_{j}|\exp(-\beta_{*}/|\eta_{j}|^{\alpha})}{\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta_{*}/|\eta_{j}|^{\alpha})}}=\frac{s}{2\sqrt{\log(es)}},$ (11)
and define $\lambda_{*}(s)=(\beta_{*}(s))_{+}^{1/\alpha}$ and $j_{2}(s)=\max\left\{j\in[d]:|\eta_{j}|\geq\lambda_{*}(s)\right\}$. The nonadaptive estimators are defined as
$\hat{L}_{s}^{*}=\sum_{j\leq j_{2}(s)}\eta_{j}y_{j}+\sum_{j>j_{2}(s)}\eta_{j}y_{j}\mathbbm{1}\left\{|\eta_{j}y_{j}|>{~}\sigma\tau\lambda_{*}(s)\right\},\text{ for }s=1,\ldots,d.$ (12)
Equation 11 differs from Equation 5 (for known sparsity) because its right-hand side involves the additional term $\sqrt{\log(es)}$ in the denominator. This extra factor is crucial for achieving adaptivity. The existence and uniqueness of $\lambda_{*}(s)$ can be ensured by Lemma 11 in the supplementary material.
The following result provides an upper bound on the risk of the nonadaptive estimator $\hat{L}_{s}^{*}$ defined in (12) when $\theta\in\Theta_{s}$ holds.
Theorem 2.
For any $\alpha,\tau>0$, there exists a constant $c_{1,0}>0$, depending only on $\alpha$ and $\tau$, such that for any integer $s\in[d]$, $\forall\,\theta\in\Theta_{s}$, and any ${\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}$, it holds that
$\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\!\left(\hat{L}_{s}^{*}-L(\theta)\right)^{2}\;\leq\;c_{1,0}\,\sigma^{2}\left[\sum_{j=1}^{d}\eta_{j}^{2}\exp\!\left(-\left(\frac{\lambda_{*}(s)}{|\eta_{j}|}\right)^{\alpha}\right)+s^{2}\lambda_{*}^{2}(s)\right].$
We now introduce the adaptive estimator that does not rely on knowledge about $\|\theta\|_{0}$. Motivated by [10], the adaptive estimator is selected from the collection of nonadaptive estimators $\{\hat{L}_{s}^{*}\}$ via a Lepski type scheme. Define $s_{*}=\max\left\{s\in[d]:\lambda_{*}(s)>0\right\}$ and $s_{0}=s_{*}+1$. For any $s\in[1,d]$, define
$\nu_{*}(s)=\sqrt{\log(es)\sum_{j=1}^{d}\eta_{j}^{2}\exp\left(-\left(\frac{\lambda_{*}(s)}{|\eta_{j}|}\right)^{\alpha}\right)},$
and
$\Phi_{*}(s;\eta)=\left\{\begin{array}[]{cc}s^{2}\lambda_{*}^{2}(s)+\nu_{*}^{2}(s),&s\leq s_{0};\\ \Phi_{*}(s_{0};\eta),&{\rm otherwise.}\end{array}\right$.
Our Lepski type selection uses the thresholds $\omega_{s}=\left[\zeta\sigma^{2}{\Phi}_{\rm adp}(s;\eta)\right]^{1/2}$ where
${\Phi}_{\rm adp}(s;\eta)=\left\{\begin{array}[]{cc}\Phi_{*}(s;\eta)\vee\Phi_{*}(1;\eta)\log^{2}(e(s\wedge s_{0})),&0<\alpha<2,\\ \Phi_{*}(s;\eta),&{\rm otherwise,}\end{array}\right$. (13)
and $\zeta>0$ is a constant that will be chosen large enough. The selected index $\hat{s}$ is defined as
$\hat{s}=\min\left\{s\in\{1,\cdots,s_{*}\}:|\hat{L}_{s}^{*}-\hat{L}_{s^{\prime}}|\leq\omega_{s^{\prime}}\text{ for all integers}\,s^{\prime}>s\right\}$ (14)
with the convention that $\hat{s}=s_{0}$ if the set in (14) is empty. Intuitively, $\hat{s}$ is chosen as the smallest sparsity level $s$ such that the estimate $\hat{L}_{s}^{*}$ is already “stable,” in the sense that increasing the assumed sparsity level does not lead to substantially different estimates. The adaptive estimator is then defined as
$\tilde{L}_{*}=\hat{L}_{\hat{s}}^{*}.$ (15)
The following theorem establishes an upper bound on the risk of $\tilde{L}_{*}$.
Theorem 3.
For any $\alpha,\tau>0$, we can choose $\zeta$ sufficiently large such that for any integer $s\in[d]$,
$\sup_{\theta\in\Theta_{s}}\;\sup_{{\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}}\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\!\left(\tilde{L}_{*}-L(\theta)\right)^{2}\;\leq\;C\sigma^{2}\Phi_{\rm adp}(s;\eta),$ (16)
for some absolute constant $C>0$.
The definition of $\Phi_{\rm adp}(s;\eta)$ involves two terms, among which $\Phi_{*}(s;\eta)$ typically dominates, as detailed below. $\Phi_{*}(s;\eta)$ naturally connects with the quantities $\Phi_{\rm o}(\cdot;\eta)$ introduced in Section 2. Following an argument analogous to that of Proposition 1, we can show
$\Phi_{*}(s;\eta)\;\asymp\;\log(es)\sum_{j\leq j_{2}(s)}\eta_{j}^{2}\;+\;s^{2}\lambda_{*}^{2}(s),\forall s\leq s_{0}.$
Since Equation 11 is the same as Equation 5 if $s$ is replaced by $s/\sqrt{\log(es)}$, we have $j_{2}(s)=j_{1}(s/\sqrt{\log(es)})$ and $\lambda_{*}(s)=\lambda_{o}(s/\sqrt{\log(es)})$. It then follows that
$\Phi_{*}(s;\eta)\asymp\log(es)\bigg[\sum_{j\leq j_{2}(s)}\eta_{j}^{2}+\bigg(\frac{s}{\sqrt{\log(es)}}\bigg)^{2}\lambda_{*}^{2}(s)\bigg]\asymp\log(es)\,\Phi_{\rm o}\biggl(\frac{s}{\sqrt{\log(es)}};\eta\biggr).$ (17)
Note that $\Phi_{*}(1;\eta)=\Phi_{\rm o}(1;\eta)$. To ensure that $\Phi_{\rm adp}(s;\eta)\asymp\Phi_{*}(s;\eta)$, it suffices to require
$\Phi_{\rm o}(1;\eta)\,\log(es)\;\lesssim\;\Phi_{\rm o}\!\left(\frac{s}{\sqrt{\log(es)}};\eta\right).$ (18)
This suggests that if the minimax rate $\Phi_{\rm o}(s;\eta)$ grows at least on the order of $\log(s)$, then we can conveniently equate ${\Phi}_{\rm adp}(s;\eta)$ with $\Phi_{*}(s;\eta)$. In Section 4, we will have
$\Phi_{\rm o}(1;\eta)\;\lesssim\;\sum_{j\leq\lceil\log^{2/\alpha}d\rceil}\eta_{j}^{2}\quad\text{and}\quad\Phi_{\rm o}(s;\eta)\;\gtrsim\;\sum_{j\leq(s^{2}\wedge d)}\eta_{j}^{2}.$
These bounds demonstrate that the growth condition in Equation 18 is mild and can be verified in some concrete examples, such as those in Section 5.
The following result shows that for any $\eta$, $\Phi_{\rm adp}(s;\eta)$ is “almost increasing” in $s$ while the ratio $\Phi_{\rm adp}(s;\eta)/(s^{2}\log(es))$ is “almost decreasing” in $s$.
Proposition 2.
The following relationships hold for any $s,s^{\prime}\in(0,s_{0}]$ with $s\leq s^{\prime}$:
$\frac{\Phi_{*}(s;\eta)}{\log(es)}\;\lesssim\;\frac{\Phi_{*}(s^{\prime};\eta)}{\log(es^{\prime})},\qquad\frac{\Phi_{*}(s^{\prime};\eta)}{(s^{\prime})^{2}\log(es^{\prime})}\;\lesssim\;\frac{\Phi_{*}(s;\eta)}{s^{2}\log(es)}.$
Consequently, there exists a constant $c_{2,0}>0$ such that
$\Phi_{\rm adp}(s;\eta)\;\leq\;c_{2,0}\,\Phi_{\rm adp}(s^{\prime};\eta),\quad\forall s,s^{\prime}\in(0,d]\text{ with }s\leq s^{\prime},$
and
$\Phi_{\rm adp}(s;\eta)\;\leq\;c_{2,0}s^{2}\log(es)\,\Phi_{\rm o}(1;\eta),\qquad\forall s\in(0,s_{0}].$
From Proposition 2, we obtain
$\Phi_{*}(s;\eta)\;\lesssim\;\Phi_{\rm o}(s;\eta)\,\log(e(s\wedge s_{0})).$
Consequently, the risk of adaptive estimator $\tilde{L}_{*}$ is within a $\log^{2}(e(s\wedge s_{0}))$ factor of the minimax risk when $s$ is known. Moreover, if $\Phi_{*}(s;\eta)\asymp\Phi_{\rm adp}(s;\eta)$, then the adaptation cost reduces to at most a factor of $\log(e(s\wedge s_{0}))$.
According to the second part of Proposition 2, the upper bounds established for the nonadaptive estimation in Theorem 1 and for the adaptive estimation in Theorem 3 grow with $s$ at most on the order of $s^{2}\log(es)$. Controlling the upper bounds at this rate is important for the proof of Theorem 3.
As discussed earlier, [10] studied adaptive estimation under the homogeneous loading vector, i.e., $\eta_{j}=1$ for all $j\in[d]$. Theorem 3 extends their results to the case of general loading vectors. Although both [10] and our work employ Lepski’s method to construct the adaptive estimator, extending from the homogeneous case to the heterogeneous case is highly nontrivial and considerably more challenging than in the nonadaptive setting. The main difficulty lies in the fact that with general loadings, there is no closed-form expression for the threshold $\lambda_{*}(s)$, and hence the analysis in [10] cannot be directly applied. To overcome this, we develop a new analysis approach that exploits key properties of the rate $\Phi_{\rm adp}(s;\eta)$ established in Proposition 2.
We now turn to the optimality of the upper bound $\sigma^{2}\Phi_{\rm adp}(s;\eta)$ for $\tilde{L}_{*}$ established in Theorem 3. The following theorem establishes a complementary lower bound: if an estimator achieves a risk sufficiently smaller than $\Phi_{\rm adp}(s;\eta)$ over $s$-sparse vectors, then its maximal risk over $1$-sparse vectors can be significantly larger than the minimax nonadaptive rate $\Phi_{\rm o}(1;\eta)$.
Theorem 4.
For any $\alpha>0$ and $\gamma\in(0,2)$, there exist constants $\tau_{\alpha},C_{0},C_{1},C_{2}>0$ such that the following holds for all $\tau\geq\tau_{\alpha}$. If an estimator $\hat{T}$ satisfies that
$\sup_{\theta\in\Theta_{s}}\sup_{{\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}}\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left(\hat{T}-L(\theta)\right)^{2}\leq\frac{\sigma^{2}}{C_{0}}\cdot\Phi_{\rm adp}(s;\eta)\text{ for some }s\in[d],$ (19)
then its maximal risk over $\Theta_{1}=\{\theta:\|\theta\|_{0}\leq 1\}$ is lower bounded as
$\sup_{\|\theta\|_{0}\leq 1}\sup_{{\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}}\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left(\hat{T}-L(\theta)\right)^{2}\geq\sigma^{2}C_{2}\max\left\{\frac{\Phi_{\rm adp}(s;\eta)}{(s\wedge s_{0})^{\gamma}},\Phi_{\rm o}(1;\eta)\right\}.$ (20)
We illustrate the results of Theorem 4 under the homogeneous loading case. In this setting, for $s=d^{\gamma^{\prime}}$ with some $\gamma^{\prime}\in(0,1/2)$, we can compute (for more details, see Section 5)
$\Phi_{\rm adp}(s;\eta)\;\asymp\;s^{2}\log^{2/\alpha}\!\left(1+\frac{d\log(es)}{s^{2}}\right)$
and $s_{0}\asymp\sqrt{d\log d}$. Consequently, Theorem 4 (with $\gamma=1$) implies that any estimator whose risk is substantially smaller than $\Phi_{\rm adp}(s;\eta)$ for $s$-sparse vectors must incur a maximal risk of at least $d^{\gamma^{\prime}}$ over $1$-sparse vectors. Since $d^{\gamma^{\prime}}\gg\Phi_{\rm o}(1;\eta)\asymp\log d$ as $d\to\infty$, the lower bound in Equation 20 is much larger than that of $\Phi_{\rm adp}(1;\eta)$, and thus will be larger than the risk of $\tilde{L}_{*}$ in the case where $\|\theta\|_{0}\leq 1$. Hence, such an estimator cannot be considered satisfactory.
Following [26, 10], we now provide the formal definition of the adaptive rate in the asymptotic context where $d\to\infty$ and $\eta$ is a sequence of corresponding loading vectors.
Definition 2.
Suppose $d_{0}$ is a fixed integer and $\{\eta^{(d)}\in\mathbb{R}^{d}\}_{d=d_{0}}^{\infty}$ is a deterministic sequence of loading vectors. For any $d\geq d_{0}$, we consider $L(\theta)=\langle\theta,\eta^{(d)}\rangle$. We call a function $(s,d)\mapsto\psi_{*}(s;\eta^{(d)})$ the adaptive rate of convergence on the scale of classes $\left\{\Theta_{s}\times{\cal G}_{\alpha,\tau}^{\otimes}:\forall s\in[d]\right\}$ if the following holds:
  1. 1.
    There exists an estimator $\hat{L}_{s}$ such that, for all $d\geq d_{0}$,
    $\max_{s\in[d]}\sup_{\theta\in\Theta_{s}}\sup_{{\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}}\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left(\hat{L}_{s}-L(\theta)\right)^{2}/\psi_{*}(s;\eta^{(d)})\leq C,$ (21)
    where $C>0$ is a constant.
  2. 2.
    If there exists another function $s\mapsto\psi_{*}^{\prime}(s;\eta^{(d)})$ and a constant $C^{\prime}>0$ such that, for all $d\geq d_{0}$,
    $\inf_{\hat{T}}\max_{s\in[d]}\sup_{\theta\in\Theta_{s}}\sup_{{\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}}\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left(\hat{T}-L(\theta)\right)^{2}/\psi_{*}^{\prime}(s;\eta^{(d)})\leq C^{\prime},$ (22)
    and
    $\min_{s\in[d]}\frac{\psi_{*}^{\prime}(s;\eta^{(d)})}{\psi_{*}(s;\eta^{(d)})}\to 0\quad\text{as }d\to\infty,$ (23)
    then there exists (a sequence of) $\bar{s}\in\{1,\ldots,d\}$ such that
    $\frac{\psi_{*}^{\prime}(\bar{s};\eta^{(d)})}{\psi_{*}(\bar{s};\eta^{(d)})}\cdot\min_{s\in[d]}\frac{\psi_{*}^{\prime}(s;\eta^{(d)})}{\psi_{*}(s;\eta^{(d)})}\to\infty\quad\text{as }d\to\infty.$ (24)
Since $\eta^{(d)}$ is deterministic and its dependence on $d$ is clear, we drop the superscript in $\eta$ in the following discussion of adaptive rates. In Definition 2, the function $\psi_{*}(s;\eta)$ is an adaptive rate of convergence if any local improvement over this rate for some $s$ (cf. (23)) necessarily incurs a substantially larger loss for at least one other sparsity level $\bar{s}$ (cf. (24)).
Motivated by the illustration following Theorem 4, we consider $\sigma^{2}\Phi_{\rm adp}(s;\eta)$ as the candidate adaptive rate for general loading vectors $\eta$. In the homogeneous loading case in [10], both the minimax rate and the lower bound precluding any local improvement are available in closed form, which readily identifies the adaptive rate. However, such explicit formulas are typically unavailable for general $\eta$, so it is not evident that an estimator with local improvement must incur a larger loss on $\Theta_{1}$. To establish $\sigma^{2}\Phi_{\rm adp}(s;\eta)$ as the adaptive rate, we consider the following condition on $\eta$.
Condition 1.
There exist constants $c>0$ and $\gamma_{0}\in(0,2)$, and a sequence $s_{\rm cut}=s_{\rm cut}(d)\in[d]$ that diverges as $d\to\infty$ such that
$\frac{\Phi_{\rm adp}(s;\eta)}{\Phi_{\rm o}(s;\eta)}\leq c,\quad\forall 1\leq s\leq s_{\rm cut},$ (25)
and
$\frac{\Phi_{\rm adp}(s;\eta)}{\Phi_{\rm o}(1;\eta)}\gtrsim(s\wedge s_{0})^{\gamma_{0}},\quad\forall s\geq s_{\rm cut}.$ (26)
In the above condition, Equation 25 guarantees that the candidate adaptive rate $\sigma^{2}\Phi_{\rm adp}(s;\eta)$ is of the same order as the minimax (nonadaptive) rate $\sigma^{2}\Phi_{\rm o}(s;\eta)$ when the sparsity level is not too large ($s\leq s_{\rm cut}$), while condition (26) essentially ensures that the ratio $\Phi_{\rm adp}(s;\eta)/\Phi_{\rm o}(1;\eta)$ grows faster than some polynomial in $s$ (with exponent $\gamma_{0}$) for large sparsity levels ($s\geq s_{\rm cut}$).
1 only depends on the loading vector $\eta$ and is thus verifiable. In Section 5, we will show that 1 is mild and is satisfied by many loading vectors, including the homogeneous loadings.
Proposition 3.
Under 1, the adaptive rate of convergence on the scale of classes $\left\{\Theta_{s}\times{\cal G}_{\alpha,\tau}^{\otimes}:s\in[d]\right\}$ is given by $\sigma^{2}\Phi_{\rm adp}(s;\eta)$.
Proof Sketch.
Theorem 3 shows that $\sigma^{2}\Phi_{\rm adp}(s;\eta)$ is attained by $\tilde{L}_{*}$ (cf. Equation 21 in Definition 2).
  1. 1.
    If $s_{0}$ is bounded, then ${\Phi_{\rm adp}(s;\eta)}/{\Phi_{\rm o}(s;\eta)}\;\leq\;\log(es_{0}),$ which is itself bounded. Hence, $\Phi_{\rm adp}(s;\eta)$ reaches the minimax (nonadaptive) rate and is therefore optimal.
  2. 2.
    If $s_{0}\to\infty$:
    • For small sparsity levels $s\leq s_{\rm cut}$, the rate $\sigma^{2}\Phi_{\rm adp}(s;\eta)$ is optimal since Equation 25 in Assumption 1 implies that it matches the minimax (nonadaptive) rate. Therefore, the local improvement can only happen at larger sparsity levels $s\geq s_{\rm cut}$.
    • For $s\geq s_{\rm cut}$, the rate $\sigma^{2}\Phi_{\rm adp}(s;\eta)$ remains near-optimal in the sense that, by Proposition 2, it exceeds the minimax (nonadaptive) rate by at most a logarithmic factor $\log^{2}(es)$. This means that for any function $\psi^{\prime}(s;\eta)$ being the risk rate of some estimator (in the sense of Equation 22), we have $\psi^{\prime}(s;\eta)/\Phi_{\rm adp}(s;\eta)\gtrsim\log^{-2}(es)$. Suppose $\min_{s\geq s_{\rm cut}}\psi^{\prime}(s;\eta)/\Phi_{\rm adp}(s;\eta)\to 0$ as $d\to\infty$ (cf. Equation 23), then there is a diverging sequence of $\tilde{s}$ and a sequence of $\hat{T}$ such that Equation 19 in Theorem 4 holds for $s=\tilde{s}$. Theorem 4 (with $\gamma=\gamma_{0}/2$) then establishes that the maximal risk of $\hat{T}$ over the class $\Theta_{1}$ must be lower bounded by $\sigma^{2}C_{2}\Phi_{\rm adp}(\tilde{s};\eta)/{(\tilde{s}\wedge s_{0})^{\gamma_{0}/2}}$. It then follows from Equation 26 in Assumption 1 that $\psi^{\prime}(1;\eta)/\Phi_{\rm adp}(1;\eta)\gtrsim{(\tilde{s}\wedge s_{0})^{\gamma_{0}/2}}$. Since $\psi^{\prime}(1;\eta)/\Phi_{\rm adp}(1;\eta)\min_{s\geq s_{0}}[\psi^{\prime}(s)/\Phi_{\rm adp}(s;\eta)]\gtrsim(\tilde{s}\wedge s_{0})^{\gamma_{0}/2}\log^{-2}(e(s\wedge s_{0}))\to\infty$, Equation 24 in Definition 2 is met with $\bar{s}=1$.
4 General exponentially decaying noise distributions
In the previous sections, we have focused on symmetric noise distributions. In this section, we show that the symmetry assumption plays a crucial role in establishing the preceding results; once this assumption is relaxed, the minimax rate may increase.
To formalize the problem, we consider the following larger class of noise distributions that are not necessarily symmetric (referred to as non-symmetric).
Definition 3 (Distributions with exponentially decaying tails).
For some $\alpha,\tau>0$, let $\mathcal{H}_{\alpha,\tau}$ denote the class of distributions on $\mathbb{R}$ such that for any $P\in\mathcal{H}_{\alpha,\tau}$ and random variable $W\sim P$, it holds that
$\mathbb{E}(W)=0,\quad\mathbb{E}(W^{2})=1,\quad\text{and}\quad\forall t\geq 0,\quad\mathbb{P}(|W|\geq t)\leq 2\exp\!\left\{-\left(\frac{t}{\tau}\right)^{\alpha}\right\}.$
We denote by $\mathcal{H}_{\alpha,\tau}^{\otimes}$ the class of product distributions on $\mathbb{R}^{d}$ whose marginal distributions belong to $\mathcal{H}_{\alpha,\tau}$. In other words, $\mathcal{H}_{\alpha,\tau}^{\otimes}=\left\{\bigotimes_{j=1}^{d}P_{j}:P_{j}\in\mathcal{H}_{\alpha,\tau},\;j=1,\ldots,d\right\}$.
For any $s\in[d]$, let $\lambda_{\cal H}\geq 0$ be the unique solution to
$\sum_{j\geq s^{2}}\exp\!\left(-\left|\frac{\lambda}{\eta_{j}}\right|^{\alpha}\right)=s.$ (27)
The left-hand side of (27) is continuous and strictly decreasing in $\lambda$, tending to $d-s^{2}+1$ as $\lambda\to 0$ and to 0 as $\lambda\to+\infty$. Therefore, equation (27) admits a unique solution whenever $s^{2}+s\leq d+1$.
The following theorem provides a lower bound on the maximal risk of any estimator of $L(\theta)$ when the noise distribution belongs to $\mathcal{H}_{\alpha,\tau}^{\otimes}$.
Theorem 5.
For any $\alpha>0$, there exist $\tau_{\alpha}>0$ and $c>0$ such that for all $\tau\geq\tau_{\alpha}$ and $s\in[d]$, it holds that
$\inf_{\hat{L}_{s}}\sup_{\theta\in\Theta_{s}}\sup_{\mathcal{P}_{\xi}\in\mathcal{H}_{\alpha,\tau}^{\otimes}}\mathbb{E}_{\theta,\mathcal{P}_{\xi}}\left(\hat{L}_{s}-L(\theta)\right)^{2}\geq c\sigma^{2}\Phi_{\rm o}(s;\eta).$ (28)
Furthermore, if (i) $s^{2}+s\leq d+1$ and (ii) there is some constant $\bar{C}$ such that
$\sum_{j\geq s^{2}}\frac{\lambda_{\cal H}^{4}}{\eta_{j}^{4}}\exp\left(-2\left|\frac{\lambda_{\cal H}}{\eta_{j}}\right|^{\alpha}\right)\leq\bar{C},$ (29)
where $\lambda_{\cal H}$ is defined in (27), then there is some constant $c^{\prime}>0$ depending on $(\alpha,\tau,\bar{C})$ such that
$\inf_{\hat{L}_{s}}\sup_{\theta\in\Theta_{s}}\sup_{\mathcal{P}_{\xi}\in\mathcal{H}_{\alpha,\tau}^{\otimes}}\mathbb{E}_{\theta,\mathcal{P}_{\xi}}\left(\hat{L}_{s}-L(\theta)\right)^{2}\geq c^{\prime}\sigma^{2}s^{2}\lambda_{\cal H}^{2}.$ (30)
Theorem 5 establishes two complementary lower bounds. The bound in (28) follows directly from Theorem 1, since ${\cal G}_{\alpha,\tau}^{\otimes}\subseteq{\cal H}_{\alpha,\tau}^{\otimes}$ according to Definitions 1 and 3. We refer to this as the symmetric lower bound, as it is the minimax rate under symmetric noise distributions.
In contrast, the bound in (30) is obtained from a construction that explicitly exploits the asymmetry of the noise distribution; we therefore refer to it as the asymmetric lower bound. As demonstrated in Section 5, when the loading vector is homogeneous, the asymmetric lower bound can be strictly larger than its symmetric counterpart for $s=o(\sqrt{d})$, indicating that the results established in the symmetric setting no longer hold once the noise distribution loses symmetry.
The key technical ingredient of the proof for (30) is the use of the Hellinger distance, rather than the $\chi^{2}$-divergence, to control the total variation distance. This idea has appeared in earlier works such as [10] and [23], though primarily in the context of estimating the quadratic functional $\|\theta\|_{2}^{2}=\sum_{j=1}^{d}\theta_{j}^{2}$, and hence cannot be directly applied to our setting. To overcome this challenge, we again employ the random sparsity strategy and construct a novel prior. The proof may be of independent theoretical interest; detailed derivations are omitted here for brevity, and the full arguments are provided in the supplementary material.
The lower bound highlights the challenge introduced by asymmetry in the noise distribution. To complement this result, we establish an upper bound that generally matches the lower bound up to a logarithmic factor and can coincide with it in certain regimes. For any integer $s\in[d]$, define
$j_{3}(s)=\big\lceil s^{2}\log^{2/\alpha}(ed/s)\big\rceil\wedge d,$ (31)
and consider the estimator
$\hat{L}_{\mathcal{H}}=\sum_{j\leq j_{3}(s)}\eta_{j}y_{j}+\sum_{j>j_{3}(s)}\eta_{j}y_{j}\,\mathbbm{1}\!\left\{|y_{j}|>c_{\mathcal{H}}\sigma\log^{1/\alpha}\!\left(\frac{ed}{s}\right)\right\},$ (32)
where $c_{\mathcal{H}}>0$ is a constant depending only on $\alpha$ and $\tau$.
The next theorem gives an upper bound on the risk of the estimator $\hat{L}_{\cal H}$.
Theorem 6.
Let $s$ and $d$ be integers such that $1\leq s\leq d$. Then, for any $\alpha>0,\tau>0$, there exist some constants $c_{\cal H}$ and $C$ depending on $\alpha$ and $\tau$ such that the following holds for the estimator in Equation 32:
$\sup_{\theta\in\Theta_{s}}\sup_{\mathcal{P}_{\xi}\in\mathcal{H}_{\alpha,\tau}^{\otimes}}\mathbb{E}_{\theta,\mathcal{P}_{\xi}}\left(\hat{L}_{\cal H}-L(\theta)\right)^{2}\leq C\sigma^{2}{\sum_{j\leq j_{3}(s)}\eta_{j}^{2}}.$ (33)
Similar to the estimator in Equation 7 for symmetric noise, the estimator $\hat{L}_{\mathcal{H}}$ defined in (32) employs a loading-dependent shrinkage strategy: it uses the plug-in estimator for components with large loadings ($j\leq j_{3}(s)$) and applies thresholding for components with small loadings ($j>j_{3}(s)$). Despite this similarity, the analysis differs substantially from the symmetric noise case. The key observation is that under symmetric noise, if $\theta_{j}=0$, the thresholding estimator is unbiased, i.e., $\mathbb{E}[\,y_{j}\mathbbm{1}\{|y_{j}|\geq\lambda\}\,]=0$ for any threshold $\lambda>0$. Under non-symmetric noise, however, this property does not hold in general. Consequently, we employ different techniques to analyze the non-symmetric case.
The upper bound established in Theorem 6 is nearly sharp. Section 5.1 shows an example where this upper bound can match the lower bounds in Theorem 5 in some regimes. In general, we have the following proposition for comparing this upper bound with the minimax rate characterized in Theorem 1.
Proposition 4.
For $\Phi_{\rm o}(s;\eta)$ defined in (8) and $j_{3}(s)$ defined in (31), we have
$\Phi_{\rm o}(s;\eta)\gtrsim\sum_{j\leq(s^{2}\wedge d)}\eta_{j}^{2}\;\gtrsim\;\frac{1}{\log^{2/\alpha}d}\sum_{j\leq j_{3}(s)}\eta_{j}^{2}.$
This proposition indicates that the cost of non-symmetry is at most a logarithmic factor. Furthermore, the symmetric lower bound $\sigma^{2}\Phi_{\mathrm{o}}(s;\eta)$ matches the upper bound in Theorem 3 up to logarithmic terms. In particular, in the dense regime where $s\gtrsim\sqrt{d}$, both the symmetric lower bound and the upper bound are of order $\sigma^{2}\sum_{j=1}^{d}\eta_{j}^{2}$, yielding the exact minimax rate. In this regime, the rate is independent of both the sparsity level $s$ and the symmetry of the noise distribution.
5 Example
We now present several examples to illustrate the rates derived above, covering both nonadaptive and adaptive estimation settings. Throughout this section, we set $\sigma=1$, and we use the notation $\Phi_{\rm o}(s;\eta)$ and $\Phi_{\rm adp}(s;\eta)$ to denote the optimal rates corresponding to minimax estimation with known $s$ and adaptive estimation, respectively. In addition, in the homogeneous loading vector example, we also examine the upper bound and asymmetric lower bound for the minimax rate $\Phi_{\rm ns}(s;\eta)$ under general (possibly non-symmetric) noise.
5.1 Homogeneous loading vector
Assume that $\eta_{1}=\cdots=\eta_{d}=1$. Up to multiplicative constants, the rate can be computed as follows:
$\Phi_{\rm o}(s;\mathbf{1}_{d})\asymp s^{2}\log^{2/\alpha}\!\left(1+\frac{d^{\alpha/2}}{s^{\alpha}}\right),\quad\Phi_{\rm adp}(s;\mathbf{1}_{d})\asymp s^{2}\log^{2/\alpha}\!\left(1+\frac{d^{\alpha/2}\log^{\alpha/2}(es)}{s^{\alpha}}\right)$ (34)
and
$\Phi_{\rm ns}(s;\mathbf{1}_{d})\asymp\left\{\begin{array}[]{cc}\displaystyle s^{2}\log^{2/\alpha}d,&\quad\displaystyle s\lesssim\frac{\sqrt{d}}{\log^{2/\alpha}d};\\ d,&\quad s\gtrsim\sqrt{d}.\end{array}\right$.
  • Adaptive rate: We can verify that Assumption 1 is satisfied (with $s_{\rm cut}\asymp d^{\gamma}$ for some $0<\gamma<1/2$) and $\gamma_{0}=1$, and thus our adaptive estimator is rate-optimal. Comparing $\Phi_{\rm o}(s;\mathbf{1}_{d})$ and $\Phi_{\rm adp}(s;\mathbf{1}_{d})$, the cost of adaptation is at most $\log d$, which is attained when $s\gtrsim\sqrt{d\log d}$.
  • Comparison with existing results: The homogeneous case has been studied previously under Gaussian noise, for both nonadaptive [9] and adaptive estimation [10]. For comparison, we set $\alpha=2$, and consider the sub-Gaussian noise class ${\cal G}_{2,\tau}$ for any $\tau>0$: the rates in Equation 34 show that the optimal rates remain the same as under Gaussian noise.
  • Comparison of symmetric and non-symmetric noise: For the non-symmetric noise setting, the asymmetric lower bound in Theorem 5 matches the minimax rate $\Phi_{\rm o}(s;\mathbf{1}_{d})$ (restricted to symmetric noise) when $s\lesssim\sqrt{d}/\log^{2/\alpha}d$. Therefore, our analysis is sharp for $s\lesssim\sqrt{d}/\log^{2/\alpha}d$ and $s\gtrsim\sqrt{d}$, but there remains a logarithmic gap between the upper and lower bounds in the narrow intermediate regime $\sqrt{d}/\log^{2/\alpha}d\lesssim s\lesssim\sqrt{d}$, which presents an interesting direction for future research. Comparing $\Phi_{\rm o}(s;\mathbf{1}_{d})$ and $\Phi_{\rm ns}(s;\mathbf{1}_{d})$, we observe that the absence of symmetry in the noise distribution increases the minimax rate from $s^{2}\log^{2/\alpha}(1+d^{\alpha/2}/s^{\alpha})$ to $s^{2}\log^{2/\alpha}d$, which differs only by a logarithmic factor.
5.2 Two-phase loading vector
We consider a loading vector $\eta$ of the form
$\eta_{j}=\begin{cases}\lambda,&j\leq d_{0},\\ 1,&j>d_{0}.\end{cases}$ (35)
This loading vector $\eta$ is divided into two parts: the first subvector consists of a small number of large components, while the second subvector consists of many small components. This structure is related to the homogeneous case, but it accommodates unbounded ratios of coordinates, a setting less studied in the literature.
For convenience, we parameterize $d_{0}$ and $\lambda$ as $d_{0}=d^{\gamma_{d}}$ and $\lambda=d^{\gamma_{\lambda}}$ for some constants $\gamma_{d}>0$ and $\gamma_{\lambda}>0$. To uncover an interesting phase transition phenomenon as discussed below, we assume $2\gamma_{\lambda}+\gamma_{d}<1$.
The minimax rate is given by
$\Phi_{\mathrm{o}}(s;\eta)$ $\asymp\begin{cases}d^{2\gamma_{\lambda}}s^{2}\log^{2/\alpha}\!\left(1+\dfrac{d^{\gamma_{d}\alpha/2}}{s^{\alpha}}\right),&s\lesssim\dfrac{d^{\gamma_{\lambda}+\gamma_{d}/2}}{\log^{1/\alpha}d},\\[6.0pt] s^{2}\log^{2/\alpha}\!\left(1+\dfrac{d^{\alpha/2}}{s^{\alpha}}\right),&s\gtrsim\dfrac{d^{\gamma_{\lambda}+\gamma_{d}/2}}{\log^{1/\alpha}d},\end{cases}$
while the adaptive rate is given by
$\Phi_{\mathrm{adp}}(s;\eta)$ $\asymp s^{2}\!\left[d^{2\gamma_{\lambda}}\log\!\left(1+\dfrac{d^{\gamma_{d}\alpha/2}\log^{\alpha/2}(es)}{s^{\alpha}}\right)+\log\!\left(1+\dfrac{d^{\alpha/2}\log^{\alpha/2}(es)}{s^{\alpha}}\right)\right].$
  • Phase transition in the minimax rate: Unlike the homogeneous case, here a phase transition arises in the minimax (nonadaptive) rate $\Phi_{\rm o}$. Specifically, when $s\lesssim d^{\gamma_{\lambda}+\gamma_{d}/2}/\log d$, the error of estimating the first subvector (with $\eta_{j}=\lambda$) dominates, and the minimax rate equals $\Phi_{\rm o}(s;\lambda\mathbf{1}_{d_{0}})$ in (34), which explicitly depends on $\gamma_{\lambda}$ and $\gamma_{d}$. In contrast, when $s\gtrsim d^{\gamma_{\lambda}+\gamma_{d}/2}/\log^{1/\alpha}d$, there are sufficiently many nonzero entries in the second subvector (with $\eta_{j}=1$) such that the overall error is dominated by the estimation error resulting from the second subvector, and the minimax rate equals $\Phi_{\rm o}(s;\mathbf{1}_{d})$ in (34), which is independent of $\gamma_{d}$ and $\gamma_{\lambda}$. Such a phase transition occurs if and only if $\gamma_{d}+2\gamma_{\lambda}<1$ (i.e., $\lambda^{2}d_{0}\ll d$); otherwise, the first subvector always dominates and the second subvector never contributes to the minimax rate.
  • Adaptive rate: We can verify that Assumption 1 is satisfied (with $s_{\rm cut}\asymp d^{\gamma}$ for some $\gamma\in(0,\gamma_{d}/2)$ and $\gamma_{0}=1$), and thus our adaptive estimator is rate-optimal.
5.3 Exponentially Decaying Loading Vector
Following [8], we consider the exponentially decaying loading vector: Let $\phi:[0,\infty)\to\mathbb{R}$ be a non-decreasing convex function such that $\phi(0)=0$, and let
$\eta_{j}=\exp\!\left(-\phi(j-1)\right),\qquad j\in\{1,\ldots,d\}.$
We allow the function $\phi$ to depend on $d$; for example, $\phi(x)=2(x/d)^{\gamma}$ for $\gamma\geq 1$. Given $\phi$, define $j_{0}=\min\{\,j\mid\eta_{j}<1/2\,\}$ if the set is nonempty; otherwise set $j_{0}=d+1$. The optimal rates are given by
$\Phi_{\rm o}(s;\eta)=s^{2}\log^{2/\alpha}\left(1+\frac{j_{0}}{s^{2}}\right)$
and
$\Phi_{\rm adp}(s;\eta)=\left\{\begin{array}[]{cc}\displaystyle s^{2}\log^{2/\alpha}\left(1+\frac{j_{0}\log(es)}{s^{2}}\right),&s\leq\sqrt{j_{0}\log(ej_{0})};\\ j_{0}\log(ej_{0}),&\text{otherwise}.\end{array}\right$.
  • If $j_{0}$ is bounded, then $\Phi_{\rm adp}(s;\eta)\asymp\Phi_{\rm o}(s;\eta)$ and therefore is optimal.
  • If $j_{0}$ goes to infinity, then we can verify 1 with $s_{0}=\sqrt{j_{0}}$ and $\gamma=1$.
Comparing the above results with the homogeneous case (34), we obtain
$\Phi_{\rm o}(s;\eta)\asymp\Phi_{\rm o}(s;\mathbf{1}_{j_{0}}),\qquad\Phi_{\rm adp}(s;\eta)\asymp\Phi_{\rm adp}(s;\mathbf{1}_{j_{0}}),$
which shows that an exponentially decaying loading vector behaves like a homogeneous vector with an effective dimension $j_{0}$.
6 Extensions
6.1 Estimation with unknown noise variance
The estimation problem with an unknown noise variance has been widely studied in the literature; see, for example, [9, 10, 11, 7]. A common approach is to first construct an estimator of $\sigma^{2}$ and then substitute it into the original procedure in place of the true noise variance. For example, [11] proposed a median-of-means estimator:
Let $\gamma_{\alpha,\tau}\in(0,1/2]$ be a constant chosen sufficiently small, depending only on $\alpha$ and $\tau$. Divide $\{1,\ldots,d\}$ into $m=\lfloor\gamma_{\alpha,\tau}d\rfloor$ disjoint subsets $B_{1},\ldots,B_{m}$, each of cardinality
$|B_{i}|\;\geq\;k:=\left\lfloor\frac{d}{m}\right\rfloor\;\geq\;\frac{1}{\gamma_{\alpha,\tau}}-1.$
The estimator of $\sigma^{2}$ is then defined as
$\hat{\sigma}^{2}=\operatorname{median}(\bar{\sigma}_{1}^{2},\ldots,\bar{\sigma}_{m}^{2}),\qquad\text{ where }\bar{\sigma}_{i}^{2}=\frac{1}{|B_{i}|}\sum_{j\in B_{i}}y_{j}^{2},\quad i=1,\ldots,m.$ (36)
[11] showed that $\hat{\sigma}^{2}$ is a consistent estimator of $\sigma^{2}$ under the assumption that the noise distribution belongs to ${\cal G}_{\alpha,\tau}$.
Based on the estimator $\hat{\sigma}^{2}$ in (36), most of our results for estimating the linear functional $L(\theta)$ extend naturally to the setting with an unknown noise variance. Due to space limitations, we restrict attention to the estimation problem with known sparsity. Replacing $\sigma$ in the estimator $\hat{L}_{s}$ defined in (7) with $\hat{\sigma}$ defined in (36) and adjusting the constant in front of the threshold, we obtain the estimator
$\hat{L}_{s}^{\prime}=\sum_{j\leq j_{1}(s)}\eta_{j}y_{j}+\sum_{j>j_{1}(s)}\eta_{j}y_{j}\,\mathbbm{1}\!\left\{|\eta_{j}y_{j}|>{~}\sqrt{2}\hat{\sigma}\tau\lambda_{o}\right\}.$
The next theorem provides an upper bound on its risk.
Theorem 7.
For any $\alpha,\tau>0$, there exist constants $\gamma_{\alpha,\tau}$ and $C$ depending only on $\alpha$ and $\tau$, such that if $1\leq s<\lfloor\gamma_{\alpha,\tau}d\rfloor/4$, then
$\sup_{\theta\in\Theta_{s}}\sup_{{\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}}\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\!\left(\hat{L}_{s}^{\prime}-L(\theta)\right)^{2}\;\leq\;C\sigma^{2}\Phi_{\rm o}(s;\eta).$
Together with Theorem 1, this theorem shows that the same minimax rate can be attained under an unknown noise level as in the case where the noise level is known, provided the sparsity level is not too large.
6.2 Linear hypothesis testing
The results on estimating the linear functional $L(\theta)$ can be applied to the problem of testing linear hypotheses over $\Theta_{s}$. For simplicity, we focus on the case of Gaussian noise, i.e., $\mathcal{P}_{\xi}={\cal N}^{\otimes}$. For any $t_{0}\in\mathbb{R}$, consider testing the null hypothesis
$H_{0}:\theta\in\Theta_{s,0}=\left\{\theta\in\Theta_{s}:L(\theta)=t_{0}\right\}$ (37)
against the alternative
$H_{1}:\theta\in\Theta_{s}(\rho)=\left\{\theta\in\Theta_{s}:|L(\theta)-t_{0}|\geq\rho\right\},$ (38)
for some $\rho>0$. Our analysis does not tie to the value of $t_{0}$.
Let $\Delta$ be a test with values in $\{0,1\}$. The risk of $\Delta$ is defined as the sum of the type I error and the maximal type II error:
$\sup_{\theta\in\Theta_{s,0}}\mathbb{P}_{\theta,{\mathcal{N}}^{\otimes}}(\Delta=1)\;+\;\sup_{\theta\in\Theta_{s}(\rho)}\mathbb{P}_{\theta,{\mathcal{N}}^{\otimes}}(\Delta=0).$
A benchmark quantity is the minimax risk of testing, defined as
$\mathcal{R}_{s}(\rho)=\inf_{\Delta}\left\{\sup_{\theta\in\Theta_{s,0}}\mathbb{P}_{\theta,{\mathcal{N}}^{\otimes}}(\Delta=1)\;+\;\sup_{\theta\in\Theta_{s}(\rho)}\mathbb{P}_{\theta,{\mathcal{N}}^{\otimes}}(\Delta=0)\right\},$
where the infimum is taken over all $\{0,1\}$-valued tests.
The minimax separation rate for testing $H_{0}:\theta\in\Theta_{s,0}$ versus $H_{1}:\theta\in\Theta_{s}(\rho)$ is the smallest $r_{\min}>0$ such that the following two properties hold:
  1. (i)
    For any $\varepsilon\in(0,1)$, there exists $A_{\varepsilon}>0$ such that, for all $A>A_{\varepsilon}$,
    $\mathcal{R}_{s}(Ar_{\min})\leq\varepsilon;$
  2. (ii)
    For any $\varepsilon\in(0,1)$, there exists $a_{\varepsilon}>0$ such that, for all $0<A<a_{\varepsilon}$,
    $\mathcal{R}_{s}(Ar_{\min})\geq 1-\varepsilon.$
The next theorem provides the exact expression for the minimax separation rate.
Theorem 8.
For any integers $s$ and $d$ such that $2\leq s\leq d$, and any $\sigma>0$, the minimax separation rate for testing $H_{0}:\theta\in\Theta_{s,0}$ versus $H_{1}:\theta\in\Theta_{s}(\rho)$ is
$r_{\min}=\sigma\sqrt{\Phi_{\rm o}(s;\eta)}.$
The proof of the lower bound in Theorem 8 follows arguments similar to those in Theorem 1; for completeness, the details are provided in Section C.3 in the supplementary material. For the upper bound, recall that the estimator $\hat{L}_{s}$ defined in (7) achieves the optimal squared error rate $\sigma^{2}\Phi_{\rm o}(s;\eta)$ over $\Theta_{s}$. Using this estimator, we construct the test
$\Delta_{s}=\mathbbm{1}\!\left\{\,|\hat{L}_{s}-t_{0}|>B\sigma\sqrt{\Phi_{\rm o}(s;\eta)}\,\right\},$ (39)
for some constant $B>0$, and we show in Section A.6 in the supplementary material that $\Delta_{s}$ achieves the minimax rate for testing with an appropriate choice of $B$. The proof also reveals that the results extend straightforwardly to the broader class of symmetric sub-Weibull noise distributions.
7 Future Directions
The sequence model in Equation 1 is closely related to the linear regression model, and our results yield new insights into the estimation of linear functionals in high-dimensional sparse regression. For high-dimensional linear regression, inference methods based on the idea of debiasing are well studied, but existing minimax theory relies on restrictive conditions on the loading vector $x_{0}$, such as bounded coordinate ratios or polynomial decay assumptions [2, 6]. Our work suggests these conditions may be relaxed since the sharp lower bounds developed here provide an improved benchmark for arbitrary loadings in orthogonal designs (see Example 2). Another related and promising direction is to consider correlated noise, which has attracted considerable recent attention [18, 19].
8 Proof of the lower bounds
In this section, we establish the lower bounds stated in Theorems 1, 4 and 5. Without loss of generality, we set $\sigma=1$ in the proofs of the lower bounds. For two probability measures $P,Q$ defined on the same measurable space $(\mathcal{X},\mathcal{U})$, we define:
$\mathrm{TV}(P,Q)\;=\;\sup_{B\in\mathcal{U}}\big|P(B)-Q(B)\big|,$
the total variation distance,
$\mathrm{H}(P,Q)\;=\;\sqrt{\int\!\left(\sqrt{\frac{{\rm d\,}P}{{\rm d\,}Q}}-1\right)^{\!2}\,{\rm d\,}Q},$
the Hellinger distance, and
$\chi^{2}(P\,\|\,Q)\;=\;\int\!\left(\frac{{\rm d\,}P}{{\rm d\,}Q}-1\right)^{\!2}{\rm d\,}Q,$
the chi-square divergence.
8.1 General Tools
For nonadaptive estimation, we prove the lower bounds using Le Cam’s method that reduces the estimation to a two-point hypothesis testing problem [22].
Let $\mu$ be a probability measure on $\mathbb{R}^{d}$. Denote by $\mathbb{P}_{\mu,\mathcal{P}_{\xi}}$ the mixture probability measure:
$\mathbb{P}_{\mu,\mathcal{P}_{\xi}}=\int_{\mathbb{R}^{d}}\mathbb{P}_{\theta,\mathcal{P}_{\xi}}\,\mu({\rm d\,}\theta).$
Specifically, we denote by $\mathbb{P}_{0,\mathcal{P}_{\xi}}$ the mixture probability measure when $\mu$ is the Dirac measure at 0, i.e., $\mathbb{P}_{0,\mathcal{P}_{\xi}}=\mathbb{P}_{\theta=0,\mathcal{P}_{\xi}}$. The following lemma, as a special case of [27, Theorem 2.15], is the key tool we will use to prove the lower bounds.
Lemma 1.
Suppose that there exist two distributions $P_{1}=\mathbb{P}_{\mu_{1},\mathcal{P}_{\xi}^{1}}$ and $P_{2}=\mathbb{P}_{\mu_{2},\mathcal{P}_{\xi}^{2}}$ such that
$\mu_{1}(L(\theta)\leq c,\theta\in\Theta)\geq 1-\beta_{1},\quad\mu_{2}(L(\theta)\geq c+2t,\theta\in\Theta)\geq 1-\beta_{2},\quad\mathrm{TV}(P_{1},P_{2})\leq\beta_{3}$
for some $c\in\mathbb{R}$, $t>0$, and $\beta_{1},\beta_{2},\beta_{3}\in(0,1)$, and some subset $\Theta\subseteq\mathbb{R}^{d}$. Then, for any estimator $\hat{T}$ of $L(\theta)$, we have
$\inf_{\hat{T}}\sup_{\theta\in\Theta}\max_{j\in\{1,2\}}\mathbb{P}_{\theta,\mathcal{P}_{\xi}^{j}}(|\hat{T}-L(\theta)|\geq t)\geq\frac{1-\beta_{1}-\beta_{2}-\beta_{3}}{2}.$
Since $\mathbb{E}(\hat{T}-L(\theta))^{2}\geq t^{2}\mathbb{P}(|\hat{T}-L(\theta)|\geq t)$, Lemma 1 implies a lower bound on the minimax mean squared error.
When bounding the total variation distance by the Hellinger distance or the $\chi^{2}$ divergence, the following lemmas are useful.
Lemma 2.
For $\alpha>0$, there exists a constant $\tau_{\alpha}>0$ and a distribution in ${\cal G}_{\alpha,\tau_{\alpha}}$ with density $f_{\alpha}^{(0)}$ such that, for $f_{\alpha}^{(1)}(\cdot):=f_{\alpha}^{(0)}(\cdot-\gamma)$ with some $\gamma\in\mathbb{R}$, it holds that
$1+\chi^{2}(f_{\alpha}^{(1)}\,\|\,f_{\alpha}^{(0)})\;\leq\;C_{\alpha,1}\exp\!\left(\left|\frac{\gamma}{C_{\alpha,2}}\right|^{\alpha}\right),$
where $C_{\alpha,1},C_{\alpha,2}$ are positive constants depending only on $\alpha$. For $\alpha\leq 1$ and $\alpha=2$, $C_{\alpha,1}=1$.
When $f_{0}$ and $f_{1}$ are Gaussian distributions, the corresponding $\chi^{2}$-divergence admits a closed-form expression, which has been widely used in the minimax theory of Gaussian models [9, 10, 2, 3, 6, 1, 24]. Lemma 2 extends this idea to the exponentially decaying noise setting, where no closed-form formula is available. Nevertheless, the optimality of our results highlights the effectiveness of Lemma 2.
Lemma 3 ([16, Theorem 7.6]).
Suppose $\mathcal{F}=\{p(\cdot;\theta):\theta\in\mathbb{R}^{p}\}$ is a family of density functions w.r.t. a $\sigma$-finite measure $\nu$ on $\mathbb{R}^{d}$ equipped with the Borel algebra $\mathcal{B}$. If the experiment $(\mathbb{R}^{d},\mathcal{B},\mathcal{F})$ is regular (see [16, Chapter 7.1] for a definition), then
$\int_{\mathbb{R}^{d}}\left[\sqrt{p(x;\theta)}-\sqrt{p(x;\theta+h)}\right]^{2}\nu({\rm d\,}x)\leq\frac{|h|^{2}}{4}\int_{0}^{1}{\rm tr}\left[I(\theta+sh)\right]{\rm d\,}s,\quad\forall\theta,h\in\mathbb{R}^{p}.$
8.2 Proof of the lower bounds in Theorem 1
Without loss of generality, assume $\sigma=1$. From Lemma 2, for any $\alpha>0$, there exists $\tau_{\alpha}>0$ and a distribution $f_{\alpha}^{0}$ such that $f_{\alpha}^{0}\in{\cal G}_{\alpha,\tau}$ for all $\tau\geq\tau_{\alpha}$. For simplicity, let $F_{\alpha}^{\otimes}$ denote the product distribution on $\mathbb{R}^{d}$ with i.i.d. marginals having density $f_{\alpha}^{0}$.
We apply Lemma 1 with $\Theta=\Theta_{s}$, taking $P_{1}=\mathbb{P}_{0,F_{\alpha}^{\otimes}}$ and $P_{2}=\mathbb{P}_{\mu,F_{\alpha}^{\otimes}}$, where $\mu$ is a distribution on $\mathbb{R}^{d}$ that will be specified later. In what follows, we establish the lower bound for the case where $s>C_{1}$ and $\lambda_{o}s+\nu>C_{2}|\eta_{1}|$ with constants $C_{1},C_{2}>0$ to be specified. The proof for the complementary case is provided in the supplementary material.
Define the probability measure $\mu$ on $\mathbb{R}^{d}$ as follows: for $\theta\sim\mu$, its coordinates are independent and
$\theta_{j}=b_{j}\gamma_{j},\quad\forall j\in[d],$
where $b_{j}={\rm Ber}(\pi_{j})$,
$\pi_{j}=c_{1}\cdot\frac{|\eta_{j}|\exp(-\beta_{+}/|\eta_{j}|^{\alpha})}{\sqrt{\sum_{i=1}^{d}\eta_{i}^{2}\exp(-\beta_{+}/|\eta_{i}|^{\alpha})}},\quad\text{and}\quad\gamma_{j}=\left\{\begin{array}[]{lc}C_{\alpha,2}{\rm sign}(\eta_{j})&j\leq j_{1}(s)\\ C_{\alpha,2}{\lambda_{o}}/{\eta_{j}}&j>j_{1}(s)\end{array}\right$.
for some constant $c_{1}>0$ to be specified later, where the constant $C_{\alpha,2}$ is as given in Lemma 2. It can be easily verified that $\pi_{j}\in(0,1)$. From Equation 5, we have
$\mathbb{E}_{\mu}\left\|\theta\right\|_{0}=\sum_{j=1}^{d}\pi_{j}\leq\frac{c_{1}}{2}s,\quad\text{and}\quad{\rm Var}_{\mu}\left\|\theta\right\|_{0}\leq\sum_{j=1}^{d}\pi_{j}(1-\pi_{j})\leq\sum_{j=1}^{d}\pi_{j}\leq\frac{c_{1}}{2}s.$
Since $\eta_{j}\gamma_{j}=C_{\alpha,2}\max(|\eta_{j}|,\lambda_{o})$ is non-increasing in $j$, we have
${\rm Var}_{\mu}[L(\theta)]=\sum_{j=1}^{d}\eta_{j}^{2}\gamma_{j}^{2}\cdot\pi_{j}[1-\pi_{j}]\leq\eta_{1}\gamma_{1}\cdot\sum_{j=1}^{d}\eta_{j}\gamma_{j}\pi_{j}=C_{\alpha,2}\max(|\eta_{1}|,\lambda_{o})\cdot\mathbb{E}_{\mu}L(\theta).$
Using $\eta_{j}\gamma_{j}\geq C_{\alpha,2}|\eta_{j}|$, we have $\mathbb{E}_{\mu}L(\theta)=\sum_{j=1}^{d}\eta_{j}\gamma_{j}\pi_{j}\geq C_{\alpha,2}\sum_{j=1}^{d}|\eta_{j}|\pi_{j}=c_{1}C_{\alpha,2}\nu$. Using $\eta_{j}\gamma_{j}\geq C_{\alpha,2}\lambda_{o}$, we have $\mathbb{E}_{\mu}L(\theta)=\sum_{j=1}^{d}\eta_{j}\gamma_{j}\pi_{j}\geq C_{\alpha,2}\lambda_{o}\sum_{j=1}^{d}\pi_{j}$. If $\lambda_{o}>0$, we have $\sum_{j=1}^{d}\pi_{j}=\frac{1}{2}c_{1}s$ and thus $\mathbb{E}_{\mu}L(\theta)\geq\frac{1}{2}c_{1}C_{\alpha,2}\lambda_{o}s$. This inequality also holds when $\lambda_{o}=0$.
Consequently, we have
$\mathbb{E}_{\mu}L(\theta)\geq\frac{c_{1}}{4}C_{\alpha,2}[\lambda_{o}s+\nu]>0.$
By Chebyshev’s inequality, we have, for any $c_{1}\in(0,2)$,
$\mu\left(\|\theta\|_{0}>s\right)\leq\mu\left(\|\theta\|_{0}-\mathbb{E}_{\mu}\|\theta\|_{0}>(1-\frac{c_{1}}{2})s\right)\leq\frac{{\rm Var}_{\mu}[\left\|\theta\right\|_{0}]}{(1-c_{1}/2)^{2}s^{2}}\leq\frac{c_{1}/2}{(1-c_{1}/2)^{2}s},$
and
$\mu\left(L(\theta)<\frac{c_{1}}{8}C_{\alpha,2}[\lambda_{o}s+\nu]\right)$ $\leq\mu\left(L(\theta)<\frac{1}{2}\mathbb{E}_{\mu}L(\theta)\right)$
$\leq\frac{4{\rm Var}_{\mu}L(\theta)}{\left[\mathbb{E}_{\mu}[L(\theta)]\right]^{2}}$
$\leq\frac{16\max(|\eta_{1}|,\lambda_{o})}{c_{1}(\lambda_{o}s+\nu)}.$
Therefore, $\forall c_{1}\in(0,2),\alpha_{1}\in(0,1)$, we can choose $C_{1},C_{2}$ large enough such that
$\forall s\geq C_{1},\lambda_{o}s+\nu\geq C_{2}|\eta_{1}|,\quad\mathbb{P}\left(\theta\in\Theta_{s},L(\theta)\geq\frac{c_{1}}{8}C_{\alpha,2}[\lambda_{o}s+\nu]\right)\geq 1-\alpha_{1}.$
Let $P_{2}=\mathbb{P}_{\mu,F_{\alpha}^{\otimes}}$, and denote by $f_{\alpha}^{j}$ the distribution of the $j$-th component under $P_{2}$. By Lemma 2, we have
$1+\chi^{2}(P_{2}\|P_{1})$ $=\prod_{j=1}^{d}\left[1+\chi^{2}\left(f_{\alpha}^{j}\|f_{\alpha}^{(0)}\right)\right]$
$\stackrel{{\scriptstyle(1)}}{{=}}\prod_{j=1}^{d}\left[1+\pi_{j}^{2}\cdot\chi^{2}(f_{\alpha}^{(0)}(\cdot-\gamma_{j})\|f_{\alpha}^{(0)})\right]$
$\stackrel{{\scriptstyle(2)}}{{\leq}}\exp\left[\sum_{j=1}^{d}\pi_{j}^{2}\chi^{2}(f_{\alpha}^{(0)}(\cdot-\gamma_{j})\|f_{\alpha}^{(0)})\right]$
$\stackrel{{\scriptstyle(3)}}{{\leq}}\exp\left[c_{1}^{2}C_{\alpha,1}\frac{e\sum_{j\leq j_{1}(s)}\eta_{j}^{2}\exp(-2\beta_{+}/|\eta_{j}|^{\alpha})+\sum_{j>j_{1}(s)}\eta_{j}^{2}\exp(-\beta_{+}/|\eta_{j}|^{\alpha})}{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta_{+}/|\eta_{j}|^{\alpha})}\right]$
$\leq\exp(c_{1}^{2}C_{\alpha,1}e),$
where we have applied Fubini’s theorem in (1), $1+x\leq e^{x}$ in (2), and Lemma 2 in (3). Therefore, for any $\alpha_{2}\in(0,1-\alpha_{1})$, we can choose some small $c_{1}$ such that
$\mathrm{TV}(P_{1},P_{2})\leq\sqrt{\chi^{2}(P_{2}\|P_{1})}/2\leq\alpha_{2}.$
Specifically, we can take $\alpha_{1}=\alpha_{2}=1/4$ and fix the constants $c_{1}$, $C_{1}$, and $C_{2}$. Using Lemma 1 with $t=2^{-3}C_{\alpha,2}(\lambda_{o}s+\nu)$, we have
$\inf_{\hat{L}_{s}}\sup_{\theta\in\Theta_{s}}\sup_{{\mathcal{P}}_{\xi}\in\mathcal{G}_{\alpha,\tau}^{\otimes}}\mathbb{P}_{\theta,{\mathcal{P}}_{\xi}}\left(\!|\hat{L}_{s}-L(\theta)|\;\geq\;2^{-3}C_{\alpha,2}(\lambda_{o}s+\nu)\right)\geq\frac{1}{4}.$
8.3 Proof of Theorem 4
The lower bound for the adaptive estimation requires the following new technical tool.
Lemma 4.
Let $P,Q_{1}$ and $Q_{2}$ be three probability measures defined over the same measurable space $(\mathcal{X},\mathcal{U})$. Then for any $q>0$, we have
$\inf_{{\cal A}\in{\cal U}}\left\{P({\cal A})q+Q_{1}({\cal A}^{c})\right\}\geq\max_{0<\tau<1}\left[\frac{q\tau}{1+q\tau}\left(1-\tau(\chi^{2}(Q_{2}||P)+1)\right)-{\rm TV}(Q_{1},Q_{2})\right].$ (40)
Lemma 4 follows directly from [10, Lemma 8] since for any $A\in\mathcal{U}$, we have $P({\cal A})q+Q_{1}({\cal A}^{c})\geq P({\cal A})q+Q_{2}({\cal A}^{c})-{\rm TV}(Q_{1},Q_{2})$.
Proof.
Theorem 1 has shown that $\Phi_{\rm o}(s;\eta)$ provides a lower bound on the minimax risk. Hence, for any fixed $\gamma\in(0,2)$, we assume that $s\leq s_{0}$ and that $\Phi_{\rm adp}(s;\eta)\;\geq\;C\,s^{\gamma}\,\Phi_{\rm o}(1;\eta)$ for some sufficiently large constant $C>0$. The case where $s>s_{0}$ follows similarly because ${\Phi}_{\rm adp}(s;\eta)=s_{0}$. Together with Proposition 2, it then suffices to restrict attention to the regime
$s\;\geq\;C_{1},\qquad\sqrt{\log(es)}\nu_{*}(s)+s\lambda_{*}(s)\;\geq\;C_{2}|\eta_{1}|^{2},$
for some large constants $C_{1},C_{2}>0$.
Define the probability measure $\mu$ on $\mathbb{R}^{d}$ as follows: for $\theta\sim\mu$, its coordinates are independent and
$\theta_{j}=b_{j}\gamma_{j},\quad\forall j\in[d],$
where $b_{j}={\rm Ber}(\pi_{j})$,
$\pi_{j}=\frac{c_{1}|\eta_{j}|\exp(-(\beta_{*})_{+}/|\eta_{j}|^{\alpha})}{\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-(\beta_{*})_{+}/|\eta_{j}|^{\alpha})}}\cdot\sqrt{\log(es)},\quad\text{and}\quad\gamma_{j}=\left\{\begin{array}[]{cc}C_{\alpha,2}{\rm sign}(\eta_{j})&j\leq j_{2}(s)\\ C_{\alpha,2}\lambda_{*}(s)/\eta_{j}&j>j_{2}(s)\end{array}\right$.
for some constant $c_{1}>0$ specified later. The construction of the measure is similar to that in Section 8.2. By following the same line of argument, we see that for any $c_{1}\in(0,2)$, one can choose $C_{1},C_{2}>0$ sufficiently large such that
$\mu\left(\theta\in\Theta_{s},L(\theta)\geq\frac{c_{1}C_{\alpha,2}}{8}[\lambda_{*}(s)s+\sqrt{\log es}\cdot\nu_{*}(s)]\right)\geq\frac{3}{4}$
for all $s\geq C_{1}$ and $\Phi_{*}(s;\eta)\geq C_{2}|\eta_{1}|$.
We now apply Lemma 4 to establish the lower bound in Theorem 4. Recall that $F_{\alpha}^{\otimes}$ denotes the product distribution introduced in Section 8.2. Let $P=\mathbb{P}_{0,F_{\alpha}^{\otimes}}$, $Q_{1}=\mathbb{P}_{\mu_{1},F_{\alpha}^{\otimes}}$ and $Q_{2}=\mathbb{P}_{\mu,F_{\alpha}^{\otimes}}$, where the $\mu_{1}$ is a restricted version of $\mu$ defined by
$\mu_{1}(A)=\frac{\mu(A\cap A_{0})}{\mu(A_{0})}$
with
$A_{0}=\left\{\theta\in\mathbb{R}^{d}:\theta\in\Theta_{s},L(\theta)\geq\frac{c_{1}C_{\alpha,2}}{8}[\lambda_{*}(s)s+\sqrt{\log es}\nu_{*}(s)]\right\}.$
Similar to the calculations in Section 8.2, we have
$1+\chi^{2}\left(Q_{2}\|P_{1}\right)$ $=\prod_{j=1}^{d}\left[1+\pi_{j}^{2}\cdot\chi^{2}\left(f_{\alpha}^{(0)}\left(\cdot-\gamma_{j}\right)\|f_{\alpha}^{(0)}\right)\right]$
$\leq\exp\left[\sum_{j=1}^{d}\pi_{j}^{2}\chi^{2}\left(f_{\alpha}^{(0)}\left(\cdot-\gamma_{j}\right)\|f_{\alpha}^{(0)}\right)\right]$
$\leq\exp\left[c_{1}^{2}C_{\alpha,1}e\log(es)\right]$
Note that
${\rm TV}(Q_{1},Q_{2})\leq{\rm TV}(\mu,\mu_{1})\leq\mu(\theta\notin A_{0})\leq\frac{1}{4}.$
Since $s\geq C_{1}$, for $\gamma\in(0,2)$ we choose $c_{1}$ small enough and $C_{1}$ larger enough such that $\chi^{2}(Q_{2}||P)\leq s^{\gamma}-1$. Then we can apply Lemma 4 with $\tau=s^{-\gamma}/2$ and $q=4s^{\gamma}$ to obtain
$\inf_{{\cal A\in U}}\left[P({\cal A})q+Q_{1}({\cal A}^{c})\right]\geq\frac{2}{3}\cdot\frac{1}{2}-\frac{1}{4}>\frac{1}{12}.$
Define the event
${\cal A}=\left\{\hat{T}\;\geq\;\frac{c_{1}C_{\alpha,2}}{16}\Bigl(\lambda_{*}(s)s+\sqrt{\log(es)}\,\nu_{*}(s)\Bigr)\right\}.$
For any estimator $\hat{T}$ satisfying
$\sup_{\theta\in\Theta_{s}}\sup_{{\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}}\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\bigl(\hat{T}-L(\theta)\bigr)^{2}\;\leq\;\frac{1}{C_{0}}\Phi_{\rm adp}(s;\eta),$
we have
$Q_{1}({\cal A}^{c})\;\leq\;\frac{16^{2}}{c_{1}^{2}C_{\alpha,2}^{2}C_{0}}\cdot\frac{\Phi_{\rm adp}(s;\eta)}{\bigl(\lambda_{*}(s)s+\sqrt{\log(es)}\,\nu_{*}(s)\bigr)^{2}}.$
Since $\Phi_{\rm adp}(s;\eta)\asymp\bigl(\lambda_{*}(s)s+\sqrt{\log(es)}\,\nu_{*}(s)\bigr)^{2}$, for any $c_{1}>0$, we can choose $C_{0}$ large enough such that $Q_{1}({\cal A}^{c})\leq 1/24$. Hence,
$P({\cal A})\;\geq\;\frac{1}{24q}\;\;\Rightarrow\;\;\sup_{\theta\in\Theta_{1}}\sup_{{\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}}\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\bigl(\hat{T}-L(\theta)\bigr)^{2}\;\geq\;\frac{c_{1}^{2}C_{\alpha,2}^{2}}{16^{2}\cdot 96}\cdot\frac{\bigl(\lambda_{*}(s)s+\sqrt{\log(es)}\,\nu_{*}(s)\bigr)^{2}}{s^{\gamma}}.$
The proof is completed since $\Phi_{\rm adp}(s;\eta)\asymp\bigl(\lambda_{*}(s)s+\sqrt{\log(es)}\,\nu_{*}(s)\bigr)^{2}$.
Supplementary material
The Supplementary Material contains the detailed proofs of all auxiliary lemmas as well as details of the proofs of Theorems.
Supplement to “Minimax and adaptive estimation of general linear functionals under sparsity”
Appendix A provides detailed proofs of all upper bound results in the paper. Appendix B provides the proof for the uniqueness of the solution to Equation 5 and includes the proofs of the propositions on the rate characteristics. Appendix C includes the proofs of Theorem 5 and the lemmas used in proving the lower bounds Section 8. Appendix D contains the details for the examples in Section 5.
In the appendix, we will denote by $C_{i},i=1,2,\ldots$, absolute positive constants, and by $C,c$ absolute positive constants that may vary from line to line.
Appendix A Proof of the upper bounds
We will use the following lemmas.
Lemma 5.
For $X\sim P$ for some $P\in{\cal G}_{\alpha,\tau}$ (or ${\cal H}_{\alpha,\tau}$), we have
$\mathbb{E}X^{2q}\leq C_{q}^{*},q\in\mathbb{N},$ (41)
where $C_{q}^{*}$ are positive constant depending on $\alpha$ and $\tau$.
Lemma 5 can be easily proved by using the tail condition in Definition 1 and integration by parts.
Lemma 6.
For $\alpha,\tau>0$, we can choose $\zeta$ large enough, then for the $\hat{s}$ defined in (14), we have
$\sup_{\theta\in\Theta_{s}}\sup_{\mathcal{P}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}}\mathbb{P}_{\theta,\mathcal{P}_{\xi}}\left(\hat{s}>s\right)\leq c_{3,0}s^{-7},\qquad\forall s\leq s_{*},$ (42)
where $c_{3,0}$ is a constant that only depend on $\alpha$ and $\tau$.
Lemma 7.
For $\alpha,\tau>0$, there exists some constant $c_{4,0}>0$ such that
$\sup_{\theta\in\Theta_{s}}\sup_{{\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}}\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left(\hat{L}_{s}^{*}-L(\theta)\right)^{4}\leq c_{4,0}\Phi_{\rm adp}^{2}(1;\eta)s^{4}\log^{2}(es),\quad\forall s\leq s_{0}.$
Lemma 8 ([11, Proposition 1]).
For any $\alpha,\tau>0$, there exist constants $\gamma_{\alpha,\tau}\in(0,1/2]$, $c_{\alpha,\tau}^{*}>0$, and $C_{\alpha,\tau}^{*}>0$ depending only on $\alpha$ and $\tau$ such that, for any integers $s$ and $d$ satisfying $1\leq s<\lfloor\gamma d\rfloor/4$ and $\hat{\sigma}$ defined in (36), we have
$\inf_{\theta\in\Theta_{s}}\inf_{{\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}}\mathbb{P}_{\theta,\mathcal{P}_{\xi}}\!\left(\frac{1}{2}\leq\frac{\hat{\sigma}^{2}}{\sigma^{2}}\leq\frac{3}{2}\right)\;\geq\;1-\exp(-c_{\alpha,\tau}^{*}d),$
and
$\sup_{\theta\in\Theta_{s}}\sup_{{\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}}\frac{\mathbb{E}_{\theta,\mathcal{P}_{\xi}}\!\bigl|\hat{\sigma}^{2}-\sigma^{2}\bigr|}{\sigma^{2}}\;\leq\;C_{\alpha,\tau}^{*}.$
In this section, we will use the notation $S=\left\{j:\theta_{j}\neq 0\right\}$ for the true parameter vector $\theta$.
A.1 Proof of the upper bound in Theorem 1
Proof.
Note that $y_{j}=\theta_{j}+\sigma\xi_{j}$, where $\xi_{j}$ are mean-zero, unit-variance random variables. If $\lambda_{o}=0$, then no thresholding is applied in $\hat{L}_{s}$, and
$\hat{L}_{s}-L(\theta)=\sum_{j=1}^{d}\eta_{j}\xi_{j}.$
In this case, it is immediate that
$\forall\,\theta\in\Theta_{s},\;{\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes},\quad\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\big(\hat{L}_{s}-L(\theta)\big)^{2}=\sigma^{2}\sum_{j=1}^{d}\eta_{j}^{2}.$
Hence, in the sequel, we restrict attention to the case where $\lambda_{o}>0$. For all ${\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}$ and $\theta\in\Theta_{s}$, we have
[…]
$+\underbrace{\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\!\left(\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S\end{subarray}}\eta_{j}y_{j}\mathbbm{1}\!\left\{|\eta_{j}y_{j}|\geq{~}\sigma\tau\lambda_{o}\right\}-\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S\end{subarray}}\eta_{j}\theta_{j}\right)^{2}}_{\text{denoted as III}}\Biggr].$
Term I.
$\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\!\left(\sum_{j\leq j_{1}(s)}\eta_{j}y_{j}-\sum_{j\leq j_{1}(s)}\eta_{j}\theta_{j}\right)^{2}=\sigma^{2}\sum_{j\leq j_{1}(s)}\eta_{j}^{2}.$
Term II.
$\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\!\left(\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S^{c}\end{subarray}}\sigma\eta_{j}\xi_{j}\mathbbm{1}\!\left\{|\eta_{j}\xi_{j}|\geq{~}\tau\lambda_{o}\right\}\right)^{2}$ $=\sigma^{2}\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S^{c}\end{subarray}}\mathbb{E}\!\eta_{j}^{2}\xi_{j}^{2}\mathbbm{1}\!\left\{|\xi_{j}|\geq\frac{{~}\tau\lambda_{o}}{|\eta_{j}|}\right\}$
$\stackrel{{\scriptstyle(*)}}{{\leq}}\sigma^{2}\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S^{c}\end{subarray}}\sqrt{\mathbb{E}(\eta_{j}^{4}\xi_{j}^{4})}\cdot\sqrt{\mathbb{P}\!\left(|\xi_{j}|\geq\frac{{~}\tau\lambda_{o}}{|\eta_{j}|}\right)}$
$\stackrel{{\scriptstyle(**)}}{{\leq}}2\sigma^{2}\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S^{c}\end{subarray}}\eta_{j}^{2}\sqrt{C_{2}^{*}}\cdot\exp(-\beta/|\eta_{j}|^{\alpha})$
$\leq 2\sqrt{C_{2}^{*}}\sigma^{2}\nu^{2},$
where inequality ()(*) follows from Hölder’s inequality, inequality ()(**) from Lemma 5 and Definition 1, and the last inequality is due to the definition of $\nu$ in Equation 8.
Term III.
$\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\!\left(\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S\end{subarray}}\eta_{j}y_{j}\mathbbm{1}\!\left\{|\eta_{j}y_{j}|\geq{~}\sigma\tau\lambda_{o}\right\}-\sum_{j>j_{1}(s)}\eta_{j}\theta_{j}\right)^{2}$
$=\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\!\left(\sigma\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S\end{subarray}}\eta_{j}\xi_{j}-\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S\end{subarray}}\eta_{j}y_{j}\mathbbm{1}\!\left\{|\eta_{j}y_{j}|<{~}\sigma\tau\lambda_{o}\right\}\right)^{2}$
$\leq 2s\cdot\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S\end{subarray}}\left[\sigma^{2}\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}(\eta_{j}\xi_{j})^{2}+\mathbb{E}\!\left(\eta_{j}y_{j}\mathbbm{1}\!\left\{|\eta_{j}y_{j}|<{~}\sigma\tau\lambda_{o}\right\}\right)^{2}\right]$
$\leq 2s\cdot\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S\end{subarray}}\left[\sigma^{2}\eta_{j}^{2}+{~}{~}\sigma^{2}\lambda_{o}^{2}\tau^{2}\right]\leq 2(1+{~}{~}\tau^{2})\sigma^{2}\lambda_{o}^{2}s^{2}.$
Combining the bounds for I, II, and III, we obtain the desired result (10) in Theorem 1. ∎
A.2 Proof of Theorem 2
Proof.
Let $\theta\in\Theta_{s},{\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}$. Similar to the proof of Theorem 1, we can assume $\lambda_{*}(s)>0$ and decompose the error as follows:
$\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}(\hat{L}_{s}^{*}-L(\theta))^{2}\leq$ $3\biggl[\underbrace{\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\biggl(\sum_{j\leq j_{2}(s)}\eta_{j}y_{j}-\sum_{j\leq j_{2}(s)}\eta_{j}\theta_{j}\biggr)^{2}}_{\text{denoted as I}}$
[…]
$\qquad\qquad+\underbrace{\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\biggl(\sum_{\begin{subarray}{c}j>j_{2}(s)\\ j\in S\end{subarray}}\eta_{j}y_{j}\mathbbm{1}\left\{|\eta_{j}y_{j}|\geq{~}\sigma\tau\lambda_{*}(s)\right\}-\sum_{\begin{subarray}{c}j>j_{2}(s)\\ j\in S\end{subarray}}\eta_{j}\theta_{j}\biggr)^{2}}_{\text{denoted as III}}\biggr].$
Term I.
$\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left(\sum_{j\leq j_{2}(s)}\eta_{j}y_{j}-\sum_{j\leq j_{1}}\eta_{j}\theta_{j}\right)^{2}=\sigma^{2}\sum_{j\leq j_{2}(s)}\eta_{j}^{2}\lesssim\sigma^{2}\sum_{j=1}^{d}\eta_{j}^{2}\exp\left(-\frac{\beta_{+}}{|\eta_{j}|^{\alpha}}\right).$
Term II.
Note that the symmetry of $\xi_{j}$ implies that $\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\bigl(\eta_{j}\xi_{j}\mathbbm{1}\left\{|\eta_{j}\xi_{j}|\geq{~}\tau\lambda_{*}(s)\right\}\bigr)=0$. Using the independence across indices, we have
$\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\biggl(\sum_{\begin{subarray}{c}j>j_{2}(s)\\ j\in S^{c}\end{subarray}}\sigma\eta_{j}\xi_{j}\mathbbm{1}\left\{|\eta_{j}\xi_{j}|\geq{~}\tau\lambda_{*}(s)\right\}\biggr)^{2}$ $=\sigma^{2}\sum_{\begin{subarray}{c}j>j_{2}(s)\\ j\in S^{c}\end{subarray}}\mathbb{E}\biggl(\eta_{j}^{2}\xi_{j}^{2}\mathbbm{1}\left\{|\xi_{j}|\geq\tau\lambda_{*}(s)/|\eta_{j}|\right\}\biggr)$
$\leq\sigma^{2}\sum_{\begin{subarray}{c}j>j_{2}(s)\\ j\in S^{c}\end{subarray}}\sqrt{\mathbb{E}(\eta_{j}^{4}\xi_{j}^{4})}\sqrt{\mathbb{P}(|\xi_{j}|\geq{~}\tau\lambda_{*}(s)/|\eta_{j}|)}$
$\leq 2\sqrt{C_{2}^{*}}\sigma^{2}\sum_{j=1}^{d}\eta_{j}^{2}\exp\left(-\frac{(\beta_{*})_{+}}{|\eta_{j}|^{\alpha}}\right).$
Term III.
$\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\biggl(\sum_{\begin{subarray}{c}j>j_{2}(s)\\ j\in S\end{subarray}}\eta_{j}y_{j}\mathbbm{1}\left\{|\eta_{j}y_{j}|\geq{~}\sigma\tau\lambda_{*}(s)\right\}-\sum_{\begin{subarray}{c}j>j_{2}(s)\\ j\in S\end{subarray}}\eta_{j}\theta_{j}\biggr)^{2}$
$=$ $\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\biggl(\sigma\sum_{\begin{subarray}{c}j>j_{2}(s)\\ j\in S\end{subarray}}\eta_{j}\xi_{j}-\sum_{\begin{subarray}{c}j>j_{2}(s)\\ j\in S\end{subarray}}\eta_{j}y_{j}\mathbbm{1}\left\{|\eta_{j}y_{j}|<{~}\sigma\tau\lambda_{*}(s)\right\}\biggr)^{2}$
$\leq$ $2s\cdot\sum_{\begin{subarray}{c}j>j_{2}(s)\\ j\in S\end{subarray}}\left[\sigma^{2}\mathbb{E}(\eta_{j}\xi_{j})^{2}+\mathbb{E}(\eta_{j}y_{j}\mathbbm{1}\left\{|\eta_{j}y_{j}|<{~}\sigma\tau\lambda_{*}(s)\right\})^{2}\right]$
$\leq$ $2s\cdot\sum_{\begin{subarray}{c}j>j_{2}(s)\\ j\in S\end{subarray}}\left[\sigma^{2}\eta_{j}^{2}+{~}{~}\sigma^{2}\tau^{2}\lambda^{2}_{*}(s)\right]\leq 2[1+{~}{~}\tau^{2}]\sigma^{2}\lambda^{2}_{*}(s)s^{2}.$
Combining the bounds for I, II, and III, we obtain the desired result in Theorem 2. ∎
A.3 Proof of Theorem 3
Proof.
For $\theta\in\Theta_{s},{\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}$, we have
$\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}(\tilde{L}_{*}-L(\theta))^{2}=\underbrace{\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left[(\tilde{L}_{*}-L(\theta))^{2}\mathbbm{1}\left\{\hat{s}\leq s\right\}\right]}_{\text{ denoted as I}}+\underbrace{\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left[(\tilde{L}_{*}-L(\theta))^{2}\mathbbm{1}\left\{\hat{s}>s\right\}\right]}_{\text{ denoted as II}}.$
Term I. Recall that $s_{0}=s_{*}+1$. By definition of $\hat{s}$, we have $\hat{s}\leq s_{0}$ and on the event $\{\hat{s}\leq s\}$,
[…]
Thus,
$\forall s<s_{0}:\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left[(\hat{L}_{\hat{s}}^{*}-L(\theta))^{2}\mathbbm{1}_{\hat{s}\leq s}\right]$ $\leq 2\left[\zeta\sigma^{2}\Phi_{\rm adp}(s;\eta)+\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left(\hat{L}_{s}^{*}-L(\theta)\right)^{2}\right],$ (44)
$\forall s\geq s_{0}:\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left[(\hat{L}_{\hat{s}}^{*}-L(\theta))^{2}\mathbbm{1}_{\hat{s}\leq s}\right]$ $\leq\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left[(\hat{L}_{\hat{s}}^{*}-L(\theta))^{2}\left(\mathbbm{1}_{\hat{s}\leq s,\hat{s}<s_{0}}+\mathbbm{1}_{\hat{s}=s_{0}}\right)\right]$
$\leq 2\left[\zeta\sigma^{2}\Phi_{\rm adp}(s;\eta)+\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left(\hat{L}_{s}^{*}-L(\theta)\right)^{2}\right]$ (45)
$\qquad+\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left(\hat{L}_{s_{0}}^{*}-L(\theta)\right)^{2}.$
For $s=1,\ldots,s_{0}-1$, by Theorem 2, we have
$\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left(\hat{L}_{s}^{*}-L(\theta)\right)^{2}\leq c_{1,0}\sigma^{2}\Phi_{\rm adp}(s;\eta)$
for some absolute constant $c_{1}>0$.
For $s\in[s_{0},d]$, $\lambda_{*}(s)=0$ and there is no thresholding in the estimator $\hat{L}_{s}^{*}$, and we have
$\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left(\hat{L}_{s}^{*}-L(\theta)\right)^{2}=\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left(\hat{L}_{s_{0}}^{*}-L(\theta)\right)^{2}=\sigma^{2}\sum_{j=1}^{d}\eta_{j}^{2}\leq\sigma^{2}\Phi_{\rm adp}(s;\eta).$
Combining with (44) and (45), we have
$\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left[\left(\hat{L}_{s}^{*}-L(\theta)\right)^{2}\mathbbm{1}\left\{\hat{s}\leq s\right\}\right]\lesssim\sigma^{2}\Phi_{\rm adp}(s;\eta),\qquad s=1,\ldots,d.$
Term II. Since by definition $\hat{s}\leq s_{0}$, we have $s<\hat{s}\leq s_{0}$ when the indicator inside the expectation equal 1. In this case, Lemmas 6 and 7 can be applied to obtain the following for any $\theta\in\Theta_{s}$:
$\sup_{\theta\in\Theta_{s}}\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left[(\hat{L}_{\hat{s}}-L(\theta))^{2}\mathbbm{1}_{\hat{s}>s}\right]$ (46)
$=$ $\sup_{\theta\in\Theta_{s}}\sum_{s<s^{\prime}\leq s_{0}}\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left[(\hat{L}_{s^{\prime}}-L(\theta))^{2}\mathbbm{1}\left\{\hat{s}=s^{\prime}\right\}\right]$
$\leq$ $\sum_{s<s^{\prime}\leq s_{0}}\sup_{\theta\in\Theta_{s}}\sqrt{\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}(\hat{L}_{s^{\prime}}-L(\theta))^{4}}\cdot\sup_{\theta\in\Theta_{s}}\sqrt{\mathbb{P}_{\theta,{\mathcal{P}}_{\xi}}\left(\hat{s}=s^{\prime}\right)}$
$\leq$ $\sum_{s<s^{\prime}\leq s_{0}}\sup_{\theta\in\Theta_{s^{\prime}}}\sup_{{\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}}\sqrt{\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}(\hat{L}_{s^{\prime}}-L(\theta))^{4}}\sup_{\theta\in\Theta_{s^{\prime}-1}}\sup_{{\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}}\sqrt{\mathbb{P}_{\theta,{\mathcal{P}}_{\xi}}\left(\hat{s}>s^{\prime}-1\right)}$
$\leq$ $\sqrt{c_{3,0}\cdot c_{4,0}}\cdot\Phi_{\rm adp}(1;\eta)\sum_{s<s^{\prime}\leq s_{0}}(s^{\prime})^{(4-7)/2}\log(es^{\prime}),$
where the first inequality is due to Hölder’s inequality, and we have applied Lemmas 6 and 7 in the last line. Therefore, we have $\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left[(\hat{L}_{\hat{s}}-L(\theta))^{2}\mathbbm{1}_{\hat{s}>s}\right]\leq c\cdot\Phi_{\rm adp}(1;\eta)$ for some constant $c>0$. The proof is completed since we have $\Phi_{\rm adp}(1;\eta)\leq c_{2,0}\Phi_{\rm adp}(s;\eta)$ and $\sum_{i=1}^{\infty}i^{-3/2}\log(ei)<\infty$. ∎
A.4 Proof of Theorem 6
Proof.
Let $\theta\in\Theta_{s}$ and ${\mathcal{P}}_{\xi}\in{\cal H}_{\alpha,\tau}^{\otimes}.$
If $j_{3}(s)=d$, then no thresholding is applied in the estimator $\hat{L}_{\cal H}$, and $\xi_{j}$ are independent noise with mean zero and unit variance. Clearly, we have
$\mathbb{E}_{\theta,{\mathcal{P}_{\xi}}}(\hat{L}_{\cal H}-L(\theta))^{2}=\mathbb{E}_{\theta,{\mathcal{P}_{\xi}}}\biggl(\sum_{j=1}^{d}\eta_{j}\xi_{j}\biggr)^{2}=\sigma^{2}\sum_{j=1}^{d}\eta_{j}^{2},$
and the desired result is proved.
When $j_{3}(s)<d$, we have $s\leq\sqrt{d}$. Let $\varepsilon=(\varepsilon_{1},\ldots,\varepsilon_{d})\in\mathbb{R}^{d}$ with
$\varepsilon_{j}=y_{j}\mathbbm{1}\left\{|y_{j}|\geq c_{\cal H}\cdot\sigma\cdot\log^{1/\alpha}\left(\frac{ed}{s}\right)\right\}-\theta_{j}.$
Similar to (43), we have
$\mathbb{E}_{\theta,{\mathcal{P}_{\xi}}}(\hat{L}_{\cal H}-L(\theta))^{2}\leq 2\biggl[\underbrace{\mathbb{E}_{\theta,{\mathcal{P}_{\xi}}}\biggl(\sum_{j\leq j_{3}(s)}\eta_{j}y_{j}-\sum_{j\leq j_{3}(s)}\eta_{j}\theta_{j}\biggr)^{2}}_{\text{denoted as I}}+\underbrace{\mathbb{E}_{\theta,{\mathcal{P}_{\xi}}}\biggl(\sum_{j>j_{3}(s)}\eta_{j}\varepsilon_{j}\biggr)^{2}}_{\text{denoted as II}}\biggr].$ (47)
Term I.
$\mathbb{E}_{\theta,{\mathcal{P}_{\xi}}}\biggl(\sum_{j\leq j_{3}(s)}\eta_{j}y_{j}-\sum_{j\leq j_{3}(s)}\eta_{j}\theta_{j}\biggr)^{2}=\sigma^{2}\sum_{j\leq j_{3}(s)}\eta_{j}^{2}.$
Term II.
$\mathbb{E}_{\theta,{\mathcal{P}_{\xi}}}\biggl(\sum_{j>j_{3}(s)}\eta_{j}\varepsilon_{j}\biggr)^{2}\leq|\eta_{j_{3}(s)+1}|^{2}\cdot\mathbb{E}_{\theta,{\mathcal{P}_{\xi}}}\|\varepsilon\|_{1}^{2}.$
For $j\in S^{c}$, we have
$\varepsilon_{j}$ $=\sigma\xi_{j}\mathbbm{1}\left\{|\xi_{j}|\geq c_{\cal H}\cdot\log^{1/\alpha}\left(\frac{ed}{s}\right)\right\}.$
Since
$\mathbb{P}(|\xi_{j}|\geq t)\leq 2\exp\left\{-\left(\frac{t}{\tau}\right)^{\alpha}\right\}\qquad\text{for any }t\geq 0,$
we have
$\mathbb{E}\varepsilon_{j}^{2}$ $=\sigma^{2}\left(\mathbb{E}\xi_{j}^{2}\mathbbm{1}\left\{|\xi_{j}|\geq c_{\cal H}\log^{1/\alpha}\left(\frac{ed}{s}\right)\right\}\right)$
$\leq\sigma^{2}\sqrt{\mathbb{E}\xi_{j}^{4}}\cdot\sqrt{\mathbb{P}\left(|\xi_{j}|\geq c_{\cal H}\cdot\log^{1/\alpha}\left(\frac{ed}{s}\right)\right)}$
$\leq\sigma^{2}\sqrt{2C_{2}^{*}\left(\frac{s}{ed}\right)^{(c_{\cal H}/{\tau})^{\alpha}}},$
where the last inequality comes from Lemma 5. Let $C_{1}=\sqrt{2C_{2}^{*}}/e^{2}$ and $c_{\cal H}=\tau 4^{1/\alpha}$. We have
$\mathbb{E}_{\theta,\mathcal{P}_{\xi}}\left\|\varepsilon_{S^{c}}\right\|_{1}^{2}$ $\leq|S^{c}|\biggl(\sum_{j\in S^{c}}\mathbb{E}\varepsilon_{j}^{2}\biggr)$
$\leq\sigma^{2}d^{2}\sqrt{2C_{2}^{*}\left(\frac{s}{ed}\right)^{4}}$
$\leq C_{1}\sigma^{2}s^{2}.$
For $j\in S$, we have
$|\varepsilon_{j}|=\left|\sigma\xi_{j}-y_{j}\mathbbm{1}\left\{|y_{j}|<c_{\cal H}\cdot\sigma\cdot\log^{1/\alpha}\left(\frac{ed}{s}\right)\right\}\right|\leq\sigma|\xi_{j}|+c_{\cal H}\cdot\sigma\cdot\log^{1/\alpha}\left(\frac{ed}{s}\right).$
The first term on the right hand side has a bounded second moment, and the number of nonzero coordinates is at most $s$. Therefore, we have $\mathbb{E}_{\theta,\mathcal{P}_{\xi}}\left\|\varepsilon_{S}\right\|_{1}^{2}\leq C_{2}\sigma^{2}s^{2}\log^{2/\alpha}\left(ed/s\right)$ for some constant $C_{2}>0$ depending on $\alpha$ and $\tau$.
Combining the above analysis, we have
$\mathbb{E}_{\theta,\mathcal{P}_{\xi}}\left(\hat{L}_{\cal H}-L(\theta)\right)^{2}\leq\sigma^{2}\left[\sum_{j\leq j_{3}(s)}\eta_{j}^{2}+2(C_{1}+C_{2})|\eta_{j_{3}(s)+1}|^{2}s^{2}\log^{2/\alpha}\left(\frac{ed}{s}\right)\right].$
Since $j_{3}(s)=s^{2}\log^{2/\alpha}(ed/s)<d$ and $|\eta_{j}|$ is decreasing, we have
$|\eta_{j_{3}(s)+1}|^{2}s^{2}\log^{2/\alpha}\left(\frac{ed}{s}\right)\leq\sum_{j\leq j_{3}(s)}\eta_{j}^{2}$
and therefore, the proof of Theorem 6 is completed.
A.5 Proof of Theorem 7
Proof.
Let $\theta\in\Theta_{s}$ and ${\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}$. As before, we only need to consider the case where $\lambda_{\rm o}>0$. Similar to (43), we decompose the risk as
[…]
$\quad+\underbrace{\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\!\left(\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S\end{subarray}}\eta_{j}y_{j}\,\mathbbm{1}\!\left\{|\eta_{j}y_{j}|\geq{~}\sqrt{2}\hat{\sigma}\tau\lambda_{o}\right\}-\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S\end{subarray}}\eta_{j}\theta_{j}\right)^{2}}_{\text{denoted as III}}\Biggr].$
Term I.
$\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\!\left(\sum_{j\leq j_{1}(s)}\eta_{j}y_{j}-\sum_{j\leq j_{1}(s)}\eta_{j}\theta_{j}\right)^{2}=\sigma^{2}\sum_{j\leq j_{1}(s)}\eta_{j}^{2}\;\lesssim\;\sigma^{2}\Phi_{\rm o}(s;\eta).$
Term II.
For $j\in S^{c}$, the random variables $\xi_{j}$ remain symmetric after conditioning on $\hat{\sigma}$. Hence,
$\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\!\left(\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S^{c}\end{subarray}}\sigma\eta_{j}\xi_{j}\,\mathbbm{1}\!\left\{\sigma|\eta_{j}\xi_{j}|\geq{~}\sqrt{2}\hat{\sigma}\tau\lambda_{o}\right\}\right)^{2}$
$=\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S^{c}\end{subarray}}\mathbb{E}\!\left[\sigma^{2}\eta_{j}^{2}\xi_{j}^{2}\,\mathbbm{1}\!\left\{\sigma|\eta_{j}\xi_{j}|\geq{~}\sqrt{2}\hat{\sigma}\tau\lambda_{o}\right\}\right]$
$=\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S^{c}\end{subarray}}\mathbb{E}\!\left[\sigma^{2}\eta_{j}^{2}\xi_{j}^{2}\,\mathbbm{1}\!\left\{\sigma|\eta_{j}\xi_{j}|\geq{~}\sqrt{2}\hat{\sigma}\tau\lambda_{o},\,\hat{\sigma}<\frac{\sigma}{\sqrt{2}}\right\}\right]$
$\quad+\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S^{c}\end{subarray}}\mathbb{E}\!\left[\sigma^{2}\eta_{j}^{2}\xi_{j}^{2}\,\mathbbm{1}\!\left\{\sigma|\eta_{j}\xi_{j}|\geq{~}\sqrt{2}\hat{\sigma}\tau\lambda_{o},\,\hat{\sigma}>\frac{\sigma}{\sqrt{2}}\right\}\right].$
For the first term, we apply Hölder’s inequality and Lemma 8 to obtain
$\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S^{c}\end{subarray}}\sigma^{2}\mathbb{E}\!\left[\eta_{j}^{2}\xi_{j}^{2}\,\mathbbm{1}\!\left\{\hat{\sigma}<\frac{\sigma}{\sqrt{2}}\right\}\right]$ $\leq\sigma^{2}\sum_{j=1}^{d}\sqrt{\mathbb{E}(\eta_{j}^{4}\xi_{j}^{4})\cdot\mathbb{P}(\hat{\sigma}<\frac{\sigma}{\sqrt{2}})}$
$\leq\sigma^{2}\exp\!\left(-\frac{c_{\alpha,\tau}^{*}}{2}d\right)\sum_{j=1}^{d}\eta_{j}^{2}$
$\leq\sigma^{2}d\exp\!\left(-\frac{c_{\alpha,\tau}^{*}}{2}d\right)\eta_{1}^{2}$
$\lesssim\;\sigma^{2}\Phi_{\rm o}(s;\eta),$
where the last inequality is because $d\exp\!\left(-\tfrac{c_{\alpha,\tau}^{*}}{2}d\right)\lesssim 1$ and $\eta_{1}^{2}\lesssim\Phi_{\rm o}(s;\eta)$ by the definition of $\Phi_{\rm o}(s;\eta)$ (either $\eta_{1}^{2}\leq\lambda_{o}^{2}$ or $\lambda_{o}(s)/|\eta_{1}|\leq 1$, which implies that $\nu^{2}\geq\eta_{1}^{2}e^{-1}$).
For the second term, we can bound as in the proof of Theorem 1:
$\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S^{c}\end{subarray}}\mathbb{E}\!\left[\sigma^{2}\eta_{j}^{2}\xi_{j}^{2}\,\mathbbm{1}\!\left\{|\eta_{j}\xi_{j}|\geq{~}\tau\lambda_{o}\right\}\right]$ $\leq\sigma^{2}\sum_{j=1}^{d}\eta_{j}^{2}\exp\!\left(-\frac{\lambda_{\rm o}^{\alpha}}{|\eta_{j}|^{\alpha}}\right)\lesssim\sigma^{2}\Phi_{\rm o}(s;\eta).$
Term III. Lemma 8 implies that $\mathbb{E}\hat{\sigma}^{2}\leq\sigma^{2}(1+C_{\alpha,\tau}^{*})$. We have
$\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\!\left(\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S\end{subarray}}\eta_{j}y_{j}\,\mathbbm{1}\!\left\{|\eta_{j}y_{j}|\geq{~}\sqrt{2}\hat{\sigma}\tau\lambda_{o}\right\}-\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S\end{subarray}}\eta_{j}\theta_{j}\right)^{2}$
$=$ $\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\!\left(\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S\end{subarray}}\eta_{j}y_{j}\,\mathbbm{1}\!\left\{|\eta_{j}y_{j}|<{~}\sqrt{2}\hat{\sigma}\tau\lambda_{o}\right\}-\sum_{\begin{subarray}{c}j>j_{1}(s)\\ j\in S\end{subarray}}\sigma\eta_{j}\xi_{j}\right)^{2}$
$\leq$ $2s\sum_{j>j_{1}(s)}\mathbb{E}\!\left[\eta_{j}^{2}y_{j}^{2}\,\mathbbm{1}\!\left\{|\eta_{j}y_{j}|<{~}\sqrt{2}\hat{\sigma}\tau\lambda_{o}\right\}+\sigma^{2}\eta_{j}^{2}\xi_{j}^{2}\right]$
$\leq$ $2s\sum_{j>j_{1}(s)}\left({~}{~}2\tau^{2}\lambda_{o}^{2}\mathbb{E}\hat{\sigma}^{2}+\sigma^{2}\eta_{j}^{2}\right)$
$\leq$ ${~}{~}2^{2}s^{2}\tau^{2}\lambda_{o}^{2}\sigma^{2}(1+C_{\alpha,\tau}^{*})+2\sigma^{2}\sum_{j>j_{1}(s)}\eta_{j}^{2}$
$\lesssim$ $\sigma^{2}\Phi_{\rm o}(s;\eta).$
Combining the bounds for I, II, and III, we obtain the desired result in Theorem 7. ∎
A.6 Proof of the upper bound in Theorem 8
Note that for the test $\Delta_{s}$ defined in (39), we have
$\forall\theta\in\Theta_{s,0},\,\mathbb{P}_{\theta,{\mathcal{N}}^{\otimes}}{(\Delta=1)=\mathbb{P}_{\theta,{\mathcal{N}}^{\otimes}}}(|\hat{L}_{s}^{*}-t_{0}|\geq B\sigma\sqrt{\Phi_{\rm o}(s;\eta)})\leq\frac{\mathbb{E}_{\theta,{\mathcal{N}}^{\otimes}}(\hat{L}_{s}-L(\theta))^{2}}{B^{2}\sigma^{2}\Phi_{\rm o}(s;\eta)}.$
For any $A>B$ and any $\theta\in\Theta_{s}(A\sigma\sqrt{\Phi_{\rm o}(s;\eta)})$, we have $|L(\theta)-t_{0}|\geq A\sigma\sqrt{\Phi_{\rm o}(s;\eta)}$. Therefore,
$\mathbb{P}_{\theta,{\mathcal{N}}^{\otimes}}(\Delta=0)$ $=\mathbb{P}_{\theta,{\mathcal{N}}^{\otimes}}(|\hat{L}_{s}^{*}-t_{0}|<B\sigma\sqrt{\Phi_{\rm o}(s;\eta)})$
$\leq\mathbb{P}_{\theta,{\mathcal{N}}^{\otimes}}(|\hat{L}_{s}^{*}-L(\theta)|>(A-B)\sigma\sqrt{\Phi_{\rm o}(s;\eta)})$
$\leq\frac{\mathbb{E}_{\theta,{\mathcal{N}}^{\otimes}}(\hat{L}_{s}-L(\theta))^{2}}{(A-B)^{2}\sigma^{2}\Phi_{\rm o}(s;\eta)}.$
By Theorem 1, there is some constant $C>0$ such that
$\sup_{\theta\in\Theta_{s}}\mathbb{E}_{\theta,{\mathcal{N}}^{\otimes}}(\hat{L}_{s}-L(\theta))^{2}\leq C\sigma^{2}\Phi_{\rm o}(s;\eta).$
For any $\varepsilon>0$, we can choose $B$ and $A_{\varepsilon}$ large enough such that $C/B^{2}<\varepsilon/2$ and $C/(A-B)^{2}<\varepsilon/2$, which implies that
$\forall A\geq A_{\varepsilon},\,\sup_{\theta\in\Theta_{s,0}}\mathbb{P}_{\theta,{\mathcal{N}}^{\otimes}}(\Delta_{s}=1)\;+\;\sup_{\theta\in\Theta_{s}(A\sigma\sqrt{\Phi_{\rm o}(s;\eta)})}\mathbb{P}_{\theta,{\mathcal{N}}^{\otimes}}(\Delta_{s}=0)\leq\varepsilon.$
This completes the proof.
A.7 Proofs of the Lemmas
Proof of Lemma 6.
Let $\theta\in\Theta_{s}$ and ${\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}$. Below, we drop the subscript in $\mathbb{P}_{\theta,\mathcal{P}_{\xi}}$.
By definition of the estimator $\hat{L}_{s}^{*}$ in (12), we have the following for any pairs of $s,s^{\prime}$ such that $s<s^{\prime}\leq s_{0}$:
$\hat{L}_{s^{\prime}}-\hat{L}_{s}^{*}=$ $\sum_{\begin{subarray}{c}j_{2}(s)<j\leq j_{2}(s^{\prime})\\ j\in S\end{subarray}}\eta_{j}y_{j}\mathbbm{1}\left\{|\eta_{j}y_{j}|\leq{~}\sigma\tau\lambda_{*}(s)\right\}$
$\quad+\sum_{\begin{subarray}{c}j>j_{2}(s^{\prime})\\ j\in S\end{subarray}}\eta_{j}y_{j}\mathbbm{1}\left\{{~}\sigma\tau\lambda_{*}(s^{\prime})<|\eta_{j}y_{j}|\leq{~}\sigma\tau\lambda_{*}(s)\right\}$
$\quad\quad+\sigma\sum_{\begin{subarray}{c}j_{2}(s)<j\leq j_{2}(s^{\prime})\\ j\notin S\end{subarray}}\eta_{j}\xi_{j}\mathbbm{1}\left\{|\eta_{j}\xi_{j}|\leq{~}\tau\lambda_{*}(s)\right\}$
$\quad\quad\quad+\sigma\sum_{\begin{subarray}{c}j>j_{2}(s^{\prime})\\ j\notin S\end{subarray}}\eta_{j}\xi_{j}\mathbbm{1}\left\{{~}\tau\lambda_{*}(s^{\prime})<|\eta_{j}\xi_{j}|\leq{~}\tau\lambda_{*}(s)\right\}.$
The sum of the first two terms on the right hand side is bounded by
$\tau\sigma s\lambda_{*}(s)\leq{~}\tau\sigma\sqrt{\Phi_{\rm adp}(s;\eta)}\leq 2\sqrt{c_{2,0}}\tau\sigma\sqrt{\Phi_{\rm adp}(s^{\prime};\eta)},$
where the second inequality is due to Proposition 2. Therefore, once $\zeta>9\cdot 2^{2/\alpha}c_{2,0}\tau^{2}$, we have
$\biggl|$ $\sum_{\begin{subarray}{c}j_{2}(s)<j\leq j_{2}(s^{\prime})\\ j\in S\end{subarray}}\eta_{j}y_{j}\mathbbm{1}\left\{|\eta_{j}y_{j}|\leq{~}\sigma\tau\lambda_{*}(s)\right\}$
$\quad+\sum_{\begin{subarray}{c}j>j_{2}(s^{\prime})\\ j\in S\end{subarray}}\eta_{j}y_{j}\mathbbm{1}\left\{\sigma\tau\lambda_{*}(s^{\prime})<|\eta_{j}y_{j}|\leq{~}\sigma\tau\lambda_{*}(s)\right\}\biggr|\leq\frac{1}{3}\omega_{s^{\prime}},$
which implies that
$\mathbb{P}_{\theta}\left(|\hat{L}_{s^{\prime}}-\hat{L}_{s}^{*}|\geq\omega_{s^{\prime}}\right)$
$\leq$ $\underbrace{\mathbb{P}\biggl(\biggl|\sum_{\begin{subarray}{c}j_{2}(s)<j\leq j_{2}(s^{\prime})\\ j\notin S\end{subarray}}\eta_{j}\xi_{j}\mathbbm{1}\left\{|\eta_{j}\xi_{j}|\leq{~}\tau\lambda_{*}(s)\right\}\biggr|\geq\frac{\omega_{s^{\prime}}}{3\sigma}\biggr)}_{\text{denoted as I}}$
$\,+\underbrace{\mathbb{P}_{\theta}\biggl(\biggl|\sum_{\begin{subarray}{c}j>j_{2}(s^{\prime})\\ j\notin S\end{subarray}}\eta_{j}\xi_{j}\mathbbm{1}\left\{{~}\tau\lambda_{*}(s^{\prime})<|\eta_{j}\xi_{j}|\leq{~}\tau\lambda_{*}(s)\right\}\biggr|\geq\frac{\omega_{s^{\prime}}}{3\sigma}\biggr)}_{\text{denoted as II}}.$
Term I. The random variables $\eta_{j}\xi_{j}\mathbbm{1}\left\{|\eta_{j}\xi_{j}|\leq{~}\tau\lambda_{*}(s)\right\}$ are independent, mean zero and and bounded by $K={~}\tau\lambda_{*}(s)$. Let $U_{1}=\{j_{2}(s)<j\leq j_{2}(s^{\prime}):j\notin S\}$. From the Bernstein’s inequality for bounded distributions [28, Theorem 2.8.4], we have
$\mathbb{E}\left\{\biggl|\sum_{j\in U_{1}}\eta_{j}\xi_{j}\mathbbm{1}\left\{|\eta_{j}\xi_{j}|\leq\tau\lambda_{*}(s)\right\}\biggr|\geq t\right\}\leq 2\exp\left(-\frac{t^{2}}{\tilde{\sigma}^{2}+Kt/3}\right),\forall t>0,$
where
$\tilde{\sigma}^{2}=\sum_{j\in U_{1}}\mathbb{E}\eta_{j}^{2}\xi_{j}^{2}\mathbbm{1}\left\{|\eta_{j}\xi_{j}|\leq{~}\tau\lambda_{*}(s)\right\}\leq\sum_{j\leq j_{2}(s^{\prime})}\eta_{j}^{2}.$
For $t=\omega_{s^{\prime}}/(3\sigma)$, we have
$t^{2}=\frac{\zeta}{9}\Phi_{\rm adp}(s^{\prime};\eta)\;\geq\;\frac{\log(es^{\prime})}{9e}\sum_{j\leq j_{2}(s^{\prime})}\eta_{j}^{2},$
and
$\frac{t}{K}\;\geq\;\frac{\sqrt{\zeta}\,\sqrt{\Phi_{\rm adp}(1;\eta)}\,\log(es^{\prime})}{{~}\tau\lambda_{*}(s)}.$
Since $\sqrt{\Phi_{\rm adp}(1;\eta)}\gtrsim\lambda_{*}(1)\gtrsim\lambda_{*}(s)$, we may choose $\zeta$ sufficiently large such that
$\mathbb{P}\!\left(\left|\sum_{\begin{subarray}{c}j_{2}(s)<j\leq j_{2}(s^{\prime})\\ j\notin S\end{subarray}}\eta_{j}\xi_{j}\mathbbm{1}\{|\eta_{j}\xi_{j}|\leq{~}\tau\lambda_{*}(s)\}\right|\;\geq\;\frac{\sqrt{\zeta}}{3}\sqrt{\Phi_{\rm adp}(s^{\prime};\eta)}\right)\;\leq\;2(s^{\prime})^{-8}.$
Moreover, when $\alpha\geq 2$, the truncated variables $\xi_{j}\mathbbm{1}\{|\eta_{j}\xi_{j}|\leq\tau\lambda_{*}(s)\}$ are sub-Gaussian, and thus Hoeffding’s inequality [28, Theorem 2.6.3] yields the same probability bound.
Term II. Let $U_{2}=\{j>j_{2}(s^{\prime}):j\notin S\}$. For $j\in U_{2}$, we write $\xi_{j}=\varepsilon_{j}\left|\xi_{j}\right|$ where $\varepsilon_{j}$ denotes the sign of $\xi_{j}$. Let
$p_{0}$ $=\mathbb{P}\left(\left|\sum_{j\in U_{2}}\eta_{j}\xi_{j}\mathbbm{1}\left\{|\eta_{j}\xi_{j}|\geq{~}\tau\lambda_{*}(s^{\prime})\right\}\right|\geq\frac{\omega_{s^{\prime}}}{3\sigma}\right)$
$=\mathbb{E}\left[\mathbb{P}\biggl(\biggl|\sum_{j\in U_{2}}\eta_{j}\varepsilon_{j}|\xi_{j}|\mathbbm{1}\left\{|\eta_{j}\xi_{j}|\geq{~}\tau\lambda_{*}(s^{\prime})\right\}\biggr|\geq\frac{\omega_{s^{\prime}}}{3\sigma}\;\bigg||\xi_{i}|,i\in U_{2}\biggr)\right],$
where we have used the law of total expectation. Consider the function
$g(x)=\biggl|\sum_{j\in U_{2}}\eta_{j}x_{j}|\xi_{j}|\mathbbm{1}\left\{|\eta_{j}\xi_{j}|\geq{~}\tau\lambda_{*}(s^{\prime})\right\}\biggr|,$
where $x=(x_{j},j\in U)$ with $x_{j}\in\{-1,1\}$. For any $i_{0}\in U_{2}$, let $g_{i_{0},u}(x)$ denote the function value of $g(x)$ but with $x_{i_{0}}$ replaced by $u$ for $u\in\{-1,1\}$. Note that for any fixed $(|\eta_{j}|,j\in U_{2})$, we have the bounded difference condition:
$\sup_{x}|g(x)-g_{i_{0},u}(x)|\leq 2\cdot|\eta_{i_{0}}\xi_{i_{0}}|\mathbbm{1}\left\{|\eta_{i_{0}}\xi_{i_{0}}|\geq{~}\tau\lambda_{*}(s^{\prime})\right\}\stackrel{{\scriptstyle\triangle}}{{=}}2\cdot Z_{i_{0}}\quad\forall u\in\{-1,1\},i_{0}\in U_{2}.$
Conditional on $(|\xi_{j}|,j\in U_{2})$, we apply the bounded difference inequality to obtain
$\mathbb{P}\biggl(\biggl|\sum_{j\in U_{2}}\eta_{j}\varepsilon_{j}|\xi_{j}|\mathbbm{1}\left\{|\eta_{j}\xi_{j}|\geq{~}\tau\lambda_{*}(s^{\prime})\right\}\biggr|\geq\frac{\omega_{s^{\prime}}}{3\sigma}\;\bigg||\xi_{i}|,i\in U_{2}\biggr)\leq 2\exp\left(-\frac{\omega_{s^{\prime}}^{2}}{18\sigma^{2}\sum_{j\in U_{2}}Z_{j}^{2}}\right).$
Therefore,
$p_{0}\leq 2\mathbb{E}\left[\exp\left(-\frac{\zeta\Phi_{\rm adp}(s^{\prime};\eta)}{18\sum_{j\in U_{2}}Z_{j}^{2}}\right)\right]\leq 2\exp\left(-\frac{\zeta\Phi_{\rm adp}(s^{\prime};\eta)}{18\Delta}\right)+\mathbb{P}(\sum_{j\in U_{2}}Z_{i}^{2}>\Delta)\quad$ (48)
for any $\Delta>0$. We can choose $\Delta=\zeta\Phi_{\rm adp}(s^{\prime};\eta)/(144\log(es^{\prime}))$ such that the first term on the right hand side of (48) is bounded by $2(s^{\prime})^{-8}$. Now for the second term, we will use the Fuk-Nagaev inequality [25, page 78], which we state here for the reader’s convenience.
Lemma 9 (Fuk-Nagaev inequality).
Let $p>2$ and $\nu>0$. Assume that $X_{1},\ldots,X_{n}$ are independent random variables with $\mathbb{E}(X_{i})=0$ and $\mathbb{E}|X_{i}|^{p}<\infty$, $i=1,\ldots,n$. Then
$\mathbb{P}\left(\sum_{i=1}^{n}X_{i}>\nu\right)\leq(1+2/p)^{p}\sum_{i=1}^{n}\mathbb{E}|X_{i}|^{p}\nu^{-p}+\exp\left(-\frac{2\nu^{2}}{(p+2)^{2}e^{p}\sum_{i=1}^{n}\mathbb{E}X_{i}^{2}}\right).$
For $j\in U_{2}$, define $X_{j}=Z_{j}^{2}-\mathbb{E}Z_{j}^{2}$. In view of Lemma 5, the random variables $X_{i}$ satisfy that
$\mathbb{E}|X_{i}|^{p}\leq 2^{p}\mathbb{E}|Z_{i}|^{2p}$ $=2^{3p}\mathbb{E}\left(|\eta_{i}\xi_{i}|^{2p}\mathbbm{1}\left\{|\xi_{i}|\geq\frac{{~}\tau\lambda_{*}(s^{\prime})}{|\eta_{i}|}\right\}\right)$
$\leq 2^{3p}|\eta_{j}|^{2p}\sqrt{\mathbb{E}|\xi_{i}|^{4p}\mathbb{P}\left(\mathbbm{1}\left\{|\xi_{i}|\geq\frac{{~}\tau\lambda_{*}(s^{\prime})}{|\eta_{i}|}\right\}\right)}$
$\leq 2^{3p}|\eta_{j}|^{2p}\sqrt{C_{2p}^{*}}\exp\left(-\frac{(\beta_{*})_{+}}{|\eta_{i}|^{\alpha}}\right).$
Since $j\in U_{2}$, we have $|\eta_{j}|\leq\lambda_{*}(s^{\prime})$. It follows that
$\sum_{j\in U_{2}}\mathbb{E}|X_{i}|^{p}\leq$ $2^{3p}\sqrt{C_{2p}^{*}}\cdot[\lambda_{*}(s^{\prime})]^{2p-2}\sum_{j\in U_{2}}\eta_{j}^{2}\exp\left(-\frac{(\beta_{*})_{+}}{|\eta_{i}|^{\alpha}}\right)$
$\leq$ $2^{3p}\sqrt{C_{2p}^{*}}\cdot[\lambda_{*}(s^{\prime})]^{2p-2}\sum_{j=1}^{d}\eta_{j}^{2}\exp\left(-\frac{(\beta_{*})_{+}}{|\eta_{i}|^{\alpha}}\right)$
$\leq$ $2^{3p}\sqrt{C_{2p}^{*}}\cdot(s^{\prime})^{-(2p-2)}(\log(es^{\prime}))^{-1}\left[\Phi_{\rm adp}(s^{\prime};\eta)\right]^{p}$
In particular, we have
$\sum_{j\in U_{2}}\mathbb{E}X_{j}^{2}\leq$ $64\sqrt{C_{4}^{*}}\cdot\lambda_{*}^{2}(s^{\prime})\sum_{j=1}^{d}|\eta_{j}|^{2}\exp\left(-\frac{(\beta_{*})_{+}}{|\eta_{i}|^{\alpha}}\right)$
$\leq$ $64\sqrt{C_{4}^{*}}(s^{\prime})^{-2}(\log(es^{\prime}))^{-1}\left[\Phi_{\rm adp}(s^{\prime};\eta)\right]^{2}.$
Similarly, we have
$\sum_{j\in U_{2}}\mathbb{E}Z_{j}^{2}\leq\sqrt{C_{2}^{*}}\sum_{j=1}^{d}|\eta_{j}|^{2}\exp\left(-\frac{(\beta_{*})_{+}}{|\eta_{i}|^{\alpha}}\right).$
Note that
$\Delta=\frac{\zeta\Phi_{\rm adp}(s^{\prime};\eta)}{144\log(es^{\prime})}\geq\frac{\zeta}{144}\left[\sum_{j=1}^{d}\eta_{j}^{2}\exp\left(-\frac{(\beta_{*})_{+}}{|\eta_{i}|^{\alpha}}\right)+\left(\frac{s^{\prime}}{\sqrt{\log(es^{\prime})}}\right)^{2}\lambda_{*}^{2}(s^{\prime})\right].$
Therefore, we can choose $\zeta$ large enough (i.e., $\zeta\geq 288\sqrt{C_{2}^{*}}$) such that
$\nu:=\Delta-\sum_{j\in U_{2}}\mathbb{E}Z_{j}^{2}\geq\Delta/2\geq\frac{\zeta}{288}\frac{\Phi_{\rm adp}(s^{\prime};\eta)}{\log(es^{\prime})}$
We apply Lemma 9 with $p=6$ to obtain
$\mathbb{P}(\sum_{j\in U_{2}}X_{j}>\nu)\leq$ $\biggl(\frac{288}{\zeta}\cdot(1+\frac{2}{6})2^{3}\biggr)^{6}\sqrt{C_{12}^{*}}\left[\frac{\Phi_{\rm adp}(s^{\prime};\eta)}{\log(es^{\prime})}\right]^{-6}(s^{\prime})^{10}(\log(es^{\prime}))^{-1}\left[\Phi_{\rm adp}(s^{\prime};\eta)\right]^{6}$
$\qquad+\exp\left(-\frac{2(\frac{\zeta}{288}\frac{\Phi_{\rm adp}(s^{\prime};\eta)}{\log(es^{\prime})})^{2}}{8^{2}e^{6}\cdot 64\sqrt{C_{4}^{*}}(s^{\prime})^{-2}(\log(es^{\prime}))^{-1}\left[\Phi_{\rm adp}(s^{\prime};\eta)\right]^{2}}\right)$
$\leq$ $C\zeta^{-6}(\log(es^{\prime}))^{5}(s^{\prime})^{-10}+\exp\left(-C^{\prime}\zeta^{2}(s^{\prime})^{2}/\log(es^{\prime})\right),$
where $C$ and $C^{\prime}$ are absolute constants. If we choose $\zeta$ large enough (for example, $\zeta^{2}C^{\prime}\geq 1$), the left hand side of the above inequality can be bounded by $c(s^{\prime})^{-8}$ for some constant $c>0$, which can then be used to bound (48).
Combining all the above analysis, we have arrived at
$\mathbb{P}\left(|\hat{L}_{s}^{*}-\hat{L}_{s^{\prime}}|\geq\omega_{s^{\prime}}\right)\leq c(s^{\prime})^{-8},\quad\forall\theta\in\Theta_{s},$
for some absolute constant $c>0$. Recall the definition of $\hat{s}$ in Equation 14. For any $\theta\in\Theta_{s}$, we have
$\mathbb{P}\left(\hat{s}>s\right)$ $=\mathbb{P}\left(\exists s^{\prime}>s,|\hat{L}_{s^{\prime}}-L_{s}|>\omega_{s^{\prime}}\right)$
$\leq\sum_{s<s^{\prime}\leq d}\mathbb{P}\left(|\hat{L}_{s}^{*}-\hat{L}_{s^{\prime}}|\geq\omega_{s^{\prime}}\right)$
$\leq c_{3,0}\cdot s^{-7},$
for some constant $c_{3,0}>0$. This completes the proof. ∎
Proof of Lemma 7.
Let $\theta\in\Theta_{s}$ and ${\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}$.
For $s=s_{0}$, we have $\lambda_{*}(s)=0$ and $\hat{L}_{s}^{*}-L(\theta)=\sigma\sum_{j=1}^{d}\eta_{j}\xi_{j}.$ Therefore
$\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left[\hat{L}_{s}^{*}-L(\theta)\right]^{4}\leq$ $\sum_{j=1}^{d}\sigma^{4}\eta_{j}^{4}\mathbb{E}\xi_{j}^{4}+3\sum_{i\neq j}\sigma^{4}\eta_{i}^{2}\eta_{j}^{2}\mathbb{E}\xi_{i}^{2}\mathbb{E}\xi_{j}^{2}$
$\leq$ $\sigma^{4}(C_{2}^{*}\sum_{j=1}^{d}\eta_{j}^{4}+3(\sum_{j=1}^{d}\eta_{j}^{2})^{2}).$
Since now $\Phi_{\rm adp}(s;\eta)=\log(es)\sum_{j=1}^{d}\eta_{j}^{2}$, we have
$\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left[\hat{L}_{s}^{*}-L(\theta)\right]^{4}\lesssim\sigma^{4}\Phi_{\rm adp}^{2}(s;\eta).$
We next consider the case $s\leq s_{*}$. Following a similar decomposition as in Section A.2, we obtain
$\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left(\hat{L}_{s}^{*}-L(\theta)\right)^{4}$
[…]
$\quad+\underbrace{\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\biggl(\sum_{\begin{subarray}{c}j>j_{2}(s)\\ j\in S\end{subarray}}\eta_{j}y_{j}\mathbbm{1}\left\{|\eta_{j}y_{j}|\geq{~}\sigma\tau\lambda_{*}(s)\right\}-\sum_{\begin{subarray}{c}j>j_{2}(s)\\ j\in S\end{subarray}}\eta_{j}\theta_{j}\biggr)^{4}}_{\text{denoted as III}}\biggr].$
Term I. Similar to the previous analysis, we have
$\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\biggl(\sum_{j\leq j_{2}(s)}\eta_{j}\xi_{j}\biggr)^{4}\lesssim\sigma^{4}\biggl(\sum_{j\leq j_{2}(s)}\eta_{j}^{2}\biggr)^{2}\lesssim\sigma^{4}\Phi_{\rm adp}^{2}(s;\eta).$
Term II. Let $U=\left\{j>j_{2}(s)\mid j\notin S\right\}$. We have
$\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\biggl(\sum_{j\in U}\sigma\eta_{j}\xi_{j}\mathbbm{1}\left\{|\eta_{j}\xi_{j}|\geq{~}\tau\lambda_{*}(s)\right\}\biggr)^{4}$
$=$ $\sigma^{4}\left[\sum_{j\in U}\eta_{j_{1}}^{4}\mathbb{E}\biggl(\xi_{j_{1}}^{4}\mathbbm{1}\left\{|\xi_{j_{1}}|\geq\frac{{~}\tau\lambda_{*}(s)}{|\eta_{j_{1}}|}\right\}\biggr)\right$.
$\qquad+\left.3\sum_{\begin{subarray}{c}j_{1},j_{2}\in U\\ j_{1}\neq j_{2}\end{subarray}}\eta_{j_{1}}^{2}\eta_{j_{2}}^{2}\mathbb{E}\left(\xi_{j_{1}}^{2}\xi_{j_{2}}^{2}\mathbbm{1}\left\{|\xi_{j_{1}}|\geq\frac{{~}\tau\lambda_{*}(s)}{|\eta_{j_{1}}|}\right\}\mathbbm{1}\left\{|\xi_{j_{2}}|\geq\frac{{~}\tau\lambda_{*}(s)}{|\eta_{j_{2}}|}\right\}\right)^{2}\right]$
$\leq$ $\sigma^{4}\left[\sqrt{C_{4}^{*}}\sum_{j\in U}|\eta_{j}|^{4}\exp\left(-\frac{(\beta_{*})_{+}}{|\eta_{j}|^{\alpha}}\right)+3C_{2}^{*}\left(\sum_{j\in U}|\eta_{j}|^{2}\exp\left(-\frac{(\beta_{*})_{+}}{|\eta_{j}|^{\alpha}}\right)\right)^{2}\right]$
$\leq$ $\sigma^{4}\left[\sqrt{C_{4}^{*}}\lambda_{*}^{2}(s)\sum_{j=1}^{d}|\eta_{j}|^{4}\exp\left(-\frac{(\beta_{*})_{+}}{|\eta_{j}|^{\alpha}}\right)+3C_{2}^{*}\left(\sum_{j=1}^{d}|\eta_{j}|^{2}\exp\left(-\frac{(\beta_{*})_{+}}{|\eta_{j}|^{\alpha}}\right)\right)^{2}\right]$
$\lesssim$ $\sigma^{4}\Phi_{\rm adp}^{2}(s;\eta).$
Term III.
$\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\biggl(\sum_{\begin{subarray}{c}j>j_{2}(s)\\ j\in S\end{subarray}}\eta_{j}y_{j}\mathbbm{1}\left\{|\eta_{j}y_{j}|\geq{~}\sigma\tau\lambda_{*}(s)\right\}-\sum_{\begin{subarray}{c}j>j_{2}(s)\\ j\in S\end{subarray}}\eta_{j}\theta_{j}\biggr)^{4}$
$\leq$ $8s^{3}\sum_{\begin{subarray}{c}j>j_{2}(s)\\ j\in S\end{subarray}}\left[\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left(\eta_{j}y_{j}\mathbbm{1}\left\{|\eta_{j}y_{j}|<{~}\sigma\tau\lambda_{*}(s)\right\}\right)^{4}+\sigma^{4}\eta_{j}^{4}\cdot\mathbb{E}\xi_{j}^{4}\right]$
$\leq$ $8s^{3}\sum_{\begin{subarray}{c}j>j_{2}(s)\\ j\in S\end{subarray}}\left[2^{4/\alpha}\sigma^{4}\tau^{4}\lambda_{*}^{4}(s)+C_{2}^{*}\sigma^{4}\eta_{j}^{4}\right]\lesssim\sigma^{4}\Phi_{\rm adp}^{2}(s;\eta).$
Therefore, we have obtained that
$\sup_{\theta\in\Theta_{s}}\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\left(\hat{L}_{s}^{*}-L(\theta)\right)^{4}\lesssim\sigma^{4}\Phi_{\rm adp}^{2}(s;\eta).$
Combined with Proposition 2, we have completed the proof. ∎
Appendix B Auxiliary Results
B.1 Existence and uniqueness of the solution to Equation 5
In this subsection, we show that Equation 5 has a unique solution. As pointed out in Section 2, it is sufficient to shows that the left hand side of Equation 5 is a continuous and strictly decreasing function in $\beta$ on $\mathbb{R}$.
Lemma 10.
The following function $\phi:\mathbb{R}\to\mathbb{R}$ is continuous and strictly decreasing:
$\phi(\beta)=\frac{\sum_{j=1}^{d}|\eta_{j}|\exp(-\beta/|\eta_{j}|^{\alpha})}{\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})}}.$
Proof of Lemma 10.
The function $\phi$ is continuous and differentiable. Its derivative can be written as
$\phi^{\prime}(\beta)$
$=$ $\frac{1}{2}\Big(\sum_{j=1}^{d}\eta_{j}^{2}e^{-\beta/|\eta_{j}|^{\alpha}}\Big)^{-3/2}\times$
$\qquad\left\{\Big(\sum_{j=1}^{d}|\eta_{j}|e^{-\beta/|\eta_{j}|^{\alpha}}\Big)\Big(\sum_{j=1}^{d}|\eta_{j}|^{2-\alpha}e^{-\beta/|\eta_{j}|^{\alpha}}\Big)-2\Big(\sum_{j=1}^{d}|\eta_{j}|^{1-\alpha}e^{-\beta/|\eta_{j}|^{\alpha}}\Big)\Big(\sum_{j=1}^{d}\eta_{j}^{2}e^{-\beta/|\eta_{j}|^{\alpha}}\Big)\right\}$
$=$ $\frac{1}{2}\Big(\sum_{j=1}^{d}\eta_{j}^{2}e^{-\beta/|\eta_{j}|^{\alpha}}\Big)^{-3/2}\Big(\sum_{j=1}^{d}{|\eta_{j}|^{2-\alpha}}e^{-\beta/|\eta_{j}|^{\alpha}}\Big)^{2}$
$\quad\times\Bigg\{\frac{\sum_{j=1}^{d}|\eta_{j}|e^{-\beta/|\eta_{j}|^{\alpha}}}{\sum_{j=1}^{d}{|\eta_{j}|}^{1-\alpha}e^{-\beta/|\eta_{j}|^{\alpha}}}\cdot\frac{\sum_{j=1}^{d}|\eta_{j}|^{2-\alpha}e^{-\beta/|\eta_{j}|^{\alpha}}}{\sum_{j=1}^{d}{|\eta_{j}|}^{1-\alpha}e^{-\beta/|\eta_{j}|^{\alpha}}}-2\cdot\frac{\sum_{j=1}^{d}\eta_{j}^{2}e^{-\beta/|\eta_{j}|^{\alpha}}}{\sum_{j=1}^{d}{|\eta_{j}|}^{1-\alpha}e^{-\beta/|\eta_{j}|^{\alpha}}}\Bigg\}.$
Define a random variable $Y$ following a probability measure $\mu=\sum_{j=1}^{d}w_{j}\delta_{|\eta_{j}|}$, where
$w_{j}=\frac{{|\eta_{j}|}^{1-\alpha}\exp(-\beta/|\eta_{j}|^{\alpha})}{\sum_{j=1}^{d}{|\eta_{j}|}^{1-\alpha}\exp(-\beta/|\eta_{j}|^{\alpha})}.$
Since $|\eta_{j}|>0,\forall j\in[d]$, we have $\mathbb{E}Y^{\alpha+1}>0$ and $\mathbb{E}Y\cdot\mathbb{E}Y^{\alpha}\leq\mathbb{E}Y^{1+\alpha}$. Therefore,
$\phi^{\prime}(\beta)$
$=$ $\frac{1}{2}\left(\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})\right)^{-3/2}\left(\sum_{j=1}^{d}\exp(-\beta/|\eta_{j}|^{\alpha})/\eta_{j}\right)^{2}\left[\mathbb{E}Y\cdot\mathbb{E}Y^{\alpha}-2\mathbb{E}Y^{1+\alpha}\right]$
$<$ $0.$
B.2 Proof of Proposition 1
Proof.
By the definition of $j_{1}$, the following holds up to absolute constants:
$[\lambda_{o}s+\nu]^{2}\asymp$ $\lambda_{o}^{2}s^{2}+\nu^{2}$
$\asymp$ $\lambda_{o}^{2}s^{2}+\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta_{+}/|\eta_{j}|^{\alpha})$
$\asymp$ $\lambda_{o}^{2}s^{2}+\sum_{j\leq j_{1}}\eta_{j}^{2}\exp(-\beta_{+}/|\eta_{j}|^{\alpha})+\sum_{j>j_{1}}\eta_{j}^{2}\exp(-\beta_{+}/|\eta_{j}|^{\alpha})$
$\asymp$ $\lambda_{o}^{2}s^{2}+\sum_{j\leq j_{1}}\eta_{j}^{2}+\sum_{j>j_{1}}\eta_{j}^{2}\exp(-\beta_{+}/|\eta_{j}|^{\alpha}).$
If $\sum_{j>j_{1}}\eta_{j}^{2}\exp(-\beta_{+}/|\eta_{j}|^{\alpha})\leq\sum_{j\leq j_{1}}\eta_{j}^{2}\exp(-\beta_{+}/|\eta_{j}|^{\alpha})$, then we have $\lambda_{o}s+\nu]^{2}\asymp\lambda_{o}^{2}s^{2}+\sum_{j\leq j_{1}}\eta_{j}^{2}$. Otherwise, we have
$\sum_{j>j_{1}}\eta_{j}^{2}\exp(-\beta_{+}/|\eta_{j}|^{\alpha})$ $\leq\lambda_{o}\sum_{j>j_{1}}|\eta_{j}|\exp(-\beta_{+}/|\eta_{j}|^{\alpha})$
$\leq\frac{1}{2}\lambda_{o}s\cdot\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta_{+}/|\eta_{j}|^{\alpha})}$
$\leq\frac{1}{2}\lambda_{o}s\cdot\sqrt{2\sum_{j>j_{1}}\eta_{j}^{2}\exp(-\beta_{+}/|\eta_{j}|^{\alpha})},$
where the first inequality holds since $|\eta_{j}|\leq\lambda_{o}$ for $j\leq j_{1}$, the second inequality holds since the definition of $\beta$ implies that $\sum_{j=1}^{d}|\eta_{j}|\exp(-\beta_{+}/|\eta_{j}|^{\alpha})\leq s/2\cdot\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta_{+}/|\eta_{j}|^{\alpha})}$. Therefore, we have $\sum_{j>j_{1}}\eta_{j}^{2}\exp(-\beta_{+}/|\eta_{j}|^{\alpha})\lesssim\lambda_{o}^{2}s^{2}$, which completes the proof. ∎
B.3 Proof of Proposition 2
Proof.
Note that
$\Phi_{*}(s;\eta)=$ $\log(es)\left[\sum_{j=1}^{d}\eta_{j}^{2}\exp\left(-\frac{(\beta_{*})_{+}}{|\eta_{j}|^{\alpha}}\right)+\lambda^{2}_{*}(s)\left(\frac{s}{\sqrt{\log(es)}}\right)^{2}\right]$
$\asymp$ $\log(es)\left[\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp\left(-\frac{(\beta_{*})_{+}}{|\eta_{j}|^{\alpha}}\right)}+\lambda_{*}(s)\cdot\frac{\sum_{j=1}^{d}|\eta_{j}|\exp\left(-{\beta_{*}}/{|\eta_{j}|^{\alpha}}\right)}{\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp\left(-{\beta_{*}}/{|\eta_{j}|^{\alpha}}\right)}}\right]^{2}$
where the second equation uses the definition of $\lambda_{*}(s)$ in (11). Note that $\beta_{*}$ is a non-increasing function of $s$. Let
$\phi_{1}(\beta)=\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})}+\beta^{1/\alpha}\cdot\frac{\sum_{j=1}^{d}|\eta_{j}|\exp(-\beta/|\eta_{j}|^{\alpha})}{\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})}}$
and
$\phi_{2}(\beta)=\frac{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})}{\sum_{j=1}^{d}|\eta_{j}|\exp(-\beta/|\eta_{j}|^{\alpha})}.$
Since $\log(e(s\wedge s_{0}))$ is nondecreasing, whereas $\log(es)/{s^{2}}$ is nonincreasing, we only need to show that there exists a constant $c>0$ such that for all $\beta_{1}\geq\beta_{2}>0$,
$\phi_{1}(\beta_{1})\leq c\cdot\phi_{1}(\beta_{2})\quad\text{and}\quad\phi_{2}(\beta_{1})\geq c\cdot\phi_{2}(\beta_{2}).$
The proof for $\phi_{2}(\beta)$ is simple, as we have
$\phi_{2}^{\prime}(\beta)=\Biggl(\sum_{j=1}^{d}\eta_{j}\exp\!\bigl(-\beta/|\eta_{j}|^{\alpha}\bigr)\Biggr)^{-2}\Biggl[$ $\sum_{j=1}^{d}\eta_{j}^{2}\exp\!\bigl(-\beta/|\eta_{j}|^{\alpha}\bigr)\!\sum_{j=1}^{d}|\eta_{j}|^{1-\alpha}\exp\!\bigl(-\beta/|\eta_{j}|^{\alpha}\bigr)\Biggr$.
$\Biggl.-\sum_{j=1}^{d}|\eta_{j}|^{2-\alpha}\exp\!\bigl(-\beta/|\eta_{j}|^{\alpha}\bigr)\!\sum_{j=1}^{d}|\eta_{j}|\exp\!\bigl(-\beta/|\eta_{j}|^{\alpha}\bigr)\Biggr].$
The above term is non-negative following the proof of Lemma 10. Therefore, $\phi_{2}(\beta)$ is non-decreasing in $\beta$. The rest analysis is on $\phi_{1}(\beta)$. Following the notation in Section B.1, we have
$\phi_{1}(\beta)=\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})}+\beta^{1/\alpha}\phi(\beta).$
There are two components in $\phi_{1}(\beta)$, and the first term
$\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})}$
is clearly strictly decreasing in $\beta$. Therefore, for $\beta=\beta_{1}$, we can assume
$\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta^{2}/|\eta_{j}|^{\alpha})}<\left(\frac{\alpha\beta}{2}\right)^{1/\alpha}\phi(\beta).$
Otherwise, we have
$\phi_{1}(\beta_{1})=$ $\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta_{1}/|\eta_{j}|^{\alpha})}+\beta_{1}^{1/\alpha}\phi(\beta_{1})$
$\leq$ $\left[1+\left(\frac{2}{\alpha}\right)^{1/\alpha}\right]\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta_{1}/|\eta_{j}|^{\alpha})}$
$\leq$ $\left[1+\left(\frac{2}{\alpha}\right)^{1/\alpha}\right]\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta_{2}/|\eta_{j}|^{\alpha})}$
$\leq$ $\left[1+\left(\frac{2}{\alpha}\right)^{1/\alpha}\right]\phi_{1}(\beta_{2}).$
Define
$\beta_{\min}=\min\left\{\beta^{\prime}\geq\beta_{2}:\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta^{2}/|\eta_{j}|^{\alpha})}\leq\left(\frac{\alpha\beta}{2}\right)^{1/\alpha}\phi(\beta)\,\,\text{holds}\,\,\forall\beta\in[\beta^{\prime},\beta_{1}]\right\}.$
The existence of such $\beta_{*}$ is guaranteed by the continuity of functions involved in the above definition.
Below we are going to show that $\phi_{1}(\beta)$ is strictly decreasing in $\beta\in[\beta_{\min},\beta_{1}]$.
Following the calculations in Section B.1, we have
$\phi^{\prime}_{1}(\beta)=-\frac{\sum_{j=1}^{d}|\eta_{j}|^{2-\alpha}\exp(-\beta/|\eta_{j}|^{\alpha})}{2\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})}}+\frac{\beta^{1/\alpha-1}\sum_{j=1}^{d}|\eta_{j}|\exp(-\beta/|\eta_{j}|^{\alpha})}{\alpha\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})}}+\beta^{1/\alpha}\phi^{\prime}(\beta)$
$=$ ${\tiny-\frac{\sum_{j=1}^{d}|\eta_{j}|^{2-\alpha}\exp(-\beta/|\eta_{j}|^{\alpha})}{2\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})}}+\frac{\beta^{1/\alpha-1}\sum_{j=1}^{d}|\eta_{j}|\exp(-\beta/|\eta_{j}|^{\alpha})}{\alpha\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})}}-\beta^{1/\alpha}\frac{\sum_{j=1}^{d}|\eta_{j}|^{1-\alpha}\exp(-\beta/|\eta_{j}|^{\alpha})}{2\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})}}}$
$+$ $\beta^{1/\alpha}\left[\frac{{\tiny\left(\sum_{j=1}^{d}|\eta_{j}|\exp(-\beta/|\eta_{j}|^{\alpha})\right)\left(\sum_{j=1}^{d}|\eta_{j}|^{2-\alpha}\exp(-\beta/|\eta_{j}|^{\alpha})\right)-\left(\sum_{j=1}^{d}|\eta_{j}|^{1-\alpha}\exp(-\beta/|\eta_{j}|^{\alpha})\right)\left(\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})\right)}}{2\cdot\left(\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})\right)^{3/2}}\right].$
The last term is non-positive following the proof of Lemma 10. And the first three terms can be rearranged as follows:
$-\frac{\sum_{j=1}^{d}|\eta_{j}|^{2-\alpha}\exp(-\beta/|\eta_{j}|^{\alpha})}{2\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})}}+\frac{\beta^{1/\alpha-1}\sum_{j=1}^{d}|\eta_{j}|\exp(-\beta/|\eta_{j}|^{\alpha})}{\alpha\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})}}-\frac{\beta^{1/\alpha}\sum_{j=1}^{d}|\eta_{j}|^{1-\alpha}\exp(-\beta/|\eta_{j}|^{\alpha})}{2\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})}}$
$=$ $-\frac{\beta^{1/\alpha}}{2\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})}}\sum_{j=1}^{d}\left[\left(\beta^{-1/\alpha}|\eta_{j}|-\frac{2}{\alpha\beta}|\eta_{j}|^{\alpha}+1\right)|\eta_{j}|^{1-\alpha}\exp(-\beta/|\eta_{j}|^{\alpha})\right].$
Proposition 5.
For any $\alpha>0$, we have
$f(t)=\alpha t^{2}-(\alpha+1)t+t^{1-\alpha}\geq 0\quad\forall t\geq 0.$
Applying Proposition 5 with $t={|\eta_{j}|}(\frac{2}{\alpha\beta})^{1/\alpha}$, we obtain
$\alpha|\eta_{j}|^{2}\left(\frac{2}{\alpha\beta}\right)^{2/\alpha}-(\alpha+1)|\eta_{j}|\left(\frac{2}{\alpha\beta}\right)^{1/\alpha}+|\eta_{j}|^{1-\alpha}\left(\frac{2}{\alpha\beta}\right)^{1/\alpha-1}\;\geq\;0.$
Rearranging yields
$|\eta_{j}|^{1-\alpha}-\frac{2}{\alpha\beta}\,|\eta_{j}|\;\geq\;\frac{2}{\beta}\left[\,|\eta_{j}|-\left(\frac{2}{\alpha\beta}\right)^{1/\alpha}|\eta_{j}|^{2}\right].$
From the definition of $\beta_{\min}$, we have
$\sum_{j=1}^{d}\left[\left(\frac{\alpha\beta}{2}\right)^{1/\alpha}|\eta_{j}|-\eta_{j}^{2}\right]\exp(-\beta/|\eta_{j}|^{\alpha})\geq 0,\forall\beta\in[\beta_{\min},\beta_{1}].$
Therefore, we have
$\sum_{j=1}^{d}\left(|\eta_{j}|^{1-\alpha}-\frac{2}{\alpha\beta}|\eta_{j}|\right)\exp(-\beta/\eta_{j}^{2})$
$\geq$ $\frac{2}{\beta}\left(\frac{2}{\alpha\beta}\right)^{\alpha/2}\sum_{j=1}^{d}\left[\left(\frac{\alpha\beta}{2}\right)^{1/\alpha}|\eta_{j}|-\eta_{j}^{2}\right]\exp(-\beta/\eta_{j}^{2})\geq 0.$
which implies that $\phi_{1}^{\prime}(\beta)\leq 0$ for all $\beta\in[\beta_{\min},\beta_{1}]$, and hence $\phi_{1}(\beta)$ is strictly decreasing on the interval $\beta_{\min},\beta_{1}]$.
If $\beta_{\min}=\beta_{2}$, then the proof is completed. Otherwise, from the definition of $\beta_{\min}$, we have
$\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta_{\min}^{2}/|\eta_{j}|^{\alpha})}=\left(\frac{\alpha\beta}{2}\right)^{1/\alpha}\phi(\beta_{\min})$
and therefore,
$\phi_{1}(\beta_{1})\leq\phi_{1}(\beta_{\min})=$ $\left[1+\left(\frac{2}{\alpha}\right)^{1/\alpha}\right]\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta_{\min}^{2}/|\eta_{j}|^{\alpha})}$
$\leq$ $\left[1+\left(\frac{2}{\alpha}\right)^{1/\alpha}\right]\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}\exp(-\beta_{2}^{2}/|\eta_{j}|^{\alpha})}$
$\leq$ $\left[1+\left(\frac{2}{\alpha}\right)^{1/\alpha}\right]\phi_{1}(\beta_{2}).$
B.4 Proof of Proposition 4
Proof.
Without loss of generality, we assume that $s^{2}\leq\lceil s^{2}t^{2}\rceil\leq d$. Then we have
$\sum_{j\leq\lceil s^{2}t^{2}\rceil}\eta_{j}^{2}=\sum_{j\leq s^{2}}\eta_{j}^{2}+\sum_{s^{2}<j\leq\lceil s^{2}t^{2}\rceil}\eta_{j}^{2},$
where the second term has at most $2s^{2}t^{2}$ terms and each term is at most $\eta_{s^{2}}^{2}$. Therefore, we have
$\sum_{j\leq\lceil s^{2}t^{2}\rceil}\eta_{j}^{2}$ $\leq\sum_{j\leq s^{2}}\eta_{j}^{2}+2s^{2}t^{2}\cdot\eta_{s^{2}}^{2}\leq(1+2t^{2})\sum_{j\leq s^{2}}\eta_{j}^{2}.$
Now the proof is completed. ∎
Appendix C Additional proof of the lower bounds
C.1 Proof of Theorem 5
Let $f_{0}:\mathbb{R}\to[0,\infty)$ be a probability density with the following properties: $f_{0}$ is continuously differentiable, symmetric about 0, supported on [3/2,3/2][-3/2,3/2], with variance $1$ and finite Fisher information $I_{f_{0}}=\int(f_{0}^{\prime}(x))^{2}f_{0}(x){{\rm d\,}}x<\infty$. The existence of such a function is guaranteed by [11, Lemma 7]. Without loss of generality, we assume that $\sigma=1$.
Note that based on (28) and Proposition 4, we only consider the case where
$s^{2}\lambda^{2}_{\cal H}\geq C\sum_{j\leq s^{2}}\eta_{j}^{2}$
for some sufficiently large constant $C>0$. The above inequality implies that
$\lambda_{\cal H}\geq\sqrt{C}\cdot\sqrt{\frac{1}{s^{2}}\sum_{j\leq s^{2}}\eta_{j}^{2}}\geq\sqrt{C}\cdot|\eta_{s^{2}}|,$
which implies $\lambda_{\cal H}/|\eta_{j}|\geq\sqrt{C}$, for any $j\geq{s^{2}}$.
Let $\theta_{1},\ldots,\theta_{d}$ be independently Bernoulli random variables with the probability of success $\mathbb{P}(\theta_{j}=1)=\pi_{j},\forall j=1,\ldots,d$, where $\pi_{j}$ is defined as follows:
$\pi_{j}=\left\{\begin{array}[]{cc}0,&j\leq s^{2}\\ \displaystyle c\cdot\exp\left(-\left|\frac{\lambda_{\cal H}}{\eta_{j}}\right|^{\alpha}\right),&j\geq s^{2}\end{array}\right$.,
where $c>0$ is a constant specified later. Then by (27), we have $\sum_{j=1}^{d}\pi_{j}=c\cdot s.$ Let $\mu$ denote the distribution of $\theta=(\gamma_{1}\theta_{1},\ldots,\gamma_{d}\theta_{d})$ with $\gamma_{j}=c_{2}\cdot\lambda_{\cal H}/\eta_{j},\forall j\in[d]$, where $c_{2}$ is some sufficiently small constant.
Note that for $j\geq s^{2}$, we have
$\gamma_{j}\cdot\pi_{j}=\displaystyle c\cdot c_{2}\cdot\frac{\lambda_{\cal H}}{\eta_{j}}\cdot\exp\left(-\left|\frac{\lambda_{\cal H}}{\eta_{j}}\right|^{\alpha}\right),$
and
$\gamma_{j}^{2}\cdot\pi_{j}(1-\pi_{j})\leq\displaystyle c\cdot c_{2}^{2}\cdot(\frac{\lambda_{\cal H}}{\eta_{j}})^{2}\cdot\exp\left(-\left|\frac{\lambda_{\cal H}}{\eta_{j}}\right|^{\alpha}\right).$
Since $\sup_{x}|x^{k}\exp(-|x|^{\alpha})|=(e\alpha/k)^{-1/\alpha}$ only depends on $\alpha$, one can find $c_{2,1}$ such that for any $c_{2}\leq c_{2,1}$, the followings hold for all $j$:
$|\gamma_{j}\pi_{j}|<1/2,\qquad\gamma_{j}^{2}\cdot\pi_{j}(1-\pi_{j})<1/4.$ (49)
Then, we have
$L(\theta)$ $=\sum_{j=1}^{d}\eta_{j}\cdot\theta_{j}=c_{2}\cdot\lambda_{\cal H}\cdot\|\theta\|_{0}$
$\mathbb{E}_{\mu}\left[\exp(t\|\theta\|_{0})\right]$ $=\prod_{j=1}^{d}\left[1+\pi_{j}\cdot(\exp(t)-1)\right],\quad t\in\mathbb{R}.$
Using the basic inequality that $1+x\leq\exp(x)$ for $\forall x\in\mathbb{R}$, we have
$\mathbb{P}_{\mu}(\left\|\theta\right\|_{0}>s)\leq$ $\exp(-s)\cdot\mathbb{E}_{\mu}\left[\exp(\|\theta\|_{0})\right]$
$=$ $\exp(-s)\cdot\prod_{j=1}^{d}\left[1+\pi_{j}\cdot(e-1)\right]$
$\leq$ $\exp(-s)\exp\left[(e-1)\cdot\sum_{j=1}^{d}\pi_{j}\right]=\exp(-s\cdot((e-1)c-1)),$
and
$\mathbb{P}_{\mu}(\left\|\theta\right\|_{0}<c_{1}\cdot s)\leq$ $\exp(c_{1}\cdot s)\mathbb{E}_{\mu}\left[\exp(-\|\theta\|_{0})\right]$
$=$ $\exp(c_{1}\cdot s)\cdot\prod_{j=1}^{d}\left[1+\pi_{j}\cdot(e^{-1}-1)\right]$
$\leq$ $\exp(c_{1}\cdot s)\cdot\exp\left[(e^{-1}-1)\sum_{j=1}^{d}\pi_{j}\right]=\exp(s\cdot(c_{1}-(1-e^{-1})\cdot c)).$
For any $\delta\in(0,1)$, we may take $c_{1}=e^{-1}$ and $c=1+e\log(2/\delta)$ such that the right hand sides of the above two inequalities are both bounded by $\delta/2$. It follows that
$\mathbb{P}_{\mu}(\theta\in\Theta_{s},L(\theta)\geq c_{1}\cdot c_{2}\cdot s\cdot\lambda_{\cal H})=\mathbb{P}_{\mu}(c_{1}s\leq\theta\in\Theta_{s})\geq 1-\delta.$
With a little abuse of notation, we let $f_{j}$ be the density of the random variable $\tilde{\delta}_{j}=\sigma_{j}\delta_{j}+\gamma_{j}\cdot(\theta_{j}-\pi_{j})$ with $\delta_{j}$ being i.i.d. random variables with density $f_{0}$ and
$\sigma_{j}=\sqrt{1-\gamma_{j}^{2}\cdot\pi_{j}(1-\pi_{j})}.$
$\sigma_{j}$ is well-defined in view of Equation 49. By construction, we have
$\mathbb{E}\,\tilde{\delta}_{j}=0\quad\text{and}\quad\mathbb{E}\,\tilde{\delta}_{j}^{2}=1,\quad\forall j=1,\ldots,d.$
Equation 49 also implies that when $\theta_{j}=0$, the absolute value of $\tilde{\delta}_{j}=\sigma_{j}\delta_{j}-\gamma_{j}\cdot\pi_{j}$ is bounded by $2$ since $\sigma_{j}\delta_{j}\in[-3/2,3/2]$.
Define $F_{0}=f_{0}^{\otimes d}$ and $F_{1}=\bigotimes_{i=1}^{d}f_{j}$. Therefore,
$\mathbb{P}(|\tilde{\delta}_{j}|\geq t)\leq\mathbb{P}\left(\theta_{j}=1,\gamma_{j}\geq t-1\right)\leq c\cdot\exp\left(-\left|\frac{\lambda_{\cal H}}{\eta_{j}}\right|^{\alpha}\right)\mathbbm{1}\left\{c_{2}\cdot\frac{\lambda_{\cal H}}{\eta_{j}}\geq t-1\right\}.$
We bound the right hand side of the above inequality in two cases.
(i). For any $t\in(0,2)$, we only require $\tau>2/\log^{1/\alpha}2$ to have
$\mathbb{P}(|\tilde{\delta}_{j}|\geq t)\leq 2\exp\left\{-\left(\frac{t}{\tau}\right)^{\alpha}\right\}.$
Let’s choose $\tau_{\alpha}=2/\log^{1/\alpha}2$.
(ii). For any $t\geq 2$, we have $t-1\geq t/2$ and the right hand side can be bounded by
$c\cdot\exp\left\{-\left(\frac{t}{2c_{2}}\right)^{\alpha}\right\}.$
If we pick a $c_{2}$ sufficiently small such that $(c_{2})^{-\alpha}-(2/\tau_{\alpha})^{\alpha}>\log(c/2)$, then for any $\tau\geq\tau_{\alpha}$, we have
$\mathbb{P}(|\tilde{\delta}_{j}|\geq t)\leq c\cdot\exp\left\{-\left(\frac{t}{2c_{2}}\right)^{\alpha}\right\}\leq 2\exp\left\{-\left(\frac{t}{\tau}\right)^{\alpha}\right\}.$
We conclude that $F_{1}\in{\cal G}_{\alpha,\tau}^{\otimes}$. Define $P_{1}=\mathbb{P}_{0,F_{1}}$ and $P_{2}=\mathbb{P}_{\mu,F_{0}}$. To complete the proof using Lemma 1, it remains to control the total variation distance between $P_{1}$ and $P_{2}$ by $cc_{2}$ (up to a constant). This is achieved by using the condition in (29) and the following lemma.
Lemma 11.
In the proof of Equation 30, we have ${\rm TV}(P_{1},P_{2})\leq\bar{C}^{\prime}cc_{2}$ for some constant $\bar{C}^{\prime}$ that depends on the condition (29) and $I_{f_{0}}$.
This completes the proof of Equation 30 with Lemma 1.
C.2 Proof of Lemma 2
Proof.
Consider the scale family $\{F_{\alpha}(\sigma)\}$ on $\mathbb{R}$ with density
$f_{\alpha}(x;\sigma)=\frac{1}{c_{\sigma}}\exp\!\left(-\frac{|x|^{\alpha}}{\sigma^{\alpha}}\right),\qquad\alpha>0,\ \sigma>0,$
where the normalizing constant is
$c_{\sigma}=\int_{\mathbb{R}}\exp\!\left(-\frac{|x|^{\alpha}}{\sigma^{\alpha}}\right)\,{\rm d\,}x=\sigma\!\int_{\mathbb{R}}e^{-|y|^{\alpha}}\,{\rm d\,}y=\sigma\cdot\frac{2}{\alpha}\,\Gamma\!\left(\frac{1}{\alpha}\right).$
We choose $\sigma=\sigma_{\alpha}$ so that $\mathbb{E}_{F_{\alpha}(\sigma_{\alpha})}[X^{2}]=1$. Using the change of variables $x=\sigma y$,
$\int_{\mathbb{R}}x^{2}e^{-|x|^{\alpha}/\sigma^{\alpha}}\,{\rm d\,}x=\sigma^{3}\int_{\mathbb{R}}y^{2}e^{-|y|^{\alpha}}\,{\rm d\,}y=\sigma^{3}\cdot\frac{2}{\alpha}\,\Gamma\!\left(\frac{3}{\alpha}\right),$
and
$\int_{\mathbb{R}}e^{-|x|^{\alpha}/\sigma^{\alpha}}\,{\rm d\,}x=\sigma\int_{\mathbb{R}}e^{-|y|^{\alpha}}\,{\rm d\,}y=\sigma\cdot\frac{2}{\alpha}\,\Gamma\!\left(\frac{1}{\alpha}\right).$
The condition $\mathbb{E}[X^{2}]=1$ is equivalent to
$\frac{1}{c_{\sigma}}\int_{\mathbb{R}}x^{2}e^{-|x|^{\alpha}/\sigma^{\alpha}}\,{\rm d\,}x=1\quad\Longleftrightarrow\quad\int_{\mathbb{R}}x^{2}e^{-|x|^{\alpha}/\sigma^{\alpha}}\,{\rm d\,}x=\int_{\mathbb{R}}e^{-|x|^{\alpha}/\sigma^{\alpha}}\,{\rm d\,}x,$
which yields
$\sigma_{\alpha}^{2}=\frac{\displaystyle\int_{\mathbb{R}}e^{-|y|^{\alpha}}\,{\rm d\,}y}{\displaystyle\int_{\mathbb{R}}y^{2}e^{-|y|^{\alpha}}\,{\rm d\,}y}=\frac{\frac{2}{\alpha}\Gamma(1/\alpha)}{\frac{2}{\alpha}\Gamma(3/\alpha)}=\frac{\Gamma(1/\alpha)}{\Gamma(3/\alpha)}.$
Hence, taking
$\sigma_{\alpha}=\sqrt{\frac{\Gamma(1/\alpha)}{\Gamma(3/\alpha)}}$
makes $f_{\sigma_{\alpha}}$ mean-zero (by symmetry) with unit variance. In order to satisfy the tail decay condition, $\tau$ is required to satisfy
$\mathbb{E}_{F_{\alpha}(\sigma_{\alpha})}\exp\left(\frac{|X|^{\alpha}}{\tau^{\alpha}}\right)=\int_{\mathbb{R}}\frac{1}{c_{\sigma_{\alpha}}}\exp\left[\left(\frac{1}{\tau^{\alpha}}-\frac{1}{\sigma_{\alpha}^{\alpha}}\right)|x|^{\alpha}\right]{\rm d\,}x\leq 2.$ (50)
The existence of such $\tau$ is guaranteed by the fact that the above integral is increasing in $\tau$ and when $\tau\to\infty$, the integral converges to 1. In fact, if we choose $\tau_{\alpha}=\sigma_{\alpha}(1-2^{-\alpha})^{-1/\alpha}$, then for any $\tau\geq\tau_{\alpha}$, Equation 50 holds; we defer the derivation of $\tau_{\alpha}$ to the end of this proof. Therefore, we already find a distribution $F_{\alpha}(\sigma_{\alpha})\in{\cal G}_{\alpha,\tau_{\alpha}}$.
The remaining is to bound the $\chi^{2}$ divergence between $f_{\alpha}^{(1)}(\cdot)=f_{\alpha}(\cdot-\gamma;\sigma_{\alpha})$ and $f_{\alpha}^{(0)}(\cdot)=f_{\alpha}(\cdot-\gamma;\sigma_{\alpha})$. We have
$\chi^{2}(f_{\alpha}^{(1)}\|f_{\alpha}^{(0)})=\int_{\mathbb{R}}\frac{f_{\alpha}^{(1)}(x)}{f_{\alpha}^{(0)}(x)}\,{\rm d\,}x$ $=\frac{1}{c_{\sigma_{\alpha}}}\int_{\mathbb{R}}\exp\left(\frac{|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}-\frac{2|x-\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,{\rm d\,}x.$
  • When $0<\alpha\leq 1$, we have $|x|^{\alpha}\leq|x-\gamma|^{\alpha}+|\gamma|^{\alpha}$ and therefore, we have
    $\chi^{2}(f_{\alpha}^{(1)}\|f_{\alpha}^{(0)})$ $=\frac{1}{c_{\sigma_{\alpha}}}\int_{\mathbb{R}}\exp\left(\frac{2|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}-\frac{2|x-\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\exp\left(-\frac{|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,{\rm d\,}x.$
    $\leq\frac{1}{c_{\sigma_{\alpha}}}\int_{\mathbb{R}}\exp\left(\frac{2|\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\exp\left(-\frac{|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,{\rm d\,}x\leq\exp\left(\frac{2|\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)$
  • When $\alpha>1$, we have $|x|^{\alpha}\leq(1-s)^{1-\alpha}|x-\gamma|^{\alpha}+s^{1-\alpha}|\gamma|^{\alpha}$ for any $s\in(0,1)$ and therefore, we have
    $\chi^{2}(f_{\alpha}^{(1)}\|f_{\alpha}^{(0)})$ $=\frac{1}{c_{\sigma_{\alpha}}}\int_{\mathbb{R}}\exp\left(\frac{|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}-\frac{2|x-\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,{\rm d\,}x.$
    $\leq\frac{1}{c_{\sigma_{\alpha}}}\int_{\mathbb{R}}\exp\left(\frac{s^{1-\alpha}|\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\exp\left(-\frac{(2-(1-s)^{1-\alpha})|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,{\rm d\,}x$
    $\leq\exp\left(\frac{2s^{1-\alpha}|\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\cdot\frac{1}{c_{\sigma_{\alpha}}}\int_{\mathbb{R}}\exp\left(-\frac{(1-2(1-s)^{1-\alpha})|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,{\rm d\,}x$
    We can choose $s=s_{\alpha}\in(0,1)$ such that
    $\int_{\mathbb{R}}\exp\left(-\frac{(1-2(1-s_{\alpha})^{1-\alpha})|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,{\rm d\,}x<\infty.$
    And therefore, we have
    $\chi^{2}(f_{\alpha}^{(1)}\|f_{\alpha}^{(0)})\leq\frac{1}{c_{\sigma_{\alpha}}}\int_{\mathbb{R}}\exp\left(-\frac{(1-2(1-s_{\alpha})^{1-\alpha})|x|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right)\,{\rm d\,}x\cdot\exp\left(\frac{2s_{\alpha}^{1-\alpha}|\gamma|^{\alpha}}{\sigma_{\alpha}^{\alpha}}\right).$
Necessary and sufficient condition for Equation 50.
We prove that Equation 50 holds if and only if $\tau\geq\sigma_{\alpha}(1-2^{-\alpha})^{-1/\alpha}$. In the following, the expectation on the left hand side of Equation 50 is denoted as $I(\tau)$.
It is clear that $\tau>\sigma_{\alpha}$ is needed for Equation 50 to hold. For such a $\tau$, define $a(\tau)=\frac{1}{\sigma_{\alpha}^{\alpha}}-\frac{1}{\tau^{\alpha}}>0$. We have
$I(\tau)=$ $\int_{\mathbb{R}}\frac{1}{c_{\sigma_{\alpha}}}\exp\left[\left(\frac{1}{\tau^{\alpha}}-\frac{1}{\sigma_{\alpha}^{\alpha}}\right)|x|^{\alpha}\right]{\rm d\,}x$
$=$ $\frac{2}{c_{\sigma_{\alpha}}}\int_{0}^{\infty}e^{-a(\tau)x^{\alpha}}{\rm d\,}x$
$=$ $\frac{2}{c_{\sigma_{\alpha}}}\frac{1}{\alpha}\int_{0}^{\infty}y^{1/\alpha-1}e^{-a(\tau)y}{\rm d\,}y$
$=$ $\frac{2}{c_{\sigma_{\alpha}}}\frac{1}{\alpha}\Gamma(1/\alpha)[a(\tau)]^{-1/\alpha}.$
Since $c_{\sigma_{\alpha}}=\sigma_{\alpha}\cdot\frac{2}{\alpha}\Gamma(1/\alpha)$, we have
$I(\tau)$ $=\frac{1}{\sigma_{\alpha}}\left(\frac{1}{\sigma_{\alpha}^{\alpha}}-\frac{1}{\tau^{\alpha}}\right)^{-1/\alpha}$
$=\left(1-\frac{\sigma_{\alpha}^{\alpha}}{\tau^{\alpha}}\right)^{-1/\alpha}.$
Therefore, it is straightforward to see that
$I(\tau)\leq 2\Leftrightarrow 1-\frac{\sigma_{\alpha}^{\alpha}}{\tau^{\alpha}}\geq 2^{-\alpha}\Leftrightarrow\tau\geq\sigma_{\alpha}(1-2^{-\alpha})^{-1/\alpha}.$
This completes the proof.
C.3 Additional proof of the lower bounds in  Theorem 1
Proof.
We now prove the lower bound in Theorem 1 when
$s\leq C_{1}\quad\text{or}\quad\lambda_{o}s+\nu\leq C_{2}|\eta_{1}|,$
for some constants $C_{1},C_{2}>0$ and for any $\alpha>0$, $\tau\geq\tau_{\alpha}$.
Case 1: $\lambda_{o}s+\nu\leq C_{2}|\eta_{1}|$. We need to show that
$\inf_{\hat{T}}\sup_{\theta\in\Theta_{s}}\sup_{{\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}}\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\!\left(|\hat{T}-L(\theta)|^{2}\right)\;\gtrsim\;\eta_{1}^{2}.$
Let $c>0$ be a small constant to be specified later, and let $\mu$ be the point mass at $(c\,\mathrm{sgn}(\eta_{1}),0,\ldots,0)$. Recall the definition of $F_{\alpha}^{\otimes}$ in Section 8.2, and set
$P_{1}=\mathbb{P}_{\mu_{1},F_{\alpha}^{\otimes}},\qquad P_{2}=\mathbb{P}_{\mu,F_{\alpha}^{\otimes}}.$
From the proof of Lemma 2, the location family $\{f_{\alpha}^{(0)}(\cdot-\gamma):\gamma\in\mathbb{R}\}$ is regular and has finite Fisher information $I_{0}>0$. By Lemma 3,
${\bf H}^{2}(P_{2},P_{1})={\bf H}^{2}\!\big(f_{\alpha}^{(0)}(\cdot-c),f_{\alpha}^{(0)}\big)\;\leq\;c^{2}I_{0}.$
Choosing $c<\alpha_{1}I_{0}^{-1/2}$ for some absolute $\alpha_{1}\in(0,1)$ gives
$\mathrm{TV}(P_{1},P_{2})\leq{\bf H}(P_{2},P_{1})\leq\alpha_{1}.$
Moreover,
$\mu\!\left(L(\theta)\geq c|\eta_{1}|,\theta\in\Theta_{s}\right)=1.$
Thus, applying Lemma 1 with $t=\frac{1}{2}c|\eta_{1}|$ yields
$\inf_{\hat{T}}\sup_{\theta\in\Theta_{s}}\mathbb{P}\!\left(|\hat{T}-L(\theta)|\geq\frac{1}{2}c|\eta_{1}|\right)\;\geq\;\frac{1-\alpha_{1}}{2}>0.$
Case 2: $s\leq C_{1}$. We first show that it suffices to prove that
$\inf_{\hat{T}}\sup_{\theta\in\Theta_{s}}\sup_{{\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}}\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\!\left(|\hat{T}-L(\theta)|^{2}\right)\;\gtrsim\;\lambda_{o}^{2},\qquad\text{if }\lambda_{o}\geq C_{3}|\eta_{1}|$ (51)
for some sufficiently large constant $C_{3}>0$.
We already know that
$\inf_{\hat{T}}\sup_{\theta\in\Theta_{s}}\sup_{{\mathcal{P}}_{\xi}\in{\cal G}_{\alpha,\tau}^{\otimes}}\mathbb{E}_{\theta,{\mathcal{P}}_{\xi}}\!\left(|\hat{T}-L(\theta)|^{2}\right)\;\gtrsim\;|\eta_{1}|^{2}.$
Note that $\lambda_{o}^{2}s^{2}\asymp\lambda_{o}^{2}$ when $s\leq C_{1}$. From (5), define
$\pi_{j}=\frac{2}{s}\cdot\frac{|\eta_{j}|e^{-\beta/|\eta_{j}|^{\alpha}}}{\sqrt{\sum_{j=1}^{d}\eta_{j}^{2}e^{-\beta/|\eta_{j}|^{\alpha}}}},\quad\sum_{j=1}^{d}\pi_{j}=1.$
Then
$\nu^{2}=\sum_{j=1}^{d}\eta_{j}^{2}e^{-\beta_{+}/|\eta_{j}|^{\alpha}}=\sum_{j=1}^{d}\frac{s\pi_{j}}{2}\eta_{j}\nu\;\leq\;\frac{s|\eta_{1}|}{2}\nu\;\;\Rightarrow\;\;\nu\lesssim|\eta_{1}|.$
Hence, if $\lambda_{o}\lesssim|\eta_{1}|$, we have $\lambda_{o}^{2}s^{2}+\nu^{2}\lesssim|\eta_{1}|^{2}$. Thus, it remains to prove (51).
To this end, let $p\sim\mathrm{Mult}((\pi_{1},\ldots,\pi_{d}),1)$ and define $\theta$ by
$\theta_{j}=c\frac{\lambda_{o}}{\eta_{j}}\,\mathbbm{1}\{j=p\},\qquad j=1,\ldots,d,$
where $c>0$ is a constant to be specified later. Let $\mu$ denote the distribution of $\theta$. Then
$\mu\!\left(L(\theta)\geq c\lambda_{o},\theta\in\Theta_{s}\right)=1.$
We now compute the $\chi^{2}$-divergence between $P_{1}=\mathbb{P}_{\mu_{1},F_{\alpha}^{\otimes}}$ and $P_{2}=\mathbb{P}_{\mu,F_{\alpha}^{\otimes}}$. Conditioned on $p=j$, we have for $x=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}$:
$\mathbb{P}_{\mu,F_{\alpha}^{\otimes}}(x\mid p=j)=f_{\alpha}^{(0)}\!\left(x_{j}-c\frac{\lambda_{o}}{\eta_{j}}\right)\prod_{k\neq j}f_{\alpha}^{(0)}(x_{k}).$
Hence,
$1+\chi^{2}(P_{2}\|P_{1})$ $=\int_{\mathbb{R}^{d}}\Bigg[\sum_{j=1}^{d}\pi_{j}\frac{f_{\alpha}^{(0)}\!\left(x_{j}-c\frac{\lambda_{o}}{\eta_{j}}\right)}{f_{\alpha}^{(0)}(x_{j})}\Bigg]^{2}\,{\rm d\,}P_{1}$
$=\sum_{i\neq j}\pi_{i}\pi_{j}\int_{\mathbb{R}^{2}}f_{\alpha}^{(0)}\!\left(x_{i}-c\frac{\lambda_{o}}{\eta_{i}}\right)f_{\alpha}^{(0)}\!\left(x_{j}-c\frac{\lambda_{o}}{\eta_{j}}\right)\,{\rm d\,}x_{i}\,{\rm d\,}x_{j}$
$\quad+\sum_{j=1}^{d}\pi_{j}^{2}\int_{\mathbb{R}}\frac{f_{\alpha}^{(0)}\!\left(x_{j}-c\frac{\lambda_{o}}{\eta_{j}}\right)^{2}}{f_{\alpha}^{(0)}(x_{j})}\,{\rm d\,}x_{j}$
$\leq 1+\sum_{j=1}^{d}\pi_{j}^{2}\Bigg[C_{1}^{\alpha}\exp\!\left(\Big|\frac{c}{C_{\alpha,2}}\cdot\frac{\lambda_{o}}{\eta_{j}}\Big|^{\alpha}\right)-1\Bigg]$
$\leq 1+\frac{\sum_{j=1}^{d}\eta_{j}^{2}\exp\!\left(\big((c/C_{\alpha,2})^{\alpha}-2\big)\beta_{+}/|\eta_{j}|^{\alpha}\right)}{\sum_{j=1}^{d}\eta_{j}^{2}e^{-\beta_{+}/|\eta_{j}|^{\alpha}}}.$
Since
$\beta_{+}/|\eta_{j}|^{\alpha}\geq(\lambda_{o}/|\eta_{1}|)^{\alpha}\geq C_{3}^{\alpha},$
we can choose $c\in(0,1)$ and $C_{3}>0$ sufficiently large so that
$\mathrm{TV}(P_{1},P_{2})\leq\sqrt{\chi^{2}(P_{2}\|P_{1})}/2\leq\alpha_{1}$
for some $\alpha_{1}\in(0,1)$. Applying Lemma 1 with $t=\frac{1}{2}c\lambda_{o}$ completes the proof.
C.4 Proof of the lower bound in Theorem 8
Based on the following lemma, the proof of the lower bound is essential the same as that of Theorem 1.
Lemma 12 ([9, Lemma 3]).
If $\mu_{1}$ and $\mu_{2}$ are probability measures supported on $\Theta_{s,0}$ and $\Theta_{s}(\rho)$, respectively, then
$\inf_{\Delta}\left\{\sup_{\theta\in\Theta_{s,0}}\mathbb{P}_{\theta,{\cal N}^{\otimes}}(\Delta=1)+\sup_{\theta\in\Theta_{s}(\rho)}\mathbb{P}_{\theta,{\mathcal{N}}^{\otimes}}(\Delta=0)\right\}\;\geq\;1-{\rm TV}(\mathbb{P}_{\mu_{1},{\cal N}^{\otimes}},\mathbb{P}_{\mu_{2},{\cal N}^{\otimes}}),$
where $\inf_{\Delta}$ denotes the infimum over all $\{0,1\}$-valued statistics.
Since $s\geq 2$, we can select some $\theta_{0}\in\Theta_{s,0}$ such that $\|\theta_{0}\|_{0}\leq\lfloor s/2\rfloor$. We then let $\mu_{1}$ be the point mass at $\theta_{0}$, and construct $\mu_{2}$ as the distribution obtained by shifting $\theta_{0}$ with the least favorable alternatives used in the proof of Theorem 1.
C.5 Proofs of lemmas in Section C.1
Proof of Lemma 11.
We can write $P_{1}=\tilde{P}_{1}*\mu$ and $P_{2}=\tilde{P}_{2}*\mu$ where $\tilde{P}_{1}$ and $\tilde{P}_{2}$ denote the distribution of $(\sigma_{1}{\delta}_{1}-\pi_{1}\gamma_{1},\ldots,\sigma_{d}{\delta}_{d}-\pi_{d}\gamma_{d})$ and of $(\delta_{1},\ldots,\delta_{d})$, respectively. It suffices to prove upper bound ${\bf H}^{2}(\tilde{P}_{1},\tilde{P}_{2})$ using properties of the Hellinger distance for product measures:
${\bf H}^{2}(\tilde{P}_{1},\tilde{P}_{2})\leq 2\left(1-\prod_{j=1}^{d}(1-h_{j}^{2}/2)\right)\leq\sum_{j=1}^{d}h_{j}^{2}$
where $h_{j}^{2}=\int\left(\sqrt{f_{0}(x)}-\sqrt{\frac{f_{0}((x+\pi_{j}\gamma_{j})/\sigma_{j})}{\sigma_{j}}}\right)^{2}{\rm d\,}x$. Let $I_{1}(t)$ and $I_{2}(t;\sigma_{j})$ be the Fisher information of the scale family $\{f_{0}(x/t)/t:t>0\}$ and of the location family $\{f_{0}((x+t)/\sigma_{j})/\sigma_{j}:t\in[0,1]\}$, respectively. We have
$I_{1}(t)=t^{-2}I_{f_{0}},\quad I_{2}(t;\sigma_{j})=I_{1}(\sigma_{j})=\sigma_{j}^{-2}I_{f_{0}}.$
Note that Equation 49 implies $\sigma_{j}^{2}\geq 3/4$. By Lemma 3, we have
$\frac{1}{2}h_{j}^{2}$ $\leq\int\left(\sqrt{f_{0}(x)}-\sqrt{\frac{f_{0}(x/\sigma_{j})}{\sigma_{j}}}\right)^{2}{\rm d\,}x+\int\left(\sqrt{\frac{f_{0}(x/\sigma_{j})}{\sigma_{j}}}-\sqrt{\frac{f_{0}((x+\pi_{j}\gamma_{j})/\sigma_{j})}{\sigma_{j}}}\right)^{2}{\rm d\,}x$
$\leq\frac{(1-\sigma_{j})^{2}}{4}\sup_{t\in[\sigma_{j},1]}I_{1}(t)+\frac{\pi_{j}^{2}\gamma_{j}^{2}}{4}\sup_{t\in[0,\pi_{j}]}I_{2}(t;\sigma_{j})$
$\leq\frac{1}{3}I_{f_{0}}\sum_{j=1}^{d}\left((1-\sigma_{j})^{2}+\gamma_{j}^{2}\pi_{j}^{2}\right)$
$\leq\frac{1}{3}I_{f_{0}}\sum_{j=1}^{d}\left((1-\sigma_{j}^{2})^{2}+\gamma_{j}^{2}\pi_{j}^{2}\right)$
$\leq\frac{1}{3}I_{f_{0}}\sum_{j=1}^{d}\left(\gamma_{j}^{2}+\gamma_{j}^{4}\right)\pi_{j}^{2}.$
By the condition (29), $\sum_{j=1}^{d}\left(\gamma_{j}^{2}+\gamma_{j}^{4}\right)\pi_{j}^{2}$ can be bounded by some constant times $c^{2}c_{2}^{2}$. This completes the proof of the lemma. ∎
Appendix D Proof of Examples
D.1 Homogeneous loading vector
In this case, we have $\eta_{j}=1$ for all $j\in[d]$.
  • $\Phi_{\rm o}(s;\eta)$: From Equation 5, we obtain
    $\frac{s}{2}\;=\;\frac{d\exp(-\beta)}{\sqrt{d\exp(-\beta)}}\;=\;\sqrt{d}\,\exp(-\beta/2),$
    which implies
    $\beta=2\log\!\left(\frac{2\sqrt{d}}{s}\right).$
    • If $s\geq 2\sqrt{d}$, then $\lambda_{\mathcal{N}}=0$ and
      $\Phi_{\rm o}(s;\eta)\;\asymp\;d\;\asymp\;s^{2}\log^{2/\alpha}\!\left(1+\frac{d^{\alpha/2}}{s^{\alpha}}\right).$
    • If $s<2\sqrt{d}$, then
      $\Phi_{\rm o}(s;\eta)\;\asymp\;d\exp(-\beta)+s^{2}\lambda_{\mathcal{N}}^{2}\;\asymp\;s^{2}\log^{2/\alpha}\!\left(\frac{2\sqrt{d}}{s}\right)\;\asymp\;s^{2}\log^{2/\alpha}\!\left(1+\frac{d^{\alpha/2}}{s^{\alpha}}\right).$
    Therefore,
    $\Phi_{\rm o}(s;\eta)\;\asymp\;s^{2}\log^{2/\alpha}\!\left(1+\frac{d^{\alpha/2}}{s^{\alpha}}\right).$
  • $\Phi_{\rm adp}(s;\eta)$: From (17), we have
    $\Phi_{*}(s;\eta)\;\asymp\;\log(es)\,\Phi\!\left(\frac{s}{\sqrt{\log(es)}};\eta\right)\;\asymp\;s^{2}\log^{2/\alpha}\!\left(1+\frac{d^{\alpha/2}\log^{\alpha/2}(es)}{s^{\alpha}}\right)\;\gtrsim\;\Phi_{*}(1;\eta)\,\log^{2/\alpha}(es).$
    Hence,
    $\Phi_{\rm adp}(s;\eta)\;\asymp\;\Phi_{*}(s;\eta)\;\asymp\;s^{2}\log^{2/\alpha}\!\left(1+\frac{d^{\alpha/2}\log^{\alpha/2}(es)}{s^{2}}\right).$
  • $\Phi_{\rm ns}(s;\eta)$: From Theorem 6, the upper bound is
    $\sum_{j\leq(s^{2}\wedge d)}\eta_{j}^{2}\;=\;j_{3}(s)=\big\lceil s^{2}\log^{2/\alpha}(ed/s)\big\rceil\wedge d.$
    • If $s\lesssim{\sqrt{d}}/{\log^{2/\alpha}(ed)}$, then by (27),
      $\lambda_{\cal H}\;=\;\log^{1/\alpha}\!\left(\frac{d-s^{2}}{s}\right)\;\asymp\;\log^{1/\alpha}d.$
      In this regime, condition (29) holds, and thus the asymmetric lower bound is $s^{2}\log^{2/\alpha}(ed),$ as given in (30).
    • If $s\gtrsim\sqrt{d}$, then the lower bound reduces to $d,$ from (28).
D.2 Two-phase loading vector
In this case, we have $\eta_{j}=d^{\gamma_{\lambda}}$ for $1\leq j\leq d^{\gamma_{d}}$ and $\eta_{j}=1$ for $d^{\gamma_{d}}<j\leq d$. Let $\eta^{(1)},\eta^{(2)}$ denote the two subvectors of $\eta$, i.e.,
$\eta^{(1)}\in\mathbb{R}^{d^{\gamma_{d}}},\quad\eta^{(1)}_{j}=d^{\gamma_{\lambda}},\ \forall j\leq d^{\gamma_{d}},$
and
$\eta^{(2)}\in\mathbb{R}^{d-d^{\gamma_{d}}},\quad\eta^{(2)}_{j}=1,\ \forall j\leq d-d^{\gamma_{d}}.$
Following the analysis in the homogeneous case, we obtain
$\Phi_{\rm o}(s;\eta^{(1)})\;\asymp\;d^{2\gamma_{\lambda}}s^{2}\log^{2/\alpha}\!\left(1+\frac{d^{\gamma_{d}\alpha/2}}{s^{\alpha}}\right)$
and
$\Phi_{\rm o}(s;\eta^{(2)})\;\asymp\;s^{2}\log^{2/\alpha}\!\left(1+\frac{(d-d^{\gamma_{d}})^{\alpha/2}}{s^{\alpha}}\right).$
Write $\Theta^{(1)}=\{\theta\in\Theta_{s}:\theta_{j}=0,\forall j>d^{\gamma_{d}}\}$ and $\Theta^{(2)}=\{\theta\in\Theta_{s}:\theta_{j}=0,\forall j\leq d^{\gamma_{d}}\}$. We then have $\Theta^{(1)}\cup\Theta^{(2)}\subset\Theta_{s}$. It is straightforward to see from the definition of minimax rate that $\Phi_{\rm o}(s;\eta)\;\gtrsim\;\Phi_{\rm o}(s;\eta^{(i)})$ for $i=1,2$.
We also have $\Theta_{s}\subset\Theta^{(1)}+\Theta^{(2)}$: For any $\theta$, we can write $\theta=\theta^{(1)}+\theta^{(2)}$ so that $\theta^{(i)}\in\Theta^{(i)}$. It follows that $L(\theta)=\langle\eta,\theta\rangle=\langle\eta^{(1)},\theta^{(1)}\rangle+\langle\eta^{(2)},\theta^{(2)}\rangle$. Therefore, for any minimax estimators for $\langle\eta^{(1)},\theta^{(1)}\rangle$ and $\langle\eta^{(2)},\theta^{(2)}\rangle$, the sum of their worst case risks is an upper bound on the minimax risk for $L(\theta)$. This implies that $\Phi_{\rm o}(s;\eta)\lesssim\Phi_{\rm o}(s;\eta^{(1)})+\Phi_{\rm o}(s;\eta^{(2)})$.
To sum up, the minimax rate for estimating $\langle\eta,\theta\rangle$ satisfies
$\Phi_{\rm o}(s;\eta)\;\asymp\;\Phi_{\rm o}(s;\eta^{(1)})+\Phi_{\rm o}(s;\eta^{(2)}).$
Therefore,
$\Phi_{\rm o}(s;\eta)\;\asymp\;d^{2\gamma_{\lambda}}s^{2}\log^{2/\alpha}\!\left(1+\frac{d^{\gamma_{d}\alpha/2}}{s^{\alpha}}\right)+s^{2}\log^{2/\alpha}\!\left(1+\frac{d^{\alpha/2}}{s^{\alpha}}\right),$ (52)
and
$\Phi_{\rm adp}(s;\eta)$ $\;\asymp\;\log(es)\,\Phi_{\rm o}\!\left(\frac{s}{\sqrt{\log(es)}};\eta\right)\;\vee\;\Phi_{\rm o}(1;\eta)\log^{(2)}(es)$
$\;\asymp\;d^{2\gamma_{\lambda}}s^{2}\log^{2/\alpha}\!\left(1+\frac{d^{\gamma_{d}\alpha/2}\log^{\alpha/2}(es)}{s^{\alpha}}\right)+s^{2}\log^{2/\alpha}\!\left(1+\frac{d^{\alpha/2}\log^{\alpha/2}(es)}{s^{\alpha}}\right).$
We further simplify Equation 52 as follows.
  1. 1.
    When $s\lesssim d^{\gamma_{d}/2}$, both the logarithmic factors are equivalent to $\log(d)$ so the first term dominates.
  2. 2.
    When $d^{\gamma_{d}/2}\ll s\ll d^{\gamma_{\lambda}+\gamma_{d}/2}$ (which is $\ll\sqrt{d}$ by assumption), the second logarithmic factor remains the same as $\log(d)$ but the first factor becomes $\asymp\left(\frac{d^{\gamma_{d}/2}}{s}\right)^{\alpha}$. In this case, the two terms are $d^{2\gamma_{\lambda}+\gamma_{d}}$ and $s^{2}\log^{2/\alpha}(d)$. This suggests that the first term dominates if and only if $s\lesssim\frac{d^{2\gamma_{\lambda}+\gamma_{d}}}{\log^{2/\alpha}(d)}$.
D.3 Exponentially decaying loading vector
Some of the calculations for this example are borrowed from Example E.4 in [8].
Recall that $j_{0}=\min\{\,j\mid\eta_{j}<1/2\,\}$. We first show that for any $\gamma\geq 1$, the following relation holds:
$\sum_{j\geq j_{0}}\eta_{j}^{\gamma}\exp\left(-\beta/\eta_{j}^{\alpha}\right)\lesssim\sum_{j<j_{0}}\eta_{j}^{\gamma}\exp\left(-\beta/\eta_{j}^{\alpha}\right).$ (53)
We write $j_{*}=j_{0}-1$ and see that if $j_{*}<d$ then $\phi(j_{*})>\log(2)$. If $j_{*}=d$, the relation is obvious. Otherwise, for any $\beta\in\mathbb{R}$, we can use the monotonicity and the convexity of $\phi$ to derive as follows:
$\sum_{j\geq j_{0}}\eta_{j}^{\gamma}\exp\left(-\beta/\eta_{j}^{\alpha}\right)=$ $\sum_{j\geq j_{0}}\exp\left(-\beta/\eta_{j}^{\alpha}-\gamma\phi(j-1)\right)$
$(\because\phi\text{ is non-decreasing })\qquad\leq$ $\sum_{j\geq j_{*}}\exp\left(-\beta/\eta_{j_{0}}^{\alpha}-\gamma\phi(j)\right)$
$(\because\phi\left(j_{*}+\ell\right)\geq\phi\left(j_{*}\right)+\ell\phi^{\prime}\left(j_{*}\right),\quad\forall\ell\geq 0)\qquad\leq$ $\sum_{\ell\geq 0}\exp\left(-\beta/\eta_{j_{0}}^{\alpha}-\gamma\phi(j_{*})-\gamma\ell\phi^{\prime}(j_{*})\right)$
$\leq\frac{\exp\left(-\beta/\eta_{j_{0}}^{\alpha}-\gamma\phi(j_{*})\right)}{1-\exp\left(-\gamma\ell\phi^{\prime}(j_{*}))\right)}$
$(\because\text{definition of }j_{*})$ $\leq\frac{\exp\left(-\beta/\eta_{j_{0}}^{\alpha}\right)}{1-\exp\left(-\gamma\ell\phi^{\prime}(j_{*}))\right)}\times 2^{-\gamma}.$
Furthermore, the monotonicity and the convexity of $\phi$ implies
$\phi(j_{*})\leq\phi(j)\leq\frac{\phi\left(j_{*}\right)}{j_{*}}\cdot j,\quad\forall j\leq j_{*}.$
Therefore,
$\sum_{1\leq j<j_{0}}\eta_{j}^{\gamma}\exp\left(-\beta/\eta_{j}^{\alpha}\right)=$ $\sum_{0\leq j\leq j_{*}}\exp\left(-\beta/\eta_{j+1}^{\alpha}-\gamma\phi(j)\right)$
$\geq$ $\sum_{j\leq j_{*}}\exp\left(-\beta/\eta_{j_{0}}^{\alpha}-\gamma j\phi(j_{*})/j_{*}\right)$
$=$ $\exp\left(-\beta/\eta_{j_{0}}^{\alpha}\right)\frac{1-\exp\left(-\gamma j_{*}\phi(j_{*})/j_{*}\right)}{1-\exp\left(-\gamma\phi(j_{*})/j_{*}\right)}$
$(\because\text{definition of }j_{*})\qquad\geq$ $\exp\left(-\beta/\eta_{j_{0}}^{\alpha}\right)\frac{1-2^{-\gamma}}{1-\exp\left(-\gamma\phi(j_{*})/j_{*}\right)}$
$(\because\phi(j_{*})/j_{*}\leq\phi^{\prime}(j_{*}))\qquad\geq$ $\exp\left(-\beta/\eta_{j_{0}}^{\alpha}\right)\frac{1-2^{-\gamma}}{1-\exp\left(-\gamma\phi^{\prime}(j_{*})\right)}.$
Therefore, for any $\beta$, we have
$\frac{\sum_{j\geq j_{0}}\eta_{j}^{\gamma}\exp\left(-\beta/\eta_{j}^{\alpha}\right)}{\sum_{1\leq j<j_{0}}\eta_{j}^{\gamma}\exp\left(-\beta/\eta_{j}^{\alpha}\right)}\leq 2^{\gamma}-1.$ (54)
We take $\gamma=1$ and $\gamma=2$ to simplify the numerator and denominator in Equation 5 into
$\sum_{j<j_{0}}|\eta_{j}|\exp(-\beta/|\eta_{j}|^{\alpha}),\text{ and }\sqrt{\sum_{j<j_{0}}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})}$
up to a constant. Furthermore, for any $j<j_{0}$, we have $1/2<\eta_{j}\leq 1$. Therefore, the solution $\beta=\beta(s;\eta)$ to Equation 5 satisfies that
$\frac{s}{2}\asymp\frac{\sum_{j<j_{0}}|\eta_{j}|\exp(-\beta/|\eta_{j}|^{\alpha})}{\sqrt{\sum_{j<j_{0}}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})}}\asymp\frac{\sum_{j<j_{0}}\exp(-\beta/|\eta_{j}|^{\alpha})}{\sqrt{\sum_{j<j_{0}}\exp(-\beta/|\eta_{j}|^{\alpha})}},$ (55)
which is lower bounded by $\sqrt{j_{0}\exp(-2^{\alpha}\beta)}$ and upper bounded $\sqrt{j_{0}\exp(-\beta)}$ up to absolute constants.
We discuss as follows:
  1. 1.
    Suppose $\beta>1$, by (55), we have
    $\sum_{j<j_{0}}\exp(-\beta/|\eta_{j}|^{\alpha})\asymp s^{2}.$
    Since $\eta_{j}\in(1/2,1)$ for any $j<j_{0}$, we have $\lambda_{o}=\beta^{1/\alpha}\asymp\log^{1/\alpha}(1+j_{0}^{\alpha/2}/s^{\alpha}).$ By Equation 54 with $\gamma=2$, we have
    $\nu^{2}(s)\asymp\sum_{j<j_{0}}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})\asymp\sum_{j<j_{0}}\exp(-\beta/|\eta_{j}|^{\alpha})\asymp s^{2}.$
    Therefore, we have
    $\Phi_{\rm o}(s;\eta)\asymp s^{2}\lambda_{o}^{2}+\nu^{2}\asymp s^{2}\lambda_{o}^{2}\asymp s^{2}\log^{1/\alpha}(1+j_{0}^{\alpha/2}/s^{\alpha}).$
  2. 2.
    Suppose $\beta\leq 1$, we have $\lambda_{o}\asymp\log^{1/2}_{+}(j_{0}/s^{2})$ and therefore, $s^{2}\lambda_{o}^{2}\leq j_{0}$. By Equation 54 with $\gamma=2$, we have
    $\nu^{2}\asymp\sum_{j<j_{0}}\eta_{j}^{2}\exp(-\beta/|\eta_{j}|^{\alpha})\asymp\sum_{j<j_{0}}\exp(-\beta/|\eta_{j}|^{\alpha})\asymp j_{0}.$
    Therefore, we have
    $\Phi_{\rm o}(s;\eta)\asymp s^{2}\lambda_{o}^{2}+\nu^{2}\asymp j_{0}\asymp s^{2}\log^{1/\alpha}(1+j_{0}^{\alpha/2}/s^{\alpha})$
Combining the two cases, we have $\Phi_{\rm o}(s;\eta)\asymp\Phi_{\rm o}(s;\mathbf{1}_{j_{0}})$. It then follows that $\Phi_{\rm adp}(s;\eta)\asymp\Phi_{\rm adp}(s;\mathbf{1}_{j_{0}})$.